### v1

# Exact Random Graph Matching with Multiple Graphs

Abstract
This work studies fundamental limits for recovering the underlying correspondence among multiple correlated random graphs. We identify a necessary condition for any algorithm to correctly match all nodes across all graphs, and propose two algorithms for which the same condition is also sufficient. The first algorithm employs global information to simultaneously match all the graphs, whereas the second algorithm first partially matches the graphs pairwise and then combines the partial matchings by transitivity. Both algorithms work down to the information theoretic threshold. Our analysis reveals a scenario where exact matching between two graphs alone is impossible, but leveraging more than two graphs allows exact matching among all the graphs. Along the way, we derive independent results about the $k$-core of Erdős-Rényi graphs.
1 Introduction
The information age has ushered an abundance of correlated networked data. For instance, the network structure of two social networks such as Facebook and Twitter is correlated because users are likely to connect with the same individuals in both networks. This wealth of correlated data presents both opportunities and challenges. On one hand, information from various datasets can be combined to increase the fidelity of data - translating to better performance in downstream learning tasks. On the other hand, the interconnected nature of this data also raises privacy and security concerns. Linkage attacks, for instance, exploit correlated data to identify individuals in an anonymized network by linking to other sources [NS09]. This poses a significant threat to user privacy.
Graph matching is the problem of recovering the underlying latent correspondence between correlated networks. The problem finds many applications in machine learning: de-anonymizing social networks [NS08, NS09], identifying similar functional components between species by matching their protein-protein interaction networks [BSI06, KHGPM16], object detection [SS05] and tracking [YYL+16] in computer vision, and textual inference for natural language processing [HNM05]. In most applications of interest, data is available in the form of several correlated networks. For instance, social media users are active each month on 6.7 social platforms on average [Ind23]. Similarly, reconciling protein-protein interaction networks among multiple species is an important problem in computational biology [SXB08]. As a first step toward this objective, many research works have studied the problem of matching two correlated graphs.
1.1 Related Work
The theoretical study of graph matching algorithms and their performance guarantees has primarily focused on Erdős-Rényi (ER) graphs. Pedarsani and Grossglauser [PG11] introduced the subsampling model to generate two such correlated graphs. The model entails twice subsampling each edge independently from a parent ER graph to obtain two sibling graphs, both of which are marginally ER graphs themselves. The goal is then to match nodes between the two graphs to recover the underlying latent correspondence. This has been the framework of choice for many works that study graph matching. For example, Cullina and Kiyavash studied the problem of exactly matching two ER graphs, where the objective is to match all vertices correctly [CK16, CK17]. They identified a threshold phenomenon for this task: exact recovery is possible if the problem parameters are above a threshold, and impossible otherwise. Subsequently, threshold phenomena were also identified for partial graph matching between ER graphs - where the objective is to match only a positive fraction of nodes [GML21, HM23, WXY22, DD23]. The case of almost-exact recovery - where the objective is to match all but a negligible fraction of nodes - was studied by Cullina and co-authors: a necessary condition for almost exact recovery was identified, and it was shown that the same condition is also sufficient for the $k$-core estimator [CKMP19]; the estimator is described formally in Section 3. This estimator proved useful to uncover the fundamental limits for graph matching in other contexts such as the stochastic block model [GRS22] and inhomogeneous random graphs [RS23]. Ameen and Hajek [AH23] showed some robustness properties of the $k$-core estimator in the context of matching ER graphs under node corruption. The estimator plays an important role in the present work as well.
A sound understanding of ER graphs inspires algorithms for real-world networks. Various efficient algorithms have been proposed, including algorithms based on the spectrum of the graph adjacency matrices [FMWX22], node degree and neighborhood based algorithms [DCKG19, DMWX21, MRT23] as well as algorithms based on iterative methods [DL23] and counting subgraphs [MWXY23, BCL+19]. Some of these are discussed in Section 5 in relation to the present work.
Incorporating information from multiple graphs to match them has been recognized as an important research direction, for instance in the work of Gaudio and co-authors [GRS22]. To our knowledge, the only other work to consider matchings among multiple graphs is the work of Rácz and Sridhar [RS21]. However, the authors solve a different problem of matching stochastic block models, and note that it is possible to exactly match $m$ graphs whenever it is possible to exactly match any two graphs by pairwise matching all the graphs exactly. In contrast, we show that under appropriate conditions, it is possible to exactly match $m$ ER graphs even when no two graphs can be pairwise matched exactly.
Contributions
In this work, we investigate the problem of combining information from multiple correlated networks to boost the number of nodes that are correctly matched among them. We consider the natural generalization of the subsampling model to generate $m$ correlated random graphs, and identify a threshold such that it is impossible for any algorithm to match all nodes correctly across all graphs when the problem parameters are below this threshold. Conversely, we show that exact recovery is possible above the threshold. This characterization generalizes known results for exact graph matching when $m=2$. Subsequently, we show that there is a region in parameter space for which exactly matching any two graphs is impossible using only the two graphs, and yet exact graph matching is possible among $m>2$ graphs using all the graphs.
We present two algorithms and prove their optimality for this task. The first algorithm matches all $m$ graphs simultaneously based on global information about the graphs. In contrast, the second algorithm first pairwise matches graphs, and then combines them to match all nodes across all graphs. We show that both algorithms correctly match all the graphs all the way down to the information theoretic threshold. Finally, we illustrate through simulation that our subroutine to combine information from pairwise comparisons between networks works well when paired with efficient algorithms for graph matching. Our analysis also yields some theoretical results about the $k$-core of ER graphs that are of independent interest.
2 Preliminaries and Setup
Notation
In this work, $G\sim\mathsf{ER}(n,p)$ denotes that the graph $G$ is sampled from the Erdős-Rényi distribution with parameters $n$ and $p$, i.e. $G$ has $n$ nodes and each edge is independently present with probability $p$. For a graph $G$, we denote the set of its vertices by $V\equiv V(G)$ and its edges by $E(G)$. The edge status of each vertex pair $\{i,j\}$ with $i\neq j$ is denoted by $G\{i.j\}$, so that $G\{i,j\}=1$ if $\{i,j\}\in E(G)$ and $G\{i,j\}=0$ otherwise. The degree of a node $v$ in graph $G$ is denoted $\delta_{G}(v)$. Let $\pi$ denote a permutation on $V(G)=\{1,\cdots,n\}$. For a graph $G$, denote by $G^{\pi}$ the graph obtained by permuting the nodes of $G$ according to $\pi$, so that
$G\{i,j\}=G^{\pi}\left\{\pi(i),\pi(j)\right\}\ \forall\ i,j\in V(G)\text{ such that }i\neq j.$
Standard asymptotic notation $(O(\cdot),o(\cdot),\cdots)$ is used throughout and it is implicit that $n\to\infty$.
Subsampling model
Consider the subsampling model for correlated random graphs [PG11], which has a natural generalization to the setting of $m$ graphs. In this model, a parent graph $G$ is sampled from the Erdős-Rényi distribution $\mathsf{ER}(n,p)$. The $m$ graphs $G_{1},G_{2}^{\prime},\cdots,G_{m-1}^{\prime},G_{m}^{\prime}$ are obtained by independently subsampling each edge from $G$ with probability $s$. Finally, the graphs $G_{2},\cdots,G_{m}$ are obtained by permuting the nodes of each of the graphs $G_{2}^{\prime},\cdots,G_{m}^{\prime}$ respectively according to independent permutations $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ sampled uniformly at random from the set of all permutations on $[n]$, i.e.
$G_{j}=(G_{j}^{\prime})^{\pi^{*}_{1j}}\ \text{ for all }\ j\in\{2,\cdots,m\}.$
Figure 1 illustrates this process of obtaining correlated graphs using the subsampling model. In this work, we are interested in the setting where $s$ is constant and $p=C\log(n)/n$ for some $C>0$.
Objective 1.
Determine conditions on parameters $C$, $s$ and $m$ so that given correlated graphs $G_{1},\cdots,G_{m}$ from the subsampling model, it is possible to exactly recover the underlying correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ with probability $1-o(1)$.
Stated thus, the underlying correspondences use the graph $G_{1}$ as a reference. Thus, for ease of notation, we will use $G_{1}$ and $G_{1}^{\prime}$ interchangeably. Note that the underlying correspondence between all the graphs is fixed upon fixing $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$: for any two graphs $G_{i}$ and $G_{j}$, their underlying correspondence is given by $\pi^{*}_{ij}:=\pi^{*}_{1j}\circ(\pi^{*}_{1i})^{-1}$.
Formally, a matching $(\mu_{12},\cdots,\mu_{1m})$ is a collection of injective functions with domain $\mathsf{dom}(\mu_{1i})\subseteq V$ for each $i$, and co-domain $V$. An estimator is simply a mechanism to map any collection of graphs $(G_{1},\cdots,G_{m})$ to a matching. We say that an estimator completely matches the graphs if the output mappings $\mu_{12},\cdots\mu_{1m}$ are all complete, i.e. they are all permutations on $\{1,\cdots,n\}$.
Figure 1: Illustration of obtaining $m$ correlated graphs from the subsampling model
3 Main Results and Algorithm
This section presents necessary and sufficient conditions to meet 1.
Theorem 2 (Impossibility).
Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively. Suppose that
$Cs\left(1-(1-s)^{m-1}\right)<1.$
The output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ of any estimator satisfies
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=o(1).$
Theorem 2 implies that the condition $Cs(1-(1-s)^{m-1}>1$ is a necessary condition to exactly match $m$ graphs with probability bounded away from 00. We show that this condition is also sufficient to exactly match $m$ graphs with probability going to $1$.
Theorem 3 (Achievability).
Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively. Suppose that
$Cs\left(1-(1-s)^{m-1}\right)>1.$
There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies
$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1).$
Theorems 2 and 3 together characterize the threshold for exact recovery. A few remarks are in order.
  1. 1.
    For $m=2$, the condition $Cs(1-(1-s)^{m-1})>1$ reduces to $Cs^{2}>1$, which is known to be necessary and sufficient for exactly matching two graphs [CK17, WXY22].
  2. 2.
    For any $m>2$, there is a non-empty region in the parameter space defined by
    $Cs(1-(1-s)^{m-1})>1>Cs^{2}.$
    For any $C$ and $s$ in this region, it is impossible to exactly match any two graphs $G_{i}$ and $G_{j}$ without using the other $m-2$ graphs as side information. Upon using them, however, it is possible to exactly match all nodes across the $m$ graphs. This is illustrated in Figure 2.
(a) $m=3$
(b) $m=10$
Figure 2: Regions in parameter space. Orange: Exactly matching $m$ graphs is impossible even with $m$ graphs. Blue: Exactly matching $2$ graphs is possible with $2$ graphs. Striped: Impossible to match $2$ graphs using only the $2$ graphs, but possible using $m$ graphs as side information.
3.1 Algorithms for exact recovery
For any two graphs $H_{1}$ and $H_{2}$ on the same vertex set $V$, denote by $H_{1}\vee H_{2}$ their union graph and by $H_{1}\wedge H_{2}$ their intersection graph. An edge $\{i,j\}$ is present in $H_{1}\vee H_{2}$ if it is present in either $H_{1}$ or $H_{2}$. Similarly, the edge is present in $H_{1}\wedge H_{2}$ if it is present in both $H_{1}$ and $H_{2}$.
A natural starting point is to study the maximum likelihood estimator (MLE) because it is optimal. To that end, we compute the log-likelihood function; the details are deferred to Appendix A.
Theorem 4.
Let $\pi_{12},\cdots,\pi_{1m}$ denote a collection of permutations on $\{1,\cdots,n\}$. Then
$\log\mathbb{P}\left(G_{1},\cdots,G_{m}\ |\ \pi^{*}_{12}=\pi_{12},\cdots,\pi^{*}_{1m}=\pi_{1m}\right)\propto\text{const.}-\left|E\left(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}\right)\right|,$
where const. depends only on $p,s$ and $G_{1},\cdots,G_{m}$.
Theorem 4 reveals that the MLE for exactly matching $m$ graphs has a neat interpretation: simply pick $\pi_{12},\cdots,\pi_{1m}$ to minimize the number of edges in the corresponding union graph. This is presented as Algorithm 1. Despite this nice interpretation of the MLE, its analysis is quite cumbersome. We instead present and analyze a different estimator, presented as Algorithm 2.
require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$
1 for $(\pi_{12},\pi_{13},\cdots,\pi_{1m})$ such that each $\pi_{1j}$ is a permutation on $[n]$ do
2       $W(\pi_{12},\cdots,\pi_{1m})\leftarrow|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|$
3
4 end for
return $(\widehat{\pi}_{12}^{\text{ML}},\cdots,\widehat{\pi}_{1m}^{\text{ML}})\in\operatorname*{arg\,max}_{\pi_{12},\cdots,\pi_{1m}}W(\pi_{12},\cdots,\pi_{1m})$
Algorithm 1 Maximum likelihood estimator
require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$, Integer $k$
// Step 1: Pairwise matching
1 for $\{i,j\}$ in $\{1,\cdots,m\}$ such that $i<j$ do
2       $\widehat{\nu}_{ij}\leftarrow\operatorname*{arg\,max}_{\pi}|\mathsf{core}_{k}\left(G_{i}\wedge G_{j}^{\pi}\right)|$
       $\widehat{\mu}_{ij}\leftarrow\widehat{\nu}_{ij}\text{ with domain restricted to}\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$
        // $k$-core estimator
3
4 end for
// Step 2: Boosting through transitive closure
5 for $v\in V$ do
6       for $j=2,\cdots,m$ do
7             if  there is a sequence of indices $1=k_{1},\cdots,k_{\ell}=j$ in $[m]$ such that $\widehat{\mu}_{k_{\ell-1},j}\circ\cdots\circ\widehat{\mu}_{k_{2},k_{3}}\circ\widehat{\mu}_{1,k_{2}}(v)=v^{\prime}$ for some $v^{\prime}\in[n]$ then
8                   Set $\widehat{\pi}_{1j}(v)=v^{\prime}$
9
10             end if
11
12       end for
13
14 end for
return $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$
Algorithm 2 Matching through transitive closure
Algorithm 2 runs in two steps: In step 1, the $k$-core estimator, for a suitable choice of $k$, is used to pairwise match all the graphs. For any $i$ and $j$, the $k$-core estimator selects a permutation $\widehat{\nu}_{ij}$ to maximize the size of the $k$-core [^1] The $k$-core of a graph $G$ is the largest subset of vertices $\mathsf{core}_{k}(G)$ such that the induced subgraph has minimum degree at least $k$. of $G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}}$. It then outputs a matching $\widehat{\mu}_{ij}$ by restricting the domain of $\widehat{\nu}_{ij}$ to $\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$. These matchings $\widehat{\mu}_{ij}$ need not be complete - in fact, each of them is a partial matching with high probability whenever $Cs^{2}<1$. In step 2, these partial matchings are boosted as follows: If a node $v$ is unmatched between two graphs $G_{i}$ and $G_{j}$, then search for a sequence of graphs $G_{i},G_{k_{1}},\cdots,G_{k_{\ell}},G_{j}$ such that $v$ is matched between any two consecutive graphs in the sequence. If such a sequence exists, then extend $\widehat{\mu}_{i,j}$ to include $v$ by transitively matching it from $G_{i}$ to $G_{j}$.
In Section 4.2, we show that Algorithm 2 correctly matches all nodes across all graphs with probability $1-o(1)$, whenever the necessary condition $Cs(1-(1-s)^{m-1})>1$ holds. We remark that this also implies that Algorithm 1 succeeds under the same condition, because the MLE is optimal. Note that the MLE selects all permutations $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ simultaneously based on their union graph. In contrast, Algorithm 2 only ever makes pairwise comparisons between graphs. Perhaps surprisingly, it turns out that this is sufficient for exact recovery. An analysis of Algorithm 2 is presented in Section 4. Along the way, independent results of interest on the $k$-core of Erdős-Rényi graphs are obtained.
4 Proof Outlines and Key Insights
4.1 Impossibility of exact graph matching (Theorem 2)
This result has a simple proof following a genie-aided converse argument. The idea is to reduce the problem to that of matching two graphs by providing extra information to the estimator.
Proof of Theorem 2. If the correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1,m-1}$ were provided as extra information to an estimator, then the estimator must still match $G_{m}$ with the union graph $G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$. This can be viewed as an instance of matching two graphs obtained by asymmetric subsampling: the graph $G_{m}$ is obtained from a parent graph $G\sim\mathsf{ER}(n,C\log(n)/n)$ by subsampling each edge independently with probability $s_{1}:=s$, and the graph $\widetilde{G}_{m-1}:=G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$ is obtained from $G$ by subsampling each edge independently with probability $s_{2}:=1-(1-s)^{m-1}$. Cullina and Kiyavash studied this model for matching two graphs: Theorem 2 of [CK17] establishes that matching $G_{m}$ and $\widetilde{G}_{m-1}$ is impossible if $Cs_{1}s_{2}<1$, or equivalently if $Cs(1-(1-s)^{m-1})<1$. ∎
4.2 Achievability of exact graph matching (Theorem 3)
Algorithm 2 succeeds if both step 1 and step 2 succeed, i.e.
  1. 1.
    Each instance of pairwise matching using the $k$-core estimator is correct on its domain, i.e.
    $\widehat{\mu}_{ij}(v)=\pi^{*}_{ij}(v)\ \forall v\in\mathsf{dom}(\widehat{\mu}_{ij}),\ \forall i,j.$
  2. 2.
    For each node $v$ and any two graphs $G_{i}$ and $G_{j}$, there is a sequence of graphs such that $v$ can be transitively matched through those graphs between $G_{i}$ and $G_{j}$.
On step 1
This falls back to the regime of analyzing the performance of the $k$-core estimator in the setting of two graphs. Cullina and co-authors [CKMP19] showed that the $k$-core estimator is precise: For any two correlated graphs $G_{i}$ and $G_{j}$ with $p=C\log(n)/n$ and constant $s$, the $k$-core estimator correctly matches all nodes in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ with probability $1-o(1)$. In fact, this is true for any $C>0$ and for any $k\geq 13$ [RS23]. Therefore, using the fact that the number of instances of pairwise matchings is constant whenever $m$ is constant, a union bound reveals
[…]
$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\mathbb{P}\left(\widehat{\mu}_{i,j}(v)\neq\pi^{*}_{i,j}(v)\text{ for some }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right)$
$=o(1).$
We have proved the following.
Proposition 5.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model. Let $k\geq 13$ and let $\widehat{\mu}_{ij}$ denote the matching output by the $k$-core estimator on graphs $G_{i}$ and $G_{j}$. Then,
( $\mathbb{P}(\exists\ 1\leq i<j\leq m,\text{ and }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime}))\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))=o(1).$
On step 2
The challenging part of the proof is to show that boosting through transitive closure matches all the nodes with probability $1-o(1)$ if $Cs(1-(1-s)^{m-1})>1$. It is instructive to visualize this using transitivity graphs.
Definition 6 (Transitivity graph, $\mathcal{H}(v)$).
For each node $v\in V$, let $\mathcal{H}(v)$ denote the graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that an edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$.
On the event that each instance of pairwise matching using the $k$-core is correct, the edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v$ is correctly matched using the $k$-core estimator between $G_{i}$ and $G_{j}$, i.e. $\pi^{*}_{1i}(v)$ is matched to $\pi^{*}_{1j}(v)$. Thus, in order for Step 2 to succeed (i.e. to exactly match all vertices across all graphs), it suffices that the graph $\mathcal{H}(v)$ is connected for each node $v\in V$. However, studying the connectivity of the transitivity graphs is challenging because in any graph $\mathcal{H}(v)$, no two edges are independent. This is because the $k$-cores of any two intersection graphs $G_{a}^{\prime}\wedge G_{b}^{\prime}$ and $G_{c}^{\prime}\wedge G_{d}^{\prime}$ are correlated, because all the graphs $G_{a},G_{b},G_{c}$ and $G_{d}$ are themselves correlated. To overcome this, we introduce another graph ~$\widetilde{\mathcal{H}}(v)$ that relates to $\mathcal{H}(v)$ and is amenable to analysis.
Definition 7.
For each node $v\in V$, let ~$\widetilde{\mathcal{H}}(v)$ denote a complete weighted graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that the weight on any edge $\{g_{i},g_{j}\}$ is $\widetilde{c}_{v}\left(i,j\right):=\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v).$
The relationship between the graphs $\mathcal{H}(v)$ and ~$\widetilde{\mathcal{H}}(v)$ stems from a useful relationship between the degree of node $v$ in $G_{i}^{\prime}\wedge G_{j}^{\prime}$ and the inclusion of $v$ in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ for each $i$ and $j$. Since this result is of independent interest in the study of random graphs, we state it below for general Erdős-Rényi graphs.
Lemma 8.
Let $n$ and $k$ be positive integers and let $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$ for some $\alpha>0$. Let $v$ be a node of $G$ and let $\delta_{G}(v)$ denote the degree of $v$ in $G$. Then,
($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)=o\left(1/n\right).$ (1)
For any $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}(n,Cs^{2}\log(n)/n)$. Thus, Lemma 8 implies that with probability $1-o(1/n)$, if a pair $\{g_{i},g_{j}\}$ has edge weight $\widetilde{c}_{ij}\geq k+1/\alpha$ in ~$\widetilde{\mathcal{H}}(v)$, then the corresponding edge $\{g_{i},g_{j}\}$ is present in the transitivity graph $\mathcal{H}(v)$. Equivalently, $v$ is correctly matched between $G_{i}$ and $G_{j}$ in the instance of pairwise $k$-core matching between them.
The graph $\mathcal{H}(v)$ is not connected only if it contains a (non-empty) vertex cut $U\subset\{1,\cdots,m\}$ with no edge crossing between $U$ and $U^{c}$. Let $c_{v}(U)$ denote the number of such crossing edges in $\mathcal{H}(v)$. Furthermore, define the cost of the cut $U$ in ~$\widetilde{\mathcal{H}}(v)$ as
$\widetilde{c}_{v}(U):=\sum_{i\in U}\sum_{j\in U^{c}}\widetilde{c}_{v}\left(i,j\right).$
Lemma 8 is a statement about a single graph, but we show it can be invoked to prove the following.
Theorem 9.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model with parameters $C$ and $s$. Let $v\in V$ and let $U$ be a vertex cut of $\{1,\cdots,m\}$ such that $|U|\leq\lfloor m/2\rfloor$. Then,
($\mathbb{P}\left(\left\{c_{v}(U)=0\right\}\cap\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}\right)=o(1/n).$ (2)
It suffices therefore to analyze the probability that the graph ~$\widetilde{\mathcal{H}}(v)$ has a cut $U$ such that its cost $\widetilde{c}_{v}(U)$ is too small. To that end, we show that the bottleneck arises from vertex cuts of small size. Formally,
Theorem 10.
Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model. Let $v\in V$ and let $U_{\ell}$ denote the set $\{1,\cdots,\ell\}$ for $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$. For any vertex cut $U$ of $\{1,\cdots,m\}$, let $\widetilde{c}_{v}(U)$ denote its cost in the graph ~$\widetilde{\mathcal{H}}(v)$. The following stochastic ordering holds:
$\widetilde{c}_{v}(U_{1})\preceq\widetilde{c}_{v}(U_{2})\preceq\cdots\preceq\widetilde{c}_{v}(U_{\lfloor m/2\rfloor}).$
Theorems 9 and 10 imply that the tightest bottleneck to the connectivity of $\mathcal{H}(v)$ is the event that $\widetilde{c}_{v}(U_{1})$ is below the threshold $r:=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$, i.e. the sum of degrees of $v$ over the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$ is less than $r$. This event occurs only if the degree of $v$ is less than $r$ in each of the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$. However, under the condition $Cs(1-(1-s)^{m-1})>1$, it turns out that this event occurs with probability $o(1/n)$.
Theorem 11.
Let $G_{1},\cdots,G_{m}$ be obtained from the subsampling model with parameters $C$ and $s$. Let $r=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$. Let $v\in[n]$ and suppose that $Cs(1-(1-s)^{m-1})>1$. Then,
$\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq r\right)\leq\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\right\}\right)=o\left(1/n\right).$
4.3 Piecing it all together: Proof of Theorem 3
Proof of Theorem 3. Let $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ denote the output of Algorithm 2 with $k\geq 13$. Let $E_{1}$ (resp. $E_{2}$) denote the event that Algorithm 1 (resp. Algorithm 2) fails to match all $m$ graphs exactly, i.e.
$E_{1}=\left\{\widehat{\pi}_{12}^{\text{ML}}\neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}^{\text{ML}}\neq\pi^{*}_{1m}\right\},\quad\quad E_{2}=\left\{\widehat{\pi}_{12}\ \neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}\neq\pi^{*}_{1m}\right\}.$
First, we show that the output of Algorithm 2 is correct with probability $1-o(1)$ whenever $Cs(1-(1-s)^{m-1})>1$. If the event $E_{2}$ occurs, then either step 1 failed, i.e. there is a $k$-core matching $\widehat{\mu}_{ij}$ that is incorrect, or step 2 failed, i.e. at least one of the graphs $\mathcal{H}(v)$ is not connected. Therefore,
$\mathbb{P}\left(E_{2}\right)$ $\leq\mathbb{P}\left(\bigcup_{i,j}\ \bigcup_{v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})}\!\!\!\!\left\{\widehat{\mu}_{ij}\neq\pi^{*}_{ij}\right\}\!\right)\!+\mathbb{P}\left(\bigcup_{v\in V}\left\{\mathcal{H}(v)\text{ is not connected}\right\}\!\right)\!\leq o(1)+\!\sum_{v\in V}q_{v},$
where the last step uses Proposition 5, and $q_{v}$ denotes the probability that the transitivity graph $\mathcal{H}(v)$ is not connected. For each $\ell$ in the set $\{1,\cdots,\lfloor m/2\rfloor\}$, let $U_{\ell}$ denote the set $\left\{1,\cdots,\ell\right\}$. Then,
$q_{v}\!$ $=\mathbb{P}\left(\bigcup_{\ell=1}^{\lfloor m/2\rfloor}\left\{\exists\ U\subset\{1,\cdots,m\}:|U|=\ell\text{ and }c_{v}(U)=0\right\}\right)$
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\cdot\mathbb{P}\left(c_{v}(U_{\ell})=0\right)$
$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!\mathbb{P}\left(\left\{c_{v}(U_{\ell})=0\right\}\cap\bigg{\{}\widetilde{c}_{v}(U_{\ell})>\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}\right)\right]$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$
$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\left[o\left(\frac{1}{n}\right)+o\left(\frac{1}{n}\right)\right]=o\left(\frac{1}{n}\right).$
Here, (a) uses Theorem 9, and (b) uses the fact that for any $\ell\geq 2$, the random variable $\widetilde{c}_{v}(U_{\ell})$ stochastically dominates $\widetilde{c}_{v}(U_{1})$ (Theorem 10). Finally, (c) uses Theorem 11 and the fact that $Cs(1-(1-s)^{m-1})>1$. Therefore, a union bound over all the nodes yields
$\mathbb{P}\left(E_{2}\right)\leq o(1)+\sum_{v\in V}q_{v}\leq o(1)+n\times o(1/n)=o(1).$
Finally, by optimality of the MLE, it follows that
$\mathbb{P}\left(E_{1}\right)\leq\mathbb{P}\left(E_{2}\right)=o(1),$
whenever $Cs(1-(1-s)^{m-1})>1$. This concludes the proof. ∎
5 Discussion and Future Work
In this work, we introduced and analyzed matching through transitive closure - an approach that combines information from multiple graphs to recover the underlying correspondence between them. Despite its simplicity, it turns out that matching through transitive closure is an optimal way to combine information in the setting where the graphs are pairwise matched using the $k$-core estimator. A limitation of our algorithms is the runtime: Algorithm 2 does not run in polynomial time because it uses the $k$-core estimator for pairwise matching, which involves searching over the space of permutations. Even so, it is useful to establish the fundamental limits of exact recovery, and serve as a benchmark to compare the performance of any other algorithm.
The transitive closure subroutine (Step 2) itself is efficient because it runs in polynomial time $O(mn)$. Therefore, a natural next step is to modify Step 1 in our algorithm so that the pairwise matchings are done by an efficient algorithm. However, it is not clear if transitive closure is optimal for combining information from the pairwise matchings in this setting. For example, there is a possibility that the pairwise matchings resulting from the efficient algorithm are heavily correlated, and transitive closure is unable to boost them. In Figure 3, we show experimentally that this is not the case for two algorithms of interest: GRAMPA [FMWX22] and Degree Profiles [DMWX21].
  1. 1.
    GRAMPA is a spectral algorithm that uses the entire spectrum of the adjacency matrices to match the two graphs. The code is available in [FMWX20].
  2. 2.
    Degree Profiles associates with each node a signature derived from the degrees of its neighbors, and matches nodes by signature proximity. The code is available in [DMWX20].
Evidently, both algorithms benefit substantially from using transitive closure to boost the number of matched nodes. This suggests that transitive closure can be a practical algorithm to boost matchings between networks by using other networks as side-information. Unfortunately, both GRAMPA and Degree Profiles require the graphs to be close to isomorphic in order to perform well, and so they do not perform well when the model parameters are close to the information theoretic threshold for exact recovery. Subsequently, they cannot be used to answer the question in 1.
Our work presents several directions for future research.
Figure 3: Matching through transitive closure
  • Polynomial-time algorithms. Using a polynomial-time estimator in place of the $k$-core estimator in Step 1 of Algorithm 2 yields a polynomial-time algorithm to match $m$ graphs. It is critical that the estimator in question is able to identify for itself the nodes that it has matched correctly - this precision is present in the $k$-core estimator and enables the transitive closure subroutine to work correctly. Can the performance guarantees of the $k$-core estimator be realized through polynomial time algorithms that meet this constraint?
  • Beyond Erdős-Rényi graphs. The study of matching two ER graphs provided tools and techniques that extended to the analysis of more realistic models. For instance, the $k$-core estimator itself played a crucial role in establishing limits to matching two correlated stochastic block models [GRS22] and two inhomogeneous random graphs [RS23]. Can the techniques developed in the present work be used to identify the information theoretic limits to exact recovery in these models in the general setting of $m$ graphs?
  • Boosting for partial recovery. This work focused on exact recovery, where the objective is to match all nodes across all graphs. It would be interesting to consider a regime where any instance of pairwise matching recovers at best a small fraction of nodes. Is it possible to quantify the extent to which transitive closure boosts the number of matched nodes?
  • Robustness. Finally, how sensitive to perturbation is the transitive closure algorithm? Is it possible to quantify the extent to which an adversary may perturb edges in some of the graphs without losing the performance guarantees of the matching algorithm? Algorithms that perform well on models such as ER graphs and are further generally robust are expected to also work well with real-world networks.
Acknowledgments and Disclosure of Funding
This work was supported by NSF under Grant CCF 19-00636.
Appendix A Maximum Likelihood Estimator
Recall the form of the maximum likelihood estimator as claimed in Theorem 4. See 4
Proof.
Notice that
$\mathbb{P}\left(G_{1},\cdots,G_{m}|\pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)$ $=\prod_{e\in\binom{[n]}{2}}\mathbb{P}\left(G_{1}(e),G_{2}(\pi^{*}_{12}(e))\cdots,G_{m}(\pi^{*}_{1m}(e))\mid\pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)$
$=\prod_{e\in\binom{[n]}{2}}\mathbb{P}\left(G_{1}(e),G_{2}^{\prime}(e)\cdots,G_{m}^{\prime}(e)\right)$ (3)
where for a node pair $e=\{u,v\}$, the shorthand $\pi(e)$ denotes $\{\pi(u),\pi(v)\}$. The edge status of any node pair $e$ in the graph tuple $(G_{1},G_{2}^{\prime},\cdots,G_{m}^{\prime})$ can be any of the $2^{m}$ bit strings of length $m$, but the corresponding probability in (3) depends only on the number of ones and zeros in the bit string. For $i\in[m]$, let $\alpha_{i}$ denote the number of node pairs $e$ whose corresponding tuple $(G_{1}(e),G_{2}^{\prime}(e),\cdots,G_{m}^{\prime}(e))$ has exactly $i$ 1’s:
[…]
Two key observations are in order. First, it follows by definition that $\alpha_{0}+\alpha_{1}+\cdots+\alpha_{m}=\binom{n}{2}$. Second, by definition of $\alpha_{i}$, it follows that
$\sum_{i=0}^{m}i\alpha_{i}=\sum_{e\in\binom{[n]}{2}}\sum_{j=1}^{m}G_{j}(e)=\sum_{e\in\binom{[n]}{2}}\sum_{j=1}^{m}G^{\prime}_{j}(e)$ (4)
is constant, independent of $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$. It follows then that
[…]
$=\left(1-p+p(1-s)^{m}\right)^{\alpha_{0}}\times p^{\sum_{i=1}^{m}\alpha_{i}}\times\prod_{i=1}^{m}\left(s^{i}(1-s)^{m-i}\right)^{\alpha_{i}}$
$=\left(1-p+p(1-s)^{m}\right)^{\alpha_{0}}\times p^{\binom{n}{2}-\alpha_{0}}\times\left(\frac{s}{1-s}\right)^{\sum_{i=1}^{m}i\alpha_{i}}\times(1-s)^{m\sum_{i=1}^{m}\alpha_{i}}$
$=\left(\frac{1-p+p(1-s)^{m}}{p(1-s)^{m}}\right)^{\alpha_{0}}\times\left(p(1-s)^{m}\right)^{\binom{n}{2}}\times\left(\frac{s}{1-s}\right)^{\sum_{i=1}^{m}i\alpha_{i}}$
$\propto\left(1+\frac{1-p}{p(1-s)^{m}}\right)^{\alpha_{0}},$
where the last step uses (4). Finally, since $\frac{1-p}{p(1-s)^{m}}>0$, it follows that the log-likelihood satisfies
$\log\left(\mathbb{P}\left(G_{1},\cdots,G_{m}\ |\ \pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)\right)\propto\text{const.}+\alpha_{0},$
i.e. maximizing the likelihood corresponds to selecting $\pi_{12},\cdots,\pi_{1m}$ to maximize $\alpha_{0}$ the number of node pairs $e$ for which $G_{1}(e)=G_{2}(\pi_{12}(e))=\cdots=G_{m}(\pi_{1m}(e))=0$. This is equivalent to minimizing the number of edges in the union graph $G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}$, as desired. ∎
Remark 12.
In the case of two graphs, minimizing the number of edges in the union graph $G_{1}\vee_{\pi}G_{2}$ is equivalent to maximizing the number of edges in the intersection graph $G_{1}\wedge_{\pi}G_{2}$. This is consistent with existing literature on two graphs [CK16, WXY22].
Appendix B Concentration Inequalities for Binomial Random Variables
The following bounds for the binomial distribution are used frequently in the analysis.
Lemma 13.
Let $X\sim\mathsf{Bin}(n,p)$. Then,
  1. 1.
    For any $\delta>0$,
    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{np}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)np}.$ (5)
  2. 2.
    For any $\delta>5$,
    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq 2^{-(1+\delta)np}.$ (6)
  3. 3.
    For any $\delta\in(0,1)$,
    $\mathbb{P}\left(X\leq(1-\delta)np\right)\leq\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{np}.$ (7)
Proof.
All proofs follow from the Chernoff bound and can be found, or easily derived, from Theorems 4.4 and 4.5 of [MU17]. ∎
Appendix C Proof of Lemma 8
We restate Lemma 8 for convenience. See 8
Before presenting the proof, we present the intuition behind it. The events $\{v\notin\mathsf{core}_{k}(G)\}$ and $\left\{\delta_{G}(v)\geq k+1/\alpha\right\}$ are highly negatively correlated. However, consider the subgraph $(G-v)$ of $G$ induced on the vertex set $V-\{v\}$, and note that the $k$-core of this subgraph does not depend on the degree of $v$. Furthermore, if $v\notin\mathsf{core}_{k}(G)$, then it must be that $v$ has fewer than $k$ neighbors in $\mathsf{core}_{k}(G-v)$. Intuitively, this event has low probability if $\mathsf{core}_{k}(G-v)$ is sufficiently large.
Notice that $(G-v)\sim\mathsf{ER}(n-1,\alpha\log(n)/n)$, and so standard results about the size of the $k$-core of Erdős-Rényi graphs apply. However, we require the error probability that the $k$-core of $G-v$ is too small to be $o(1/n)$ this is crucial since we will later use a union bound over all the nodes $v$. Unfortunately, standard results such as [Łuc91] can only be invoked directly to show that the corresponding probability is $o(1)$, which is insufficient for our purpose. Later in this section, we refine the analysis in [Łuc91] to obtain the desired convergence rate. The refinement culminates in the following.
Lemma 14.
Let $\alpha>0$ and $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$. Let $v$ be a node of $G$. The size of the $k$-core of $G-v$ satisfies
$\mathbb{P}\left(\left|\mathsf{core}_{k}(G-v)\right|<n-3n^{1-\alpha}\right)=o(1/n).$
The proof of Lemma 14 is deferred to Section C.1. It remains to study the error event that $v$ has too few neighbors in $\mathsf{core}_{k}(G-v)$. To count the number of neighbors of $v$ in $\mathsf{core}_{k}(G-v)$, we exploit the independence of $\mathsf{core}_{k}(G-v)$ and $v$ as follows: each neighbor of $v$ is considered a success if it belongs to $\mathsf{core}_{k}(G-v)$ and a failure otherwise. Counting the number of successes is equivalent to sampling with replacement $\delta_{G}(v)$ elements, each of which is independently a success with probability $|\mathsf{core}_{k}(G-v)|/\left(n-1\right)$. The number of successes then follows precisely a hypergeometric distribution. This intuition is made rigorous in the proof below.
Let us recall some facts about the hypergeometric distribution because it plays an important role in the proof. Denote by $\mathsf{HypGeom}(N,K,n)$ a random variable that counts the number of successes in a sample of $n$ elements drawn without replacement from a population of $N$ individuals, of which $K$ elements are considered successes. Note that if this sampling were done with replacement, then the number of successes would follow a $\mathsf{Bin}\left(n,K/N\right)$ distribution. A result of Hoeffding [Hoe94] establishes that the $\mathsf{HypGeom}(N,K,n)$ distribution is convex-order dominated by the $\mathsf{Bin}\left(n,K/N\right)$ distribution, i.e.
$\mathbb{E}\left[f(\mathsf{HypGeom}(N,K,n))\right]\leq\mathbb{E}\left[f(\mathsf{Bin}(n,K/N))\right]\ \text{ for all convex functions $f$}.$
In particular, the function $f(x)=e^{tx}$ is convex for any value of $t$, and so Chernoff bounds that hold for the binomial distribution also hold for the corresponding hypergeometric distribution. This yields the following proposition.
Proposition 15.
Let $X\sim\mathsf{HypGeom}(N,K,n)$. It follows for any $\delta>0$ that
$\mathbb{P}\left(X>(1+\delta)\times\frac{nK}{N}\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{nK/N}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)nK/N}.$
Our final remark about the hypergeometric distribution is a symmetry property. By interchanging the success and failure states, it follows that
$\mathbb{P}\left(\mathsf{HypGeom}(N,K,n)=k\right)=\mathbb{P}\left(\mathsf{HypGeom}(N,N-K,n)=n-k\right).$
The above intuition for the proof of Lemma 8 is formalized below.
Proof of Lemma 8. Let $V$ denote the vertex set of $G$, and let $G-v$ denote the induced subgraph of $G$ on the vertex set $V-\{v\}$. For any set $A\subseteq V$, let $N_{v}(A)$ denote the set of neighbors of $v$ in the set $A$, i.e.
$N_{v}(A):=\left\{u\in A:\{u,v\}\in E(G)\right\}.$
Since $\mathsf{core}_{k}(G-v)\subseteq\mathsf{core}_{k}(G)$, it is true that
$\left\{v\notin\mathsf{core}_{k}(G)\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G))|\leq k-1\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G-v))|\leq k-1\right\}.$
It follows that
($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)\leq p_{1}+p_{2},$
where
$p_{1}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right\}\right)$
$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
It suffices to show that both $p_{1}$ and $p_{2}$ are $o(1/n)$. The term $p_{1}$ deals with the probability that the $k$-core of $G-v$ is too small. In fact, by Lemma 14, it follows directly that
$p_{1}$ $\leq\mathbb{P}\left(|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right)=o(1/n),$
Next, the probability $p_{2}$ is analyzed. Enumerate arbitrarily but independently the elements of sets $N_{v}(V)$ and $\mathsf{core}_{k}(G-v)$, so that
$N_{v}(V)=\left\{v_{1},\cdots,v_{\delta_{G}(v)}\right\},\quad\mathsf{core}_{k}(G-v)=\left\{a_{1},\cdots,a_{|\mathsf{core}_{k}(G-v)|}\right\}.$
Given that $N_{v}(V)$ has more than $k+1/\alpha$ nodes and $\mathsf{core}_{k}(G-v)$ has more than $n-3n^{1-\alpha}$ nodes, it is true that
$N_{v}(\mathsf{core}_{k}(G-v))$ $=N_{v}(V)\cap\mathsf{core}_{k}(G-v)$
$\supseteq\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}\cap\left\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\right\}=:\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v)).$
In words, $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ counts among the first $\lceil k+1/\alpha\rceil$ neighbors of $v$ those nodes that are also in the first $\lceil n-3n^{1-\alpha}\rceil$ nodes of $\mathsf{core}_{k}(G-v)$. Therefore,
$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq\ k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
($\leq\mathbb{P}\left(\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$
($\leq\mathbb{P}\left(\{\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$ (8)
Note that $\mathsf{core}_{k}(G-v)$ is entirely determined by the graph $G-v$, i.e. it is independent of the neighbors of $v$. Consequently, the two sets $\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}$ and $\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\}$ are selected independent of each other. Equivalently, given that $|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}$ and $\delta_{G}(v)\geq k+1/\alpha$, the size of the intersection set $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ follows a hypergeometric distribution with parameters $(n-1,\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)$. Therefore,
[…]
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathbb{P}\left(\mathsf{HypGeom}(n-1,n-1-\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)\geq\lceil k+1/\alpha\rceil-(k-1)\right)$
$=\mathbb{P}\left(\mathsf{HypGeom}(n-1,\lfloor 3n^{1-\alpha}\rfloor-1,\lceil k+1/\alpha\rceil)\geq 1+1/\alpha\right)$ (9)
where (a) uses the symmetry of the hypergeometric distribution. Using Proposition 15 and the fact that $n-1\geq n/2$ for any $n\geq 1$ yields
[…]
$\leq\left(\frac{6e\lceil k+1/\alpha\rceil}{1+1/\alpha}\right)^{1+1/\alpha}\times n^{-1-\alpha}$
$=o(1/n),$
whenever $\alpha>0$. ∎
C.1 Proof of Lemma 14
A key ingredient towards proving Lemma 14 is a useful result about the number of low-degree vertices in an Erdős-Rényi graph, presented next.
Proposition 16.
Let $\alpha>0$ and $G\sim\mathsf{ER}\left(n-1,\alpha\log(n)/n\right)$. Let $r$ be a positive integer and let $Z_{r}$ denote the set of vertices in $G$ with degree no more than $r$, i.e.
$Z_{r}=\left\{v\in V(G):\delta_{G}(v)\leq r\right\}.$
For any $\delta$ such that $\delta>1-\alpha$, it is true that
$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)=o(1/n).$
Proof.
Notice that
$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)$ $=\mathbb{P}\left(\exists S^{\prime}\subseteq V:\left\{|S^{\prime}|\geq n^{\delta}\right\}\cap\left\{\max_{i\in S^{\prime}}\delta_{G}(i)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\max_{i\in S}\delta_{G}(i)\leq r\right\}\right)$
$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{i\in S}\delta_{G}(i)\leq r|S|\right\}\right).$ (10)
If $|S|=n^{\delta}$, then the sum of degrees of vertices in $S$ is the total number of edges with exactly end point in $S$, plus twice the number of edges with both end points in $S$. There are exactly $\binom{|S|}{2}+|S|\left(n-1-|S|\right)\leq n^{1+\delta}$ such vertex pairs, and each of them independently has an edge with probability $\alpha\log(n)/n$. Therefore, a union bound over all possible choices of $S$ yields
[…]
$\leq\binom{n-1}{n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(n^{1+\delta},\alpha\log(n)/n\right)\leq rn^{\delta}\right)$
$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\left(\frac{ne}{n^{\delta}}\right)^{n^{\delta}}\times\left(\frac{\exp\left(r/(\alpha\log n)-1\right)}{(r/(\alpha\log n))^{r/(\alpha\log n)}}\right)^{n^{\delta}\alpha\log(n)}$
$=o(1/n),$
whenever $\delta>1-\alpha$ as desired. Note that (a) uses the Binomial concentration inequality (7) and the fact that $\binom{n-1}{k}\leq\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$. ∎
Our objective is to show that the $k$-core of $G-v$ is sufficiently large with probability $1-o(1/n)$. To that end, consider Algorithm 3 to identify a subset of the $k$-core, originally proposed by Łuczak [Łuc91].
require :  Graph $G$, Set $U\subseteq V(G)$.
1 $U_{0}\leftarrow U$
2 for $i=0,1,2,3,\cdots$ do
3       if  there exists $u\in V\setminus U_{i}$ such that $u$ has $3$ or more neighbors in $U_{i}$ then
4             $U_{i+1}\leftarrow U_{i}\cup\{u\}$
5
6       else
7             return $U_{i}$
8
9       end if
10
11 end for
Algorithm 3 Łuczak expansion
Note that the for loop eventually terminates - the set $V\setminus U_{i}$ is empty, for example when $i=n$ for any input set $U$. The key is to realize that the for loop terminates much faster when the input $U=Z_{k+1}$ , i.e the set of vertices of the input graph $G$ whose degree is $k+1$ or less. Furthermore, the complement of the set output by the algorithm is contained in the $k$-core. Formally,
Lemma 17.
Let $U_{f}$ be the output of Algorithm 3 with input graph $G-v$ and set $U=Z_{k+1}$. Then,
  1. (a)
    $U_{f}^{c}\subseteq\mathsf{core}_{k}(G-v)$.
  2. (b)
    For any $\delta>1-\alpha$,
    $\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)=o(1/n).$
Proof.
(a) The proof is by construction: Since $U_{f}$ is obtained by adding exactly $f$ nodes to $U_{0}$, it follows that $U_{f}^{c}\subseteq U_{0}^{c}=Z_{k+1}^{c}$, so each node in $U_{f}^{c}$ has degree $k+2$ or more in $G-v$. Further, each node in $U_{f}^{c}$ has at most $2$ neighbors in $U_{f}$, else the for loop would not have terminated. Thus, the subgraph of $G-v$ induced on the set $U_{f}^{c}$ has minimum degree at least $k$, and the result follows.
(b) If $|U_{f}|>3n^{\delta}$, then either $|U_{0}|>3n^{\delta}$ or there is some $M$ in $\{0,1,\cdots,f\}$ for which $|U_{M}|=3n^{\delta}$. Therefore,
$\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)$ $\leq\mathbb{P}\left(|U_{0}|>3n^{\delta}\right)+\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)$
$=o(1/n)+\underbrace{\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)}_{(\star)},$
by Proposition 16. Note that each iteration $i=0,1,\cdots,M-1$ of the for loop adds exactly $1$ vertex and at least $3$ edges to the subgraph of $G-v$ induced on $U_{M}$. Therefore, the induced subgraph $G|_{U_{M}}$ has $3n^{\delta}$ vertices and at least $3\left(|U_{M}|-|U_{0}|\right)$ edges. Thus,
(($\star$)( ⋆ ) ( $\leq\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{ and }|F|\geq 3\left(3n^{\delta}-|U_{0}|\right)\right)$
$\leq\mathbb{P}\left(|U_{0}|>n^{\delta}\right)+\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{and }|F|\geq 6n^{\delta}\right)$
$\leq o(1/n)+\underbrace{\binom{n}{3n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(\binom{3n^{\delta}}{2},\frac{\alpha\log(n)}{n}\right)>6n^{\delta}\right)}_{(\star\star)},$
where the last step uses Proposition 16 and a union bound over all possible choices of $W$. Finally, using the relation $\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$ and the concentration inequality (5) from Lemma 13 yields
[…]
$\leq(n^{1-\delta})^{3n^{\delta}}\times\left(\frac{3\alpha e\log n}{4n^{1-\delta}}\right)^{6n^{\delta}}$
$=\left(\frac{3\alpha e\log n}{4n^{(1-\delta)/2}}\right)^{6n^{\delta}}$
$=o(1/n),$
whenever $10<\delta<$ 1. The result follows. ∎
Finally, notice that Lemma 17 directly implies Lemma 14.
Appendix D Proof of Theorem 9
See 9
Proof.
For any vertex cut $U$,
$\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}$ $\stackrel{{\scriptstyle\text{(a)}}}{{\subseteq}}\left\{\widetilde{c}_{v}(U)>|U|\left(m-|U|\right)\left(k+\frac{1}{Cs^{2}}\right)\right\}$
=$=\bigg{\{}\sum_{i\in U}\sum_{j\in U^{c}}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>|U|(m-|U|)\bigg{(}k+\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}$
$\subseteq\bigcup_{i\in U}\bigcup_{j\in U^{c}}\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\},$
where (a) uses the fact that the maximum of a set of a numbers is greater than or equal to the average. On the other hand
$\left\{c_{v}(U)=0\right\}=\bigcap_{i\in U}\bigcap_{j\in U^{c}}\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}.$
Let $p_{1}$ denote the probability in the LHS of (2). It follows from the union bound that
$p_{1}\leq\sum_{i\in U}\sum_{j\in U^{c}}\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}\cap\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\}\right)=o(1/n),$
since for any choice of $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}\left(n,Cs^{2}\log(n)/n\right)$. ∎
Appendix E On Stochastic Dominance: Proof of Theorem 10
The objective of this section is to build up to a proof of Theorem 10. We start by making a simple observation about products of Binomial random variables.
Lemma 18.
Let $X_{1},\cdots,X_{m}\sim\mathsf{Bern}(s)$ be i.i.d. random variables, and let $B=X_{1}+\cdots+X_{m}$ denote their sum. For each $\ell$ in $\{1,2,\cdots,\lfloor m/2\rfloor\}$, define
$T_{\ell}=\left(X_{1}+\cdots+X_{\ell}\right)\left(X_{\ell+1}+\cdots+X_{m}\right).$
For any $\ell_{1},\ell_{2}\in\{1,2,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$, and for any $t\in\mathbb{R}$ and any $b\in\{0,1,\cdots,m\}$,
$\mathbb{P}\left(T_{{\ell}_{1}}>t\mid B=b\right)\leq\mathbb{P}\left(T_{\ell_{2}}>t\mid B=b\right).$ (11)
Proof of Lemma 18. Consider overlapping but exhaustive cases:
Case 1: $t<0$. Since $T_{\ell}\geq 0$ almost surely for all $\ell$, the inequality (11) holds.
Case 2: $t\geq b-1$. Note that conditioned on $B=b$, it follows that $T_{1}\in\{0,b-1\}$. Therefore, the left hand side of (11) equals zero, and the inequality holds.
Case 3: $b=0$ or $b=1$. In this case, $T_{\ell}$ is identically zero for all $\ell$, so (11) holds.
Case 4: $b\geq 2$ and $10\leq t<b-$ 1. For any $\ell\in\{1,2,\cdots,\lfloor m/2\rfloor\}$,
$\mathbb{P}\left(T_{\ell}>t\mid B=b\right)$ =($=\frac{\mathbb{P}\left(\left\{\left(X_{1}+\cdots+X_{\ell}\right)\left(X_{\ell+1}+\cdots+X_{m}\right)>t\right\}\cap\left\{X_{1}+\cdots+X_{m}=b\right\}\right)}{\mathbb{P}\left(X_{1}+\cdots+X_{m}=b\right)}$
$=\frac{\sum_{i:i(b-i)>t}\mathbb{P}\left(\left\{X_{1}+\cdots+X_{\ell}=i\right\}\cap\left\{X_{\ell+1}+\cdots+X_{m}=b-i\right\}\right)}{\mathbb{P}\left(X_{1}+\cdots+X_{m}=b\right)}$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\frac{\sum_{i=1}^{b-1}\mathbb{P}\left(X_{1}+\cdots+X_{\ell}=i\right)\mathbb{P}\left(X_{\ell+1}+\cdots+X_{m}=b-i\right)}{\mathbb{P}\left(X_{1}+\cdots+X_{m}=b\right)}$
=$\stackrel{{\scriptstyle\text{(b)}}}{{=}}\frac{\sum_{i=1}^{b-1}\binom{\ell}{i}\binom{m-\ell}{b-i}}{\binom{m}{b}}$
$=\frac{\sum_{i=0}^{b}\binom{\ell}{i}\binom{m-\ell}{b-i}-\binom{m-\ell}{b}-\binom{\ell}{b}}{\binom{m}{b}}$
$=\frac{\binom{m}{b}-\binom{m-\ell}{b}-\binom{\ell}{b}}{\binom{m}{b}},$ (12)
where (a) used the fact that for any $t$ such that $10\leq t<b-$ 1, it is true that
$\{i:i(b-i)>t\}=\left\{1,2,\cdots,b-1\right\}.$
Here, the notation for binomial coefficients in (b) involves setting $\binom{n}{k}=0$ whenever $k<0$ or $k>n$. Let $f_{m,b}(\ell)$ denote the numerator of (12), i.e.
$f_{m,b}(\ell):=\binom{m}{b}-\binom{m-\ell}{b}-\binom{\ell}{b}$
It suffices to show that $f_{m,b}(\ell)-f_{m,b}(\ell-1)\geq 0$ for all $\ell\in\left\{2,\cdots,\lfloor m/2\rfloor\right\}$. Indeed,
$f_{m,b}(\ell)-f_{m,b}(\ell-1)$ $=\binom{m-\ell+1}{b}-\binom{m-\ell}{b}-\left(\binom{\ell}{b}-\binom{\ell-1}{b}\right)$
=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}\binom{m-\ell}{b-1}-\binom{\ell-1}{b-1}\geq 0,$
whenever $m-\ell\geq\ell-1$, i.e. $\ell\leq\lfloor m/2\rfloor$. Here, (c) uses the identity $\binom{n}{k}=\binom{n-1}{k-1}+\binom{n-1}{k}$, and the fact that $\binom{n_{1}}{k}\geq\binom{n_{2}}{k}$ whenever $n_{1}\geq n_{2}$. This concludes the proof. ∎
Corollary 19.
Let $F$ be a collection of edges in the parent graph $G$. For any edge $e_{r}\in F$, let $X^{r}_{i}$ denote the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$. For each $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$, define
$T_{\ell}^{r}=(X_{1}^{r}+\cdots+X_{\ell}^{r})(X_{\ell+1}^{r}+\cdots+X_{m}^{r}).$
Then, for any $\ell_{1},\ell_{2}\in\{1,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$, the following stochastic ordering holds
$\sum_{r=1}^{|F|}T^{r}_{\ell_{1}}\preceq\sum_{r=1}^{|F|}T^{r}_{\ell_{2}}.$
Proof.
It suffices to show that $T^{r}_{\ell_{1}}\preceq T^{r}_{\ell_{2}}$ for each $r$, since the edges are independent. Indeed, we have for any $t$ that
$\mathbb{P}\left(T^{r}_{\ell_{1}}>t\right)=\sum_{b=0}^{m}\mathbb{P}\left(B=b\right)\mathbb{P}\left(T_{\ell_{1}}^{r}>t|B=b\right)\leq\sum_{b=0}^{m}\mathbb{P}\left(B=b\right)\mathbb{P}\left(T_{\ell_{2}}^{r}>t|B=b\right)=\mathbb{P}\left(T^{r}_{\ell_{2}}>t\right),$
which concludes the proof. ∎
With this, we are ready to prove Theorem 10. The theorem is restated for convenience.
See 10
Proof.
Let $\ell_{1},\ell_{2}\in\{1,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$. Let $t\in\mathbb{R}$. Consider the parent graph $G$ and label the set of incident edges on $v$ as $\{e_{1},\cdots,e_{\delta_{G}(v)}\}$. Denote by $X_{i}^{r}$ the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$. It follows that
$\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell_{2}})>t\right)$ $=\mathbb{P}\left(\sum_{i=1}^{\ell_{2}}\sum_{j=\ell_{2}+1}^{m}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)\geq t\right)$
$=\mathbb{P}\left(\sum_{i=1}^{\ell_{2}}\sum_{j=\ell_{2}+1}^{m}\sum_{r=1}^{\delta_{G}(v)}X^{r}_{i}X^{r}_{j}>t\right)$
$=\sum_{d=0}^{n}\mathbb{P}\left(\delta_{G}(v)=d\right)\mathbb{P}\left(\sum_{r=1}^{d}\left((X_{1}^{r}+\cdots+X_{\ell_{2}}^{r})(X_{\ell_{2}+1}^{r}+\cdots+X_{m}^{r})\right)>t\right)$
$\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\sum_{d=0}^{n}\mathbb{P}\left(\delta_{G}(v)=d\right)\mathbb{P}\left(\sum_{r=1}^{d}\left((X_{1}^{r}+\cdots+X_{\ell_{1}}^{r})(X_{\ell_{1}+1}^{r}+\cdots+X_{m}^{r})\right)>t\right)$
$=\mathbb{P}\left(\sum_{i=1}^{\ell_{1}}\sum_{j=\ell_{1}+1}^{m}\sum_{r=1}^{\delta_{G}(v)}X^{r}_{i}X^{r}_{j}>t\right)$
$=\mathbb{P}\left(\sum_{i=1}^{\ell_{1}}\sum_{j=\ell_{1}+1}^{m}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)\geq t\right)$
$=\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell_{1}})>t\right),$
as desired. Here, (a) uses Corollary 19. ∎
Appendix F On Low Degree Nodes: Proof of Theorem 11
See 11
Proof.
Consider fixed integers $r_{1},\cdots,r_{m}$ such that $0\leq r_{2},\cdots,r_{m}\leq r$. Since $r$ is constant, by a union bound argument it suffices to show
()=:($(\star)=:\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\right\}\right)=o\left(1/n\right).$
Proceed by conditioning on the degree of $v$ in $G_{1}$, which follows a $\mathsf{Bin}(n,ps)$ distribution. Since the degrees of $v$ in the intersection graphs $\{G_{1}\wedge G_{i}^{\prime}:i=2,\cdots,m\}$ are conditionally independent given the degree of $v$ in $G_{1}$, we have
(($\star$)( ⋆ ) $=\mathbb{E}_{D}\left[\mathbb{P}\Big{(}\bigcap_{i=2}^{m}\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\Big{)}\right]$
$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\right)\right]$
$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\binom{D}{r_{i}}s^{r_{i}}(1-s)^{D-r_{i}}\right].$ (13)
Using the fact that $\binom{D}{r_{i}}\leq\left(\frac{De}{r_{i}}\right)^{r_{i}}$, it follows that
[…]
Expanding out the expectation yields
[…]
Proceed by splitting the summation at $(\log n)^{2}$. The first part can be bounded as
$\sum_{d=0}^{\left(\log n\right)^{2}}L_{d}$ $\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\sum_{d=0}^{(\log n)^{2}}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
$\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\sum_{d=0}^{n}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
$=\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\mathbb{E}_{D}\left[(1-s)^{(m-1)D}\right]$
=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\left(1-\frac{Cs\left(1-(1-s)^{m-1}\right)\log n}{n}\right)^{n}$
$=o(1/n),$
whenever $Cs(1-(1-s))^{m-1}>1$. Here, (a) is obtained by evaluating the probability generating function of the $\mathsf{Bin}(n,ps)$ random variable at $(1-s)^{m-1}$ and setting $p=C\log(n)/n$.
The other part of the sum can now be bounded as follows.
$\sum_{d=\left(\log n\right)^{2}}^{n}L_{d}$ $\leq\left[\max_{d\colon(\log n)^{2}\leq d\leq n}d^{\sum_{i=2}^{m}r_{i}}(1-s)^{md}\right]\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)\geq(\log n)^{2}\right)$
$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\left[\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}(1-s)^{m\left(\log n\right)^{2}}\right]\times 2^{-\left(\log n\right)^{2}}$
$=(\log n)^{2\sum_{i=2}^{m}r_{i}}\left(\frac{(1-s)^{m}}{2}\right)^{(\log n)^{2}}$
$=o\left(1/n\right).$
whenever $C>0$. Here, (b) is true because the function $d\mapsto d^{\sum r_{i}}(a-s)^{md}$ is decreasing on the interval $[(\log n)^{2},n]$ for all sufficiently large $n$. Finally, the concentration inequality for the Binomial distribution holds by (6) in Lemma 13. The inequality applies since $p=C\log(n)/n$ and since $(\log n)^{2}>6Cs\log(n)$ for all $n$ sufficiently large. This concludes the proof. ∎

### v2

@@
-Exact Random Graph Matching with Multiple Graphs
+Aligning Multiple Inhomogeneous Random Graphs: Fundamental Limits of Exact Recovery
 Abstract
-This work studies fundamental limits for recovering the underlying correspondence among multiple correlated random graphs.
-We identify a necessary condition for any algorithm to correctly match all nodes across all graphs, and propose two algorithms for which the same condition is also sufficient.
-The first algorithm employs global information to simultaneously match all the graphs, whereas the second algorithm first partially matches the graphs pairwise and then combines the partial matchings by transitivity.
-Both algorithms work down to the information theoretic threshold.
-Our analysis reveals a scenario where exact matching between two graphs alone is impossible, but leveraging more than two graphs allows exact matching among all the graphs.
-Along the way, we derive independent results about the $k$-core of Erdős-Rényi graphs.
+This work studies fundamental limits for recovering the underlying correspondence among multiple correlated graphs.
+In the setting of inhomogeneous random graphs, we present and analyze a matching algorithm: first partially match the graphs pairwise and then combine the partial matchings by transitivity.
+Our analysis yields a sufficient condition on the problem parameters to exactly match all nodes across all the graphs.
+In the setting of homogeneous (Erdős-Rényi) graphs, we show that this condition is also necessary, i.e.
+the algorithm works down to the information theoretic threshold.
+This reveals a scenario where exact matching between two graphs alone is impossible, but leveraging more than two graphs allows exact matching among all the graphs.
+Converse results are also given in the inhomogeneous setting and transitivity again plays a role.
+Along the way, we derive independent results about the $k$-core of inhomogeneous random graphs.
 1 Introduction
@@
 This wealth of correlated data presents both opportunities and challenges.
-On one hand, information from various datasets can be combined to increase the fidelity of data - translating to better performance in downstream learning tasks.
+On one hand, information from various datasets can be combined to increase the fidelity of data, which translates to better performance in downstream tasks.
 On the other hand, the interconnected nature of this data also raises privacy and security concerns.
-Linkage attacks, for instance, exploit correlated data to identify individuals in an anonymized network by linking to other sources [NS09].
-This poses a significant threat to user privacy.
-Graph matching is the problem of recovering the underlying latent correspondence between correlated networks.
-The problem finds many applications in machine learning: de-anonymizing social networks [NS08, NS09], identifying similar functional components between species by matching their protein-protein interaction networks [BSI06, KHGPM16], object detection [SS05] and tracking [YYL+16] in computer vision, and textual inference for natural language processing [HNM05].
-In most applications of interest, data is available in the form of several correlated networks.
-For instance, social media users are active each month on 6.7 social platforms on average [Ind23].
-Similarly, reconciling protein-protein interaction networks among multiple species is an important problem in computational biology [SXB08].
-As a first step toward this objective, many research works have studied the problem of matching two correlated graphs.
+Linkage attacks, for instance, exploit correlated data to identify individuals in an anonymized network by linking to other sources (NS, 09).
+The problem of graph alignment is to recover the underlying latent node correspondence between correlated networks, i.e.
+to match the nodes across graphs.
+The problem derives its importance from the fact that nodes in networks are often unlabeled or scrambled – for example, user identities in social networks are anonymized for privacy.
+Applications include:
+  1. 1.
+    Social networks: Matching two social networks amounts to recovering the hidden node identities of users in the anonymized network.
+Narayanan and Shmatikov (NS, 09) established the efficacy of this approach: A third of the users active on both Twitter and Flickr could be identified with only a 12% error rate, simply by matching the anonymous Twitter graph to the Flickr graph.
+In a related work (NS, 08), they also showed that anonymous movie ratings of Netflix users can be recovered by matching to the public Internet Movie Database (IMDb).
+  2. 2.
+    Biological networks: The protein-protein interactomes of two different species are often correlated networks.
+Singh and co-authors (SXB, 08) showed that matching the PPI network of yeast to that of fruit flies allowed identification of conserved functional components between the two species.
+This transfer of knowledge between species is a useful tool in biology and genomics (KHGPM, 16; BSI, 06).
+Similarly, the brain connectomes of healthy humans are correlated (STK, 05), and their alignment is useful in detecting abnormalities (CPPDG, 24).
+  3. 3.
+    Natural language processing: Given a passage of text and a hypothesis, consider the objective of determining whether the hypothesis is entailed by the passage.
+Haghighi and co-authors (HNM, 05) presented an approach to solving this, by first constructing dependency trees of the text and hypothesis respectively, and then matching them to conclude with a yes/no answer after thresholding a score of the output matching.
+  4. 4.
+    Computer vision: Graph matching is employed extensively for vision tasks.
+One example is object detection: Schellewald and Schnörr (SS, 05) showed that objects in images can be detected by matching graph representations of the test image to a feature image that contains the object.
+Yet another example in pattern recognition is the problem of tracking the movement of objects in a sequence of frames (YYL+, 16).
+  5. 5.
+    Autonomous agents: Robots and self-driving cars rely on a multitude of sensors to collect data about their environment.
+Combining information from various sensors increases data fidelity.
+For example, Ye and co-authors (YDF+, 22) showed that graph alignment is useful for scene-map construction when multiple robots explore an unknown environment.
+There is a long history of utilizing domain knowledge to implement graph alignment algorithms for these tasks – see for example (YYL+, 16; VF, 13).
+Even so, the theoretical study of graph matching algorithms and their performance guarantees is a relatively recent undertaking.
+The vast majority of work has focused on pairs of random graphs, particularly Erdős-Rényi graphs.
+Presently, we study exact recovery thresholds with multiple inhomogeneous random graphs.
 1.1 Related Work
-The theoretical study of graph matching algorithms and their performance guarantees has primarily focused on Erdős-Rényi (ER) graphs.
-Pedarsani and Grossglauser [PG11] introduced the subsampling model to generate two such correlated graphs.
-The model entails twice subsampling each edge independently from a parent ER graph to obtain two sibling graphs, both of which are marginally ER graphs themselves.
+The theoretical study of graph matching algorithms and their performance guarantees has focused on the homogeneous setting of Erdős-Rényi (ER) graphs.
+Pedarsani and Grossglauser (PG, 11) introduced the subsampling model to generate two such correlated graphs: the model entails twice subsampling each edge independently from a parent ER graph to obtain two sibling graphs, both of which are marginally ER graphs themselves.
 The goal is then to match nodes between the two graphs to recover the underlying latent correspondence.
 This has been the framework of choice for many works that study graph matching.
-For example, Cullina and Kiyavash studied the problem of exactly matching two ER graphs, where the objective is to match all vertices correctly [CK16, CK17].
+For example, Cullina and Kiyavash (CK, 16, 17) studied the problem of exactly matching two ER graphs, where the objective is to match all vertices correctly.
 They identified a threshold phenomenon for this task: exact recovery is possible if the problem parameters are above a threshold, and impossible otherwise.
-Subsequently, threshold phenomena were also identified for partial graph matching between ER graphs - where the objective is to match only a positive fraction of nodes [GML21, HM23, WXY22, DD23].
-The case of almost-exact recovery - where the objective is to match all but a negligible fraction of nodes - was studied by Cullina and co-authors: a necessary condition for almost exact recovery was identified, and it was shown that the same condition is also sufficient for the $k$-core estimator [CKMP19]; the estimator is described formally in Section 3.
-This estimator proved useful to uncover the fundamental limits for graph matching in other contexts such as the stochastic block model [GRS22] and inhomogeneous random graphs [RS23].
-Ameen and Hajek [AH23] showed some robustness properties of the $k$-core estimator in the context of matching ER graphs under node corruption.
-The estimator plays an important role in the present work as well.
-A sound understanding of ER graphs inspires algorithms for real-world networks.
-Various efficient algorithms have been proposed, including algorithms based on the spectrum of the graph adjacency matrices [FMWX22], node degree and neighborhood based algorithms [DCKG19, DMWX21, MRT23] as well as algorithms based on iterative methods [DL23] and counting subgraphs [MWXY23, BCL+19].
-Some of these are discussed in Section 5 in relation to the present work.
-Incorporating information from multiple graphs to match them has been recognized as an important research direction, for instance in the work of Gaudio and co-authors [GRS22].
-To our knowledge, the only other work to consider matchings among multiple graphs is the work of Rácz and Sridhar [RS21].
-However, the authors solve a different problem of matching stochastic block models, and note that it is possible to exactly match $m$ graphs whenever it is possible to exactly match any two graphs by pairwise matching all the graphs exactly.
-In contrast, we show that under appropriate conditions, it is possible to exactly match $m$ ER graphs even when no two graphs can be pairwise matched exactly.
-Contributions
-In this work, we investigate the problem of combining information from multiple correlated networks to boost the number of nodes that are correctly matched among them.
-We consider the natural generalization of the subsampling model to generate $m$ correlated random graphs, and identify a threshold such that it is impossible for any algorithm to match all nodes correctly across all graphs when the problem parameters are below this threshold.
-Conversely, we show that exact recovery is possible above the threshold.
-This characterization generalizes known results for exact graph matching when $m=2$.
-Subsequently, we show that there is a region in parameter space for which exactly matching any two graphs is impossible using only the two graphs, and yet exact graph matching is possible among $m>2$ graphs using all the graphs.
-We present two algorithms and prove their optimality for this task.
-The first algorithm matches all $m$ graphs simultaneously based on global information about the graphs.
-In contrast, the second algorithm first pairwise matches graphs, and then combines them to match all nodes across all graphs.
-We show that both algorithms correctly match all the graphs all the way down to the information theoretic threshold.
-Finally, we illustrate through simulation that our subroutine to combine information from pairwise comparisons between networks works well when paired with efficient algorithms for graph matching.
-Our analysis also yields some theoretical results about the $k$-core of ER graphs that are of independent interest.
-2 Preliminaries and Setup
-Notation
-In this work, $G\sim\mathsf{ER}(n,p)$ denotes that the graph $G$ is sampled from the Erdős-Rényi distribution with parameters $n$ and $p$, i.e.
-$G$ has $n$ nodes and each edge is independently present with probability $p$.
-For a graph $G$, we denote the set of its vertices by $V\equiv V(G)$ and its edges by $E(G)$.
-The edge status of each vertex pair $\{i,j\}$ with $i\neq j$ is denoted by $G\{i.j\}$, so that $G\{i,j\}=1$ if $\{i,j\}\in E(G)$ and $G\{i,j\}=0$ otherwise.
-The degree of a node $v$ in graph $G$ is denoted $\delta_{G}(v)$.
-Let $\pi$ denote a permutation on $V(G)=\{1,\cdots,n\}$.
-For a graph $G$, denote by $G^{\pi}$ the graph obtained by permuting the nodes of $G$ according to $\pi$, so that
-$G\{i,j\}=G^{\pi}\left\{\pi(i),\pi(j)\right\}\ \forall\ i,j\in V(G)\text{ such that }i\neq j.$
-Standard asymptotic notation $(O(\cdot),o(\cdot),\cdots)$ is used throughout and it is implicit that $n\to\infty$.
-Subsampling model
-Consider the subsampling model for correlated random graphs [PG11], which has a natural generalization to the setting of $m$ graphs.
-In this model, a parent graph $G$ is sampled from the Erdős-Rényi distribution $\mathsf{ER}(n,p)$.
-The $m$ graphs $G_{1},G_{2}^{\prime},\cdots,G_{m-1}^{\prime},G_{m}^{\prime}$ are obtained by independently subsampling each edge from $G$ with probability $s$.
-Finally, the graphs $G_{2},\cdots,G_{m}$ are obtained by permuting the nodes of each of the graphs $G_{2}^{\prime},\cdots,G_{m}^{\prime}$ respectively according to independent permutations $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ sampled uniformly at random from the set of all permutations on $[n]$, i.e.
-$G_{j}=(G_{j}^{\prime})^{\pi^{*}_{1j}}\ \text{ for all }\ j\in\{2,\cdots,m\}.$
-Figure 1 illustrates this process of obtaining correlated graphs using the subsampling model. In this work, we are interested in the setting where $s$ is constant and $p=C\log(n)/n$ for some $C>0$.
-Objective 1.
-Determine conditions on parameters $C$, $s$ and $m$ so that given correlated graphs $G_{1},\cdots,G_{m}$ from the subsampling model, it is possible to exactly recover the underlying correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ with probability $1-o(1)$.
-Stated thus, the underlying correspondences use the graph $G_{1}$ as a reference.
-Thus, for ease of notation, we will use $G_{1}$ and $G_{1}^{\prime}$ interchangeably.
-Note that the underlying correspondence between all the graphs is fixed upon fixing $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$: for any two graphs $G_{i}$ and $G_{j}$, their underlying correspondence is given by $\pi^{*}_{ij}:=\pi^{*}_{1j}\circ(\pi^{*}_{1i})^{-1}$.
-Formally, a matching $(\mu_{12},\cdots,\mu_{1m})$ is a collection of injective functions with domain $\mathsf{dom}(\mu_{1i})\subseteq V$ for each $i$, and co-domain $V$.
+Subsequently, threshold phenomena were also identified for partial graph matching between ER graphs, where the objective is to match only a positive fraction of nodes (GML, 21; HM, 23; WXY, 22; DD, 23; Du, 25).
+The case of almost-exact recovery – where the objective is to match all but a negligible fraction of nodes – was studied by Cullina and co-authors (CKMP, 19): a necessary condition for almost exact recovery was identified, and the same condition was shown to also sufficient be for the $k$-core estimator; the estimator is described in Section 3.
+This estimator proved useful for graph matching in other contexts such as the stochastic block model (GRS, 22) and inhomogeneous random graphs (RS, 23).
+Ameen and Hajek (AH, 23) showed some robustness properties of the $k$-core estimator in the context of matching ER graphs under node corruption.
+The estimator also plays an important role in the present work.
+Other works have studied computational aspects of graph alignment: Various efficient algorithms have been proposed, including algorithms based on the spectrum of the graph adjacency matrices (FMWX, 22), node degree and neighborhood based algorithms (DCKG, 19; DMWX, 21; MRT, 23) as well as algorithms based on iterative methods (DL, 23) and counting subgraphs (MWXY, 23; BCL+, 19; CR, 24).
+While all these works study Erdős-Rényi graphs, there is recent interest in computational aspects of graph alignment in other models such as the stochastic block model (CR, 24), as well as inhomogeneous random graphs (DFW, 23).
+Still other works have studied the graph matching problem under side information – this includes settings where a subset of nodes are matched a priori and provided as seeds (PG, 11; MX, 20; LS, 18), the setting where node attributes are present (WZWW, 24; ZWWW, 24; YC, 24), and the setting where correlation is localized to a subset of nodes (HSY, 24).
+Incorporating information from multiple graphs to match them has been recognized as an important research direction, for instance in the work of Gaudio and co-authors (GRS, 22).
+For example, a user is active, on average, on 6.7 social networks each month (Ind, 23).
+Similarly, reconciling protein-protein interaction networks among multiple species is an important problem in computational biology (SXB, 08).
+These applications also necessitate going beyond ER graphs, since many real-world networks lack the homogeneity of the ER model.
+To our knowledge, a very limited number of papers consider matchings among multiple graphs: The works (JLK, 21) and (RS, 21) have different objectives, and note that it is possible to exactly match $m$ graphs whenever it is possible to exactly match any two graphs by pairwise matching all the graphs exactly.
+In contrast, we show that under appropriate conditions, it is possible to exactly match $m$ graphs even when no two graphs can be pairwise matched exactly.
+The work of Rácz and Zhang (RZ, 24) studies the interplay between graph matching and community detection for multiple correlated stochastic block models.
+The recent work of Vassaux and Massoulié (VM, 25) studies the information-theoretic limit for exact recovery in the Gaussian model, where the observations are multiple correlated Wigner matrices, and also the limit for partial recovery in the homogeneous Erdős-Rényi setting.
+Closest to our work is the work of Rácz and Sridhar (RS, 23), where sufficeint conditions for exact and partial graph matching between two inhomoegenous random graphs are studied.
+1.2 Contributions
+The present work studies fundamental limits of exact recovery with multiple inhomogeneous random graphs.
+These graphs allow for heterogeneity by having each edge $\{i,j\}$ in the graph appear independently with probability $p_{ij}$, and subsume other models such as ER graphs and stochastic block models, and are closely related to random geometric graphs.
+We consider the generalization of the matching problem to $m$ graphs and propose an algorithm to combine information among the graphs.
+A key idea is that of transitive closure – our algorithm works by pairwise matching graphs partially and then boosting the matchings by transitivity.
+Our analysis yields a sufficient condition for exact recovery of multiple inhomogeneous graphs.
+In the homogeneous case where $p_{ij}=p$, we show that this condition is also necessary, thereby characterizing the sharp reconstruction threshold for exactly matching multiple Erdős-Rényi graphs.
+In the heterogeneous setting, we also provide converse results and transitive closure plays a role here as well.
+We present simulation results to corroborate the utility of transitive closure as a black-box, efficient and optimal bridge between the binary and $m$-ary matching problems.
+2 Model
+Let $\mathbf{P}=(p_{ij})_{i,j\in[n]}$ with $0\leq p_{ij}=p_{ji}\leq 1$ denote a symmetric matrix with zero diagonals.
+A graph $G$ is an inhomogeneous random graph on the vertex set $[n]\triangleq\{1,2,\cdots,n\}$ with parameter matrix $\mathbf{P}$, denoted by $G\sim\mathrm{RG}(n,\mathbf{P})$, if each edge $\{i,j\}$ with $i<j$ exists independently with probability $p_{ij}$.
+Consider the natural extension of the subsampling model to $m$ graphs.
+Let $s\in(0,1]$, and obtain $m$ correlated inhomogeneous graphs $G_{1}^{\prime},\cdots,G_{m}^{\prime}$ by subsampling each edge of a parent graph $G\sim\mathrm{RG}(n,\mathbf{P})$ independently with probability $s$.
+These sibling graphs are themselves inhomogeneous random graphs with parameter matrix $\mathbf{P}s$.
+Let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote independent permutations that are each drawn uniformly at random from the set of all permutations on $[n]$.
+Obtain the graphs $G_{2},\cdots,G_{m}$ by permuting the nodes of $G_{k}^{\prime}$ according to $\pi^{*}_{1k}$, for each $k=2,\cdots,m$.
+For example, $\{i,j\}$ is an edge in $G_{k}^{\prime}$ if and only if $\{\pi^{*}_{1k}(i),\pi^{*}_{1k}(j)\}$ is an edge in $G_{k}$.
+An illustration of the subsampling model is provided in Figure 1.
+Figure 1: Obtaining $m$ correlated inhomogeneous graphs via the subsampling model.
+We will use the tuple $(\pi_{12},\cdots,\pi_{1m})$ of permutations to identify the correspondence between all $m$ graphs.
+Stated thus, the graph $G_{1}$ is used as a reference and so we will interchangeably use $G_{1}$ and $G_{1}^{\prime}$.
+Note that the underlying correspondence between all the graphs is fixed upon fixing $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$: for any two graphs $G_{k}$ and $G_{\ell}$, their correspondence is given by $\pi^{*}_{k\ell}\triangleq\pi^{*}_{1\ell}\circ(\pi^{*}_{1k})^{-1}$.
+In this work, we are concerned with exact recovery.
+The goal is to determine necessary and sufficient conditions on the parameters $\mathbf{P}$, $s$ and $m$ so that given correlated inhomogeneous graphs $G_{1},\cdots,G_{m}$ from the subsampling model, it is possible to exactly recover the underlying correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ with probability $1-o(1)$.
+Formally, a matching $(\mu_{12},\cdots,\mu_{1m})$ is a collection of injective functions with domain $\mathrm{dom}(\mu_{1i})\subseteq V$ for each $i$, and co-domain $V$.
 An estimator is simply a mechanism to map any collection of graphs $(G_{1},\cdots,G_{m})$ to a matching.
 We say that an estimator completely matches the graphs if the output mappings $\mu_{12},\cdots\mu_{1m}$ are all complete, i.e.
-they are all permutations on $\{1,\cdots,n\}$.
-Figure 1: Illustration of obtaining $m$ correlated graphs from the subsampling model
-3 Main Results and Algorithm
-This section presents necessary and sufficient conditions to meet 1.
-Theorem 2 (Impossibility).
-Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively.
+they are all permutations on $[n]$.
+If the matching $(\mu_{12},\cdots,\mu_{1m})$ satisfies $\mu_{1j}=\pi^{*}_{1j}$ for all $2\leq j\leq m$, then all nodes are correctly matched across all graphs, i.e.
+the graphs are exactly matched.
+Some special cases of inhomogeneous random graphs
+The inhomogeneous random graph model captures a variety of generative models. Following (RS, 23), we introduce these models conditioned on latent variables intrinsic to them for convenience.
+  1. 1.
+    Erdős-Rényi graph: Given a parameter $p\in[0,1]$, the homogeneous case of ER graphs sets $p_{ij}=p$ for all distinct $i$ and $j$.
+  2. 2.
+    Stochastic block model (SBM) (Abb, 18; HLL, 83): Given a partition $V_{1},\cdots,V_{M}$ of the node set $[n]$ into $M$ communities and edge probabilities $(q_{ab})_{1\leq a<b\leq M}$, the SBM sets $p_{ij}=q_{ab}$ for all distinct $i$ and $j$ such that $i\in V_{a}$ and $j\in V_{b}$.
+  3. 3.
+    Random geometric graph (RGG) (LR, 23): Given a collection $x_{1},\cdots,x_{n}$ of points in a set $S\subseteq\mathbb{R}^{d}$, and parameters $p\in[0,1]$ and $r>0$, the RGG sets $p_{ij}=p\cdot\mathbf{1}_{\left\{\|x_{i}-x_{j}\|_{2}\leq r\right\}}$ for distinct $i$ and $j$.
+  4. 4.
+    Chung-Lu graph (CLG) (CL, 02): Given a collection $(w_{i})_{1\leq i\leq n}$ of positive weights satisfying $w_{i}\leq\sqrt{\sum_{j\in[n]}w_{j}}$ for all $i$, the CLG model sets $p_{ij}=\frac{w_{i}w_{j}}{\sum_{k\in[n]}w_{k}}$ for distinct $i$ and $j$.
+Correlated graphs from any of the above models are obtained via the subsampling model.
+Notation
+Throughout this work, $p_{\max}$ denotes the largest entry in the matrix $\mathbf{P}$, i.e.
+$p_{\max}=\max_{i,j}p_{ij}$.
+For a graph $G$ and node $v$ in $G$, let $\delta_{G}(v)$ denote the degree of $v$.
+Given a matching $\widehat{\mu}$, we denote by $G^{\mu}$ the graph on node set given by the image of $\mu$ obtained by relabeling the nodes of $G$ according to $\mu$, i.e.
+for $i,j\in\mathrm{dom}(\mu)$:
+$\{i,j\}\in E(G)\iff\{\mu(i),\mu(j)\}\in E(G^{\mu}).$
+For any two graphs $H_{1}$ and $H_{2}$ on the same vertex set $V$, denote by $H_{1}\vee H_{2}$ their union graph and by $H_{1}\wedge H_{2}$ their intersection graph.
+An edge $\{i,j\}$ is present in $H_{1}\vee H_{2}$ if it is present in either $H_{1}$ or $H_{2}$.
+Similarly, the edge is present in $H_{1}\wedge H_{2}$ if it is present in both $H_{1}$ and $H_{2}$.
+For a random variable $X$, denote by $M_{X}(t)\triangleq\mathbb{E}[e^{tX}]$ the moment generating function of $X$.
+Standard asymptotic notation $(O(\cdot),o(\cdot),\cdots)$ is used throughout, and it is implicit that $n\to\infty$.
+3 Main Results and Algorithms
+3.1 Exact recovery threshold: Positive results
+Our first result is a sufficient condition for exactly matching all nodes across all the graphs.
+Theorem 1
+Let $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose that there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
+$\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$ (1)
+There exists an estimator, namely Algorithm 1 described in the next subsection, whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies
+$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1).$
+A few remarks are in order.
+  • The sufficient condition (1) of Theorem 1 can be slightly improved to the condition
+    $\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}-\max_{j\in[n]}e^{-d_{j}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$ (2)
+    Let $v^{*}\in\operatorname*{arg\,min}_{i\in[n]}d_{i}$.
+It can be shown that Algorithm 1 applied to the graphs $G_{k}$ yields that all nodes except $v^{*}$ can be exactly matched.
+The node $v^{*}$ is then the only unmatched node and can be manually matched without ambiguity.
+  • Rácz and Sridhar (RS, 23) studied the graph alignment problem when $m=2$.
+Setting $m=2$ in (1) yields $\sum_{i=1}^{n}e^{-d_{i}s^{2}}=o(n^{-\alpha})$, which is equivalent to their sufficient condition.
+  • Algorithm 1 does not require access to the parameter matrix $\mathbf{P}$.
+The sufficient condition (1) can be specialized to various models that are captured by inhomogeneous random graphs. Some of these are presented next.
+Corollary 2 (Correlated SBMs)
+Let $G_{1},\cdots,G_{m}$ be correlated SBMs obtained from the subsampling model.
+Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $q_{ab}=o(n^{\alpha/2-\varepsilon-1})$ for all $a$ and $b$.
 Suppose that
-$Cs\left(1-(1-s)^{m-1}\right)<1.$
-The output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ of any estimator satisfies
-$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=o(1).$
-Theorem 2 implies that the condition $Cs(1-(1-s)^{m-1}>1$ is a necessary condition to exactly match $m$ graphs with probability bounded away from 00.
-We show that this condition is also sufficient to exactly match $m$ graphs with probability going to $1$.
-Theorem 3 (Achievability).
-Let $G_{1},\cdots,G_{m}$ be correlated graphs obtained from the subsampling model with parameters $C$ and $s$, and let $\pi^{*}_{12},\cdots,\pi^{*}_{1m}$ denote the underlying latent correspondences between $G_{1}$ and $G_{2},\cdots,G_{m}$ respectively.
+$\sum_{a=1}^{M}|V_{a}|\cdot\exp\left(-s\left(1-(1-s)^{m-1}\right)\cdot\sum_{b=1}^{m}|V_{b}|q_{ab}\right)=o(n^{-\alpha}).$
+There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
+Corollary 3 (Correlated RGGs)
+Let $G_{1},\cdots,G_{m}$ be correlated RGGs obtained from the subsampling model.
+Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p=o(n^{\alpha/2-\varepsilon-1})$.
+For a node $i$, let $N_{r}(i)$ denote the number of nodes in its $r$-neighborhood, i.e.
+$N_{r}(i)\triangleq|\{j:\|x_{i}-x_{j}\|_{2}\leq r\}|$.
 Suppose that
-$Cs\left(1-(1-s)^{m-1}\right)>1.$
-There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies
-$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\ \widehat{\pi}_{13}=\pi^{*}_{13},\cdots,\ \widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1).$
-Theorems 2 and 3 together characterize the threshold for exact recovery. A few remarks are in order.
-  1. 1.
-    For $m=2$, the condition $Cs(1-(1-s)^{m-1})>1$ reduces to $Cs^{2}>1$, which is known to be necessary and sufficient for exactly matching two graphs [CK17, WXY22].
-  2. 2.
-    For any $m>2$, there is a non-empty region in the parameter space defined by
-    $Cs(1-(1-s)^{m-1})>1>Cs^{2}.$
-    For any $C$ and $s$ in this region, it is impossible to exactly match any two graphs $G_{i}$ and $G_{j}$ without using the other $m-2$ graphs as side information.
-Upon using them, however, it is possible to exactly match all nodes across the $m$ graphs.
-This is illustrated in Figure 2.
+$\sum_{i=1}^{n}e^{-ps(1-(1-s)^{m-1})\cdot N_{r}(i)}=o(n^{-\alpha}).$
+There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
+Corollary 4 (Correlated CLGs)
+Let $G_{1},\cdots,G_{m}$ be correlated CLGs obtained from the subsampling model.
+Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $w_{i}=o(n^{\alpha/2-\varepsilon-1})\sqrt{\sum_{k\in[n]}w_{k}}$ for all $i$.
+Suppose that
+$\sum_{i=1}^{n}e^{-w_{i}s(1-(1-s)^{m-1})}=o(n^{-\alpha}).$
+There is an estimator whose output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ satisfies $\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-o(1)$.
+Next, we show that the sufficient condition (2) can be improved for a class of parameter matrices if the algorithm has access to $P$.
+Theorem 5
+Suppose $\varepsilon$ and $s$ are fixed with $\varepsilon>0$ and $0<s\leq 1$ If $d$ varies with $n$ such that:
+$ds\geq(1+\epsilon)\log n,$ (3)
+then there exists a parameter matrix $P$ with $d_{i}\triangleq\sum_{j\neq i}P_{ij}\leq d$ for $i\in[n]$ and $p_{\max}\triangleq\max_{i,j}P_{i,j}=\frac{\log^{2}(n)}{n}$ such that exact recovery is possible, i.e.
+$m$ graphs $G_{1},G_{2},\cdots,G_{m}$ obtained by subsampling each edge independently with probability $s$ from a parent graph $G\sim\mathrm{RG}(n,P)$ can be exactly matched with probability $1-o(1)$ by some algorithm that has access to $P$.
+Taking $ds=(1+\epsilon)\log n$ in Theorem 5 yields the following corollary.
+Corollary 6
+There exists a parameter matrix $P$ satisfying
+$\sum_{i=1}^{n}e^{-d_{i}s(1-2\varepsilon)}-\max_{j\in[n]}e^{-d_{j}s(1-2\varepsilon)}=\Omega(n^{\varepsilon}),$ (4)
+such that exact recovery is possible by some algorithm that has access to $P$.
+Remark 7
+Corollary 6 shows that if $(1-s)^{m-1}>2\varepsilon$, then the sufficient condition (2) of Theorem 1 does not hold for some $P$, but exact recovery is nevertheless possible.
+Roughly speaking, the matrix $P$ itself is similar to an additional graph that can help in matching by some algorithm that has access to $P$.
+3.2 Exact recovery threshold: negative results
+While Corollary 6 shows that (2) is not a sharp sufficient condition for general parameter matrices, we prove that this condition is sharp in the homogeneous case of Erdős-Rényi graphs.
+This provides the sharp threshold for exactly matching $m$ ER graphs.
+Theorem 8
+Let $C>0$ and $p_{ij}=C\log n/n$ for each $i<j$, and let $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose that
+$\sum_{i=1}^{n}e^{-d_{i}s(1-(1-s)^{m-1})}=\Omega(1),$ (5)
+or equivalently that $Cs(1-(1-s)^{m-1})<1$. Then exact recovery is impossible, i.e. the output $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ of any estimator satisfies
+$\mathbb{P}\left(\widehat{\pi}_{12}=\pi^{*}_{12},\cdots,\widehat{\pi}_{1m}=\pi^{*}_{1m}\right)=1-\Omega(1).$
+In the setting of ER graphs with $p_{ij}=C\log(n)/n$, Theorems 1 and 8 establish that exact recovery is possible if and only if $Cs(1-(1-s)^{m-1})>1$.
+When $m=2$, this condition reduces to $Cs^{2}>1$, recovering the reconstruction threshold of Cullina and Kiyavash (CK, 17).
+Thus, when $m>2$, there is a regime given by
+$Cs(1-(1-s)^{m-1})>1>Cs^{2},$
+where exact recovery is impossible using only the two graphs but possible using all $m$ graphs. This is illustrated in Figure 2.
 (a) $m=3$
 (b) $m=10$
-Figure 2: Regions in parameter space.
+Figure 2: Regions in parameter space for the homogeneous setting.
 Orange: Exactly matching $m$ graphs is impossible even with $m$ graphs.
@@
 Striped: Impossible to match $2$ graphs using only the $2$ graphs, but possible using $m$ graphs as side information.
-3.1 Algorithms for exact recovery
-For any two graphs $H_{1}$ and $H_{2}$ on the same vertex set $V$, denote by $H_{1}\vee H_{2}$ their union graph and by $H_{1}\wedge H_{2}$ their intersection graph.
-An edge $\{i,j\}$ is present in $H_{1}\vee H_{2}$ if it is present in either $H_{1}$ or $H_{2}$.
-Similarly, the edge is present in $H_{1}\wedge H_{2}$ if it is present in both $H_{1}$ and $H_{2}$.
-A natural starting point is to study the maximum likelihood estimator (MLE) because it is optimal. To that end, we compute the log-likelihood function; the details are deferred to Appendix A.
-Theorem 4.
-Let $\pi_{12},\cdots,\pi_{1m}$ denote a collection of permutations on $\{1,\cdots,n\}$. Then
-$\log\mathbb{P}\left(G_{1},\cdots,G_{m}\ |\ \pi^{*}_{12}=\pi_{12},\cdots,\pi^{*}_{1m}=\pi_{1m}\right)\propto\text{const.}-\left|E\left(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}\right)\right|,$
-where const. depends only on $p,s$ and $G_{1},\cdots,G_{m}$.
-Theorem 4 reveals that the MLE for exactly matching $m$ graphs has a neat interpretation: simply pick $\pi_{12},\cdots,\pi_{1m}$ to minimize the number of edges in the corresponding union graph.
-This is presented as Algorithm 1.
-Despite this nice interpretation of the MLE, its analysis is quite cumbersome.
-We instead present and analyze a different estimator, presented as Algorithm 2.
-require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$
-1 for $(\pi_{12},\pi_{13},\cdots,\pi_{1m})$ such that each $\pi_{1j}$ is a permutation on $[n]$ do
-2       $W(\pi_{12},\cdots,\pi_{1m})\leftarrow|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|$
-3
-4 end for
-return $(\widehat{\pi}_{12}^{\text{ML}},\cdots,\widehat{\pi}_{1m}^{\text{ML}})\in\operatorname*{arg\,max}_{\pi_{12},\cdots,\pi_{1m}}W(\pi_{12},\cdots,\pi_{1m})$
-Algorithm 1 Maximum likelihood estimator
-require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $V$, Integer $k$
-// Step 1: Pairwise matching
+We turn next to negative results in the heterogeneous case. To that end, consider the following lemma about isolated nodes in an inhomogeneous random graph.
+Lemma 9
+Let $H\sim\mathrm{RG}(n,\mathbf{Q})$, and let $r_{i}=\sum_{j\neq i}q_{ij}$. Suppose that $q_{\max}=o(1/\sqrt{n})$ and
+$\left(\sum_{i=1}^{n}e^{-r_{i}}\right)-\max_{j\in[n]}e^{-r_{j}}=\Omega(1).$ (6)
+With probability bounded away from 00, there are at least two isolated nodes in $H$.
+Lemma 9 is proved in Appendix B.1.
+Applying it to $H=G_{1}$ yields a regime where there are at least two isolated nodes $u$ and $v$ in $G_{1}$, with probability bounded away from 00.
+Since it is impossible to distinguish between $u$ and $v$ in $G_{1}$, it is impossible to match them with probability greater than $1/2$.
+Exactly matching the $m$ graphs is therefore impossible in this regime.
+This implies the following converse to Corollary 6.
+Theorem 10
+Consider the model of $m$ correlated graphs with parameters $P$ and $s$. Let $d_{i}=\sum_{j\neq i}p_{ij}$. If $p_{\max}=o(1/\sqrt{n})$ and
+$\left(\sum_{i=1}^{n}e^{-d_{i}s}\right)-\max_{j\in[n]}e^{-d_{j}s}=\Omega(1),$ (7)
+then for any algorithm the probability of exactly matching all $m$ graphs is $1-\Omega(1)$.
+Remark 11
+If we ask for a given $P,s,m$ whether exact recovery is possible by some algorithm, then implicitly algorithms with access to $P$ should be included because $P$ is deterministic.
+It remains an open problem to give the sharp recovery threshold for a given general $P,s,m$.
+Theorem 1 provides a performance floor while Theorem 10 provides a performance ceiling.
+Theorem 8 shows Theorem 1 is tight for homogeneous $P$ while Corollary 6 shows Theorem 10 is tight for some choices of $P$.
+Remark 12
+Lemma 9 can also be used to give a second proof of the converse result in the homogeneous setting (i.e.
+Theorem 8) in the special case of $m=2$.
+Setting $H=G_{1}\wedge G_{2}^{\prime}\sim\mathrm{RG}(n,\mathbf{P}s^{2})$ in Lemma 9, we have that if $p_{\max}=o(1/\sqrt{n})$ and
+$\left(\sum_{i=1}^{n}e^{-d_{i}s^{2}}\right)-\max_{j\in[n]}e^{-d_{j}s^{2}}=\Omega(1),$ (8)
+then with probability bounded away from 00, there are at least two nodes $u$ and $v$ that are isolated in $G_{1}\wedge G_{2}^{\prime}$, i.e.
+$u$ and $v$ both have disjoint sets of neighbors in $G_{1}$ and $G_{2}^{\prime}$.
+This implies that the maximum likelihood for Erdős-Rényi graphs, given in Appendix A, fails.
+In the heterogeneous setting, however, the existence of two isolated nodes in the intersection graph $G_{1}\wedge G_{2}^{\prime}$ does not imply impossibility of exact matching, because the maximum likelihood estimator no longer has this simple form.
+3.3 Estimation Algorithm
+The algorithm that achieves the guarantee of Theorem 1 is based on the idea of transitive closure: First, the graphs are matched pairwise to obtain partial matchings $\widehat{\mu}_{ij}$ between each pair of graphs.
+These matchings can then be boosted via transitivity.
+The idea is formalized in Algorithm 1, which runs in two steps: In step 1, the $k$-core estimator, for a suitable choice of $k$, is used to pairwise match all the graphs.
+For any $i$ and $j$, the $k$-core estimator selects a permutation $\widehat{\nu}_{ij}$ to maximize the size of the $k$-core of $G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}}$ (the $k$-core of a graph $G$ is the largest subset of vertices $\mathrm{core}_{k}(G)$ such that the induced subgraph has minimum degree at least $k$).
+It then outputs a matching $\widehat{\mu}_{ij}$ by restricting the domain of $\widehat{\nu}_{ij}$ to $\mathrm{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$.
+Note that the matchings $\widehat{\mu}_{ij}$ need not be complete matchings.
+In step 2, these partial matchings are boosted as follows: If a node $v$ is unmatched between two graphs $G_{i}$ and $G_{j}$, then search for a sequence of graphs $G_{i},G_{k_{1}},\cdots,G_{k_{\ell}},G_{j}$ such that $v$ is matched between any two consecutive graphs in the sequence.
+If such a sequence exists, then extend $\widehat{\mu}_{ij}$ to include $v$ by transitively matching it from $G_{i}$ to $G_{j}$.
+We emphasize the following characteristics of the algorithm.
+require :  Graphs $G_{1},G_{2},\cdots,G_{m}$ on a common vertex set $[n]$, Integer $k$
+// Step 1: Pairwise matching using the $k$-core estimator
 1 for $\{i,j\}$ in $\{1,\cdots,m\}$ such that $i<j$ do
-2       $\widehat{\nu}_{ij}\leftarrow\operatorname*{arg\,max}_{\pi}|\mathsf{core}_{k}\left(G_{i}\wedge G_{j}^{\pi}\right)|$
-       $\widehat{\mu}_{ij}\leftarrow\widehat{\nu}_{ij}\text{ with domain restricted to}\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$
-        // $k$-core estimator
-3
+2       $\widehat{\nu}_{ij}\leftarrow\operatorname*{arg\,max}_{\pi}|\mathrm{core}_{k}\left(G_{i}\wedge G_{j}^{\pi}\right)|$
+3       $\widehat{\mu}_{ij}\leftarrow\widehat{\nu}_{ij}\text{ with domain restricted to}\mathrm{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$
+4
 5 end for
 // Step 2: Boosting through transitive closure
-5 for $v\in V$ do
+6 for $v\in[n]$ do
 7       for $j=2,\cdots,m$ do
@@
 9                   Set $\widehat{\pi}_{1j}(v)=v^{\prime}$
-9
+10
 11             end if
-11
+12
 13       end for
-13
+14
 15 end for
@@
 Algorithm 1 Matching through transitive closure
-Algorithm 2 runs in two steps: In step 1, the $k$-core estimator, for a suitable choice of $k$, is used to pairwise match all the graphs.
-For any $i$ and $j$, the $k$-core estimator selects a permutation $\widehat{\nu}_{ij}$ to maximize the size of the $k$-core [^1] The $k$-core of a graph $G$ is the largest subset of vertices $\mathsf{core}_{k}(G)$ such that the induced subgraph has minimum degree at least $k$.
-of $G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}}$.
-It then outputs a matching $\widehat{\mu}_{ij}$ by restricting the domain of $\widehat{\nu}_{ij}$ to $\mathsf{core}_{k}(G_{i}\wedge G_{j}^{\widehat{\nu}_{ij}})$.
-These matchings $\widehat{\mu}_{ij}$ need not be complete - in fact, each of them is a partial matching with high probability whenever $Cs^{2}<1$.
-In step 2, these partial matchings are boosted as follows: If a node $v$ is unmatched between two graphs $G_{i}$ and $G_{j}$, then search for a sequence of graphs $G_{i},G_{k_{1}},\cdots,G_{k_{\ell}},G_{j}$ such that $v$ is matched between any two consecutive graphs in the sequence.
-If such a sequence exists, then extend $\widehat{\mu}_{i,j}$ to include $v$ by transitively matching it from $G_{i}$ to $G_{j}$.
-In Section 4.2, we show that Algorithm 2 correctly matches all nodes across all graphs with probability $1-o(1)$, whenever the necessary condition $Cs(1-(1-s)^{m-1})>1$ holds.
-We remark that this also implies that Algorithm 1 succeeds under the same condition, because the MLE is optimal.
-Note that the MLE selects all permutations $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ simultaneously based on their union graph.
-In contrast, Algorithm 2 only ever makes pairwise comparisons between graphs.
-Perhaps surprisingly, it turns out that this is sufficient for exact recovery.
-An analysis of Algorithm 2 is presented in Section 4.
-Along the way, independent results of interest on the $k$-core of Erdős-Rényi graphs are obtained.
-4 Proof Outlines and Key Insights
-4.1 Impossibility of exact graph matching (Theorem 2)
-This result has a simple proof following a genie-aided converse argument. The idea is to reduce the problem to that of matching two graphs by providing extra information to the estimator.
-Proof of Theorem 2.
-If the correspondences $\pi^{*}_{12},\cdots,\pi^{*}_{1,m-1}$ were provided as extra information to an estimator, then the estimator must still match $G_{m}$ with the union graph $G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$.
-This can be viewed as an instance of matching two graphs obtained by asymmetric subsampling: the graph $G_{m}$ is obtained from a parent graph $G\sim\mathsf{ER}(n,C\log(n)/n)$ by subsampling each edge independently with probability $s_{1}:=s$, and the graph $\widetilde{G}_{m-1}:=G_{1}^{\prime}\vee G_{2}^{\prime}\vee\cdots\vee G_{m-1}^{\prime}$ is obtained from $G$ by subsampling each edge independently with probability $s_{2}:=1-(1-s)^{m-1}$.
-Cullina and Kiyavash studied this model for matching two graphs: Theorem 2 of [CK17] establishes that matching $G_{m}$ and $\widetilde{G}_{m-1}$ is impossible if $Cs_{1}s_{2}<1$, or equivalently if $Cs(1-(1-s)^{m-1})<1$.
-∎
-4.2 Achievability of exact graph matching (Theorem 3)
+  • Algorithm 1 only ever makes pairwise comparisons between graphs.
+In contrast, the maximum likelihood estimator for the simpler homogeneous setting selects the permutations $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ simultaneously based on all the graphs.
+Despite this, Algorithm 1 works down to the threshold for exact recovery.
+  • The transitive closure sub-routine (step 2) has low computational complexity, $O(m^{2}n\log n)$.
+Further, it can be used to boost the pairwise matchings obtained by any estimator in step 1, i.e.
+it can be used as a black-box bridge between binary graph matching and $m$-ary graph matching whenever partial pairwise matchings are available.
+Our choice of the $k$-core estimator for step 1 is motivated by its precision, i.e.
+the $k$-core estimator does not make mistakes with high probability (CKMP, 19; RS, 23).
+This is crucial to avoid compounding errors during the boosting step.
+This precision is quantified formally later in the analysis (see Section 5.1.1).
+4 Preliminaries: On the $k$-core of Inhomogeneous Graphs
+This section presents useful preliminaries and lemmas for the proof of Theorem 1.
+The results are stated for general inhomogeneous graphs and may be of interest more generally beyond the scope of the graph alignment problem.
+The main objective of this section is to establish that under appropriate conditions, all nodes with sufficiently large degree in an inhomogeneous graph are in its $k$-core with high probability.
+Lemma 13
+Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and $d_{i}=\sum_{j\neq i}p_{ij}$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
+$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$
+Then, for any positive integer $k$ and any $v\in[n]$,
+($\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)=o(1/n).$
+We build up to the proof of this lemma by establishing some results about the number of low-degree nodes (in Section 4.1), and about the size of the $k$-core in inhomogeneous random graphs (in Section 4.2).
+These results are then used to prove Lemma 13 in Section 4.3.
+4.1 On low degree nodes in an inhomogeneous graph
+Lemma 14
+Let $p\in[0,1]$. Consider independent random variables $X$, $Y$ and $Z$ with distributions
+$X\sim\mathrm{Bern}(p),\qquad Y\sim\mathrm{Pois}(p),\qquad Z\sim 2\cdot\mathrm{Pois}(p/2).$
+Then, for all $t$, the moment generating functions of $X$, $Y$ and $Z$ are ordered as
+$M_{X}(t)\leq M_{Y}(t)\leq M_{Z}(t).$
+Proof. We have
+$M_{X}(t)=1+p(e^{t}-1)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(p(e^{t}-1)\right)=M_{Y}(t)\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\exp\left(p\left(\frac{e^{2t}-1}{2}\right)\right)=M_{Z}(t),$
+[…]
+Lemma 15
+Let $n>0$ and let $X_{1},\cdots,X_{n}$ be independent random variables with $X_{i}\sim\text{Bern}(p_{i})$. Let $m$ be an integer so that $0\leq m\leq n$, and define
+$X:=\sum_{i\leq m}X_{i}+\sum_{j>m}2X_{j}.$
+Then, for any $t<\mathbb{E}X$,
+$\mathbb{P}\left(X\leq t\right)\leq\exp\left(-\frac{\mathbb{E}X}{2}+\frac{t}{2}\log\left(\frac{e\cdot\mathbb{E}X}{t}\right)\right)$
+Proof. Let $Y_{i}\sim\text{Pois}(p_{i})$ and $Z_{i}\sim 2\cdot\text{Pois}(p_{i}/2)$. It follows that
+$M_{X}(t)$ $=\prod_{i=1}^{m}M_{X_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{X_{j}}(2t)$
+$\leq\prod_{i=1}^{m}M_{Z_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{Y_{j}}(2t)$
+$=\prod_{i=1}^{m}M_{Z_{i}}(t)\cdot\!\!\prod_{j=m+1}^{n}\!\!M_{2Y_{j}}(t)=M_{Z}(t),$
+where $Z:=\sum_{i=1}^{m}Z_{i}+\sum_{j=m+1}^{n}2Y_{j}$ is distributed as $Z\sim 2\cdot\text{Pois}\left(\mathbb{E}X/2\right)$.
+Therefore, Chernoff tail bounds that hold for $Z$ also hold for $X$. In particular, for any $t<\mathbb{E}X$,
+$\mathbb{P}\left(X\leq t\right)\leq\inf_{\theta\geq 0}e^{\theta t}M_{X}(-\theta)\leq\inf_{\theta\geq 0}e^{\theta t}M_{Z}(-\theta)=\inf_{\theta\geq 0}\exp\left(\theta t+\frac{\mathbb{E}{X}}{2}\left(e^{-2\theta}-1\right)\right).$
+[…]
+Lemma 16
+Let $G\sim\mathrm{RG}(n,\mathbf{P})$, and let $d_{i}=\sum_{j\neq i}p_{ij}$.
+Let $r>0$ be an integer and let $Z_{r}$ denote the set of nodes of $G$ with degree less than or equal to $r$.
+Suppose that for some $\alpha>0$,
+$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$ (9)
+  • For any $\delta>1-\alpha/2$,
+    $\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)=o(1/n).$ (10)
+  • If $p_{\max}=o(n^{\alpha/2-1})$, then (10) is true for all $\delta>1-\alpha$.
+Proof. If strictly more than $n^{\delta}$ nodes all have degree less than or equal to $r$, then a subset of exactly $n^{\delta}$ nodes has average degree less than $r$, i.e.
+$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $=\mathbb{P}\left(\exists\ T\subseteq[n]\colon\left\{|T|\geq n^{\delta}\right\}\cap\left\{\max_{u\in T}\delta_{G}(u)\leq r\right\}\right)$
+$\leq\mathbb{P}\left(\exists\ T\subseteq[n]\colon\left\{|T|=n^{\delta}\right\}\cap\left\{\max_{u\in T}\delta_{G}(u)\leq r\right\}\right)$
+$\leq\mathbb{P}\left(\exists\ S\subseteq[n]\colon\left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{u\in S}\delta_{G}(u)\leq r|S|\right\}\right).$ (11)
+Let $G_{uv}$ denote the indicator random variable for the presence of edge $\{u,v\}$ in $G$. For any set $S\subseteq[n]$, define
+$\mu_{S}:=\sum_{i\in S}\mathbb{E}[\delta_{G}(i)]=\sum_{i\in S}d_{i}.$
+Since for any $i\in[n]$, we have that $e^{-d_{i}}\leq\sum_{j=1}^{n}e^{-d_{j}}\leq n^{-\alpha},$ it follows that $d_{i}\geq\alpha\log n$.
+Therefore, for any $S\subseteq[n]$, we have that $\mu_{S}\geq\alpha|S|\log n$.
+Hence,
+$\mathbb{P}\left(\sum_{u\in S}\delta_{G}(u)\leq r|S|\right)$ $\!=\mathbb{P}\left(\sum_{\begin{subarray}{c}u\in S\\ v\in[n]\setminus S\end{subarray}}\!\!\!G_{uv}+\!\!\sum_{\begin{subarray}{c}u,v\in S\\ u<v\end{subarray}}2G_{uv}\leq r|S|\right)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(\!-\frac{\mu_{S}}{2}\!+\!\frac{r|S|}{2}\log\left(\frac{e\cdot\mu_{S}}{r|S|}\right)\!\right)$
+$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\exp\left(-\frac{\alpha|S|}{2}\log(n)+\frac{r|S|}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\right)$ (12)
+where (a) uses Lemma 15 and (b) uses that the function on the right hand side of (b) is decreasing in $\mu_{S}$ on the interval $(r|S|,\infty)$, and that $r|S|\leq\alpha|S|\log n\leq\mu_{S}$ for all sufficiently large $n$.
+Finally, combining (11) and (12), and using that $|S|=n^{\delta}$ yields
+$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $\leq\!\!\sum_{S\subseteq[n]:|S|=n^{\delta}}\!\!\!\!\mathbb{P}\left(\sum_{u\in S}\delta_{G}(v)\leq r|S|\right)\leq\binom{n}{n^{\delta}}\exp\left(-\frac{\alpha n^{\delta}}{2}\log n+\frac{rn^{\delta}}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\!\right)$
+$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\exp\left(\left(1-\delta-\frac{\alpha}{2}\right)n^{\delta}\log n+\frac{rn^{\delta}}{2}\log\left(\frac{e\cdot\alpha\log n}{r}\right)\right)=o(1/n),$
+whenever $\delta>1-\alpha/2$. Here, (a) uses the fact that $\binom{n}{n^{\delta}}\leq\exp\left((1-\delta)n^{\delta}\log n\right)$. Hence, (10) follows.
+Suppose that $p_{\max}=o(n^{\alpha/2-1})$. Notice that
+$\sum_{u\in S}\delta_{G}(u)=\sum_{u\in S}\sum_{\begin{subarray}{c}v\in[n]\\ v\neq u\end{subarray}}G_{uv}\leq\sum_{u\in S}\sum_{\begin{subarray}{c}v\in[n]\setminus S\end{subarray}}G_{uv}\triangleq Y_{S},$
+and so we have that
+$\mathbb{P}\left(\sum_{u\in S}\delta_{G}(u)\leq r|S|\right)\leq\mathbb{P}\left(Y_{S}\leq r|S|\right)\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}e^{-\mathbb{E}\left[Y_{S}\right]}\cdot\left(\frac{e\cdot\mathbb{E}\left[Y_{S}\right]}{r|S|}\right)^{r|S|},$ (13)
+where (a) uses that $Y_{S}$ is a sum of independent Bernoulli random variables, and so the Chernoff tail bound (19) can be applied. On the other hand,
+$\mathbb{E}\left[Y_{S}\right]=\sum_{u\in S}d_{u}-\sum_{u\in S}\sum_{\begin{subarray}{c}v\in S\\ v\neq u\end{subarray}}2p_{uv}\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\sum_{u\in S}d_{u}-\binom{|S|}{2}\cdot(2p_{\max})\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}\alpha|S|\log(n)-|S|^{2}\cdot p_{\max}\triangleq\widetilde{\mu}_{S},$
+where (a) uses that $p_{uv}\leq p_{\max}$ and (b) uses that $\mu_{S}\geq\alpha|S|\log n$. Continuing from (13), we have
+$\mathbb{P}\left(Y_{S}\leq r|S|\right)\leq e^{-\mathbb{E}\left[Y_{S}\right]}\cdot\left(\frac{e\cdot\mathbb{E}\left[Y_{S}\right]}{r|S|}\right)^{r|S|}\leq e^{-\widetilde{\mu}_{S}}\cdot\left(\frac{e\cdot\widetilde{\mu}_{S}}{r|S|}\right)^{r|S|},$ (14)
+where the last step uses that the function $x\mapsto e^{-x}\cdot\left(\frac{e\cdot x}{r|S|}\right)^{r|S|}$ is decreasing in $x$ whenever $x\geq r|S|$.
+Thus, for any $S$ such that $|S|=n^{\delta}$, we have from (14) that
+$\mathbb{P}\left(Y_{S}\leq rn^{\delta}\right)$ $\leq\exp\left(-n^{\delta}\left(\alpha\log(n)-r-n^{\delta}p_{\max}\right)\right)\cdot\left(\frac{\alpha\log(n)-n^{\delta}p_{\max}}{r}\right)^{rn^{\delta}}$
+$=\exp\left(-n^{\delta}\left(\alpha\log(n)-r-n^{\delta}p_{\max}-r\log\left(\frac{\alpha\log(n)-n^{\delta}p_{\max}}{r}\right)\right)\right).$ (15)
+Combining this with (11), a union bound yields that
+$\mathbb{P}\left(|Z_{r}|>n^{\delta}\right)$ $\!\leq\mathbb{P}\left(\exists\ S\subseteq[n]:\left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{u\in S}\delta_{G}(u)\leq r|S|\right\}\right)\leq\binom{n}{n^{\delta}}\mathbb{P}\left(Y_{S}\leq rn^{\delta}\right)$
+$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\exp\!\left(-n^{\delta}\Big{[}(\alpha\!+\!\delta\!-\!1)\log(n)-n^{\delta}p_{\max}-r\log\Big{(}\frac{\alpha\log(n)\!-\!n^{\delta}p_{\max}}{r}\Big{)}-r-1\Big{]}\right),$ (16)
+[…]
+4.2 On the size of the $k$-core of an inhomogeneous random graph
+The following concentration inequalities for a sum of Bernoulli random variables are used frequently in the analysis.
+Lemma 17
+Let $X$ be a sum of $n$ independent Bernoulli random variables, and let $\mathbb{E}[X]=np$. Then,
+  1. 1.
+    For any $\delta>0$,
+    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{np}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)np}.$ (17)
+  2. 2.
+    For any $\delta>5$,
+    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq 2^{-(1+\delta)np}.$ (18)
+  3. 3.
+    For any $\delta\in(0,1)$,
+    $\mathbb{P}\left(X\leq(1-\delta)np\right)\leq\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{np}=e^{-np}\cdot\left(\frac{e}{1-\delta}\right)^{(1-\delta)np}.$ (19)
+[…]
+Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and let $k>0$.
+Let $v$ be any node of $G$ and consider the induced subgraph $G-v$ of $G$ on the vertex set $[n]\setminus\{v\}$.
+The objective of this section is to show that the $k$-core of $G-v$ is sufficiently large with probability $1-o(1/n)$.
+Formally, we wish to prove the following.
+Lemma 18
+Let $G\sim\mathrm{RG}(n,\mathbf{P})$, and let $v$ be a node of $G$. Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
+$\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha}).$
+Then for any $\delta>1-\alpha$, the size of the $k$-core of $G-v$ satisfies
+$\mathbb{P}\left(\left|\mathrm{core}_{k}(G-v)\right|<n-3n^{\delta}\right)=o(1/n).$
+require :  Graph $G$, Set $U\subseteq V(G)$.
+1 $U_{0}\leftarrow U$
+2 for $i=0,1,2,3,\cdots$ do
+3       if  there exists $u\in V\setminus U_{i}$ such that $u$ has $3$ or more neighbors in $U_{i}$ then
+4             $U_{i+1}\leftarrow U_{i}\cup\{u\}$
+5
+6       else
+7             return $U_{i}$
+8
+9       end if
+10
+11 end for
+Algorithm 2 Łuczak expansion
+To this end, consider Algorithm 2 to identify a subset of the $k$-core, originally proposed by Łuczak (Łuc, 91).
+Note that the for loop eventually terminates - the set $V\setminus U_{i}$ is empty, for example, when $i=n$ for any input set $U$.
+The key is to realize that the for loop terminates quickly when the input $U$ is $Z_{k+1}$ , i.e.
+the set of vertices of the input graph $G$ whose degree is $k+1$ or less.
+Furthermore, for that input, the complement of the set output by the algorithm is contained in the $k$-core.
+This is formally stated in Lemma 19 in what follows.
+Lemma 19
+Let $G\sim\mathrm{RG}(n,\mathbf{P})$ and let $v$ be a node of $G$. Let $U_{f}$ be the output of Algorithm 2 with input graph $G-v$ and $U=Z_{k+1}$. Then,
+  1. (a)
+    $U_{f}^{c}\subseteq\mathrm{core}_{k}(G-v)$.
+  2. (b)
+    If there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and $\sum_{i=1}^{n}e^{-d_{i}}=o(n^{-\alpha})$, then for any $\delta>1-\alpha$,
+    $\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)=o(1/n).$
+Proof.
+(a) The proof is by construction: Since $U_{f}$ is obtained by adding exactly $f$ nodes to $U_{0}$, it follows that $U_{f}^{c}\subseteq U_{0}^{c}=Z_{k+1}^{c}$, so each node in $U_{f}^{c}$ has degree $k+2$ or more in $G-v$.
+Further, each node in $U_{f}^{c}$ has at most $2$ neighbors in $U_{f}$, else the for loop would not have terminated.
+Thus, the subgraph of $G-v$ induced on the set $U_{f}^{c}$ has minimum degree at least $k$, and the result follows.
+(b) Let $\delta^{\prime}$ be such that $1-\alpha<\delta^{\prime}<\min\left\{\delta,1-\alpha+2\varepsilon\right\}$.
+If $|U_{f}|>3n^{\delta}$, then either $|U_{0}|>3n^{\delta^{\prime}}$ or there is some $M$ in $\{0,1,\cdots,f\}$ for which $|U_{M}|=3n^{\delta^{\prime}}$.
+Therefore,
+$\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)$ $\leq\mathbb{P}\left(|U_{0}|>3n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta^{\prime}}\right)$
+$=o(1/n)+\underbrace{\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta^{\prime}}\right)}_{(\star)},$
+where the first term is $o(1/n)$ by Lemma 16.
+Note that each iteration $i=0,1,\cdots,M-1$ of the for loop adds exactly $1$ vertex and at least $3$ edges to the subgraph of $G-v$ induced on $U_{M}$.
+Therefore, the induced subgraph $G|_{U_{M}}$ has $3n^{\delta}$ vertices and at least $3\left(|U_{M}|-|U_{0}|\right)$ edges.
+Let us couple the inhomogeneous random graph $G$ with an Erdős-Rényi random graph $\widetilde{G}$ on $n$ nodes with marginal distribution $\widetilde{G}\sim\text{ER}(n,p_{\max})$, such that $G$ is a subgraph of $\widetilde{G}$.
+Thus,
+(($\star$)( ⋆ ) ( $\leq\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta^{\prime}}\text{ and }|F|\geq 3\left(3n^{\delta^{\prime}}-|U_{0}|\right)\right)$
+$\leq\mathbb{P}\left(|U_{0}|>n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\text{ subgraph $H=(W,F)$ of }G\text{ s.t. }|W|=3n^{\delta^{\prime}}\text{ and }|F|\geq 6n^{\delta^{\prime}}\right)$
+$\leq\mathbb{P}\left(|U_{0}|>n^{\delta^{\prime}}\right)+\mathbb{P}\left(\exists\text{ subgraph $\widetilde{H}\!=\!(\widetilde{W},\widetilde{F})$of }\widetilde{G}\text{ s.t.
+}|\widetilde{W}|=3n^{\delta^{\prime}}\text{ and }|\widetilde{F}|\geq 6n^{\delta^{\prime}}\right)$
+$\leq o(1/n)+\underbrace{\binom{n}{3n^{\delta^{\prime}}}\cdot\mathbb{P}\left(\text{Bin}\left(\binom{3n^{\delta^{\prime}}}{2},p_{\max}\right)>6n^{\delta^{\prime}}\right)}_{(\star\star)},$
+where the last step uses Lemma 16 and a union bound over all possible choices of $W$.
+Finally, using that $\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$ and the fact that $\text{Bin}(n,p)\preceq\text{Bin}(n^{\prime},p^{\prime})$ whenever $n^{\prime}\geq n$ and $p^{\prime}\geq p$ yields
+[…]
+$=\left(\frac{3e}{4}\cdot n^{\alpha/2-\varepsilon-(1-\delta^{\prime})/2}\right)^{6n^{\delta^{\prime}}}=o(1/n),$
+since $\alpha/2-\varepsilon-(1-\delta^{\prime})/2<0$. Note (a) uses the concentration bound (17) from Lemma 17. This concludes the proof of Lemma 19. $\Box$
+Finally, Lemma 19 directly implies Lemma 18.
+4.3 Relating high degree nodes to the $k$-core: Proof of Lemma 13
+Proof of Lemma 13. For any set $A\subseteq[n]$, let $N_{v}(A)$ denote the number of neighbors of $v$ in $A$, i.e.
+$N_{v}(A):=|\left\{u\in A:\{u,v\}\in E(G)\right\}|.$
+Fix $v\in[n]$. Let $G-v$ denote the induced subgraph of $G$ on the vertex set $[n]-\{v\}$. Let $C_{v}$ denote the $k$-core of $G-v$. Since $C_{v}$ is a subset of the $k$-core of $G$, it follows that
+$\left\{v\notin\mathrm{core}_{k}(G)\right\}\subseteq\left\{N_{v}(\mathrm{core}_{k}(G))\leq k-1\right\}\subseteq\left\{N_{v}(C_{v})\leq k-1\right\}.$
+Let $\delta\in(1-\alpha,1-\alpha+2\varepsilon)$. It follows that
+($\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)\leq q_{1}+q_{2},$ (20)
+where
+$q_{1}:=\mathbb{P}\left(|C_{v}|<n-3n^{\delta}\right)\stackrel{{\scriptstyle\text{(a)}}}{{=}}o(1/n),$
+where (a) follows from Lemma 18, and
+$q_{2}$ :=($:=\mathbb{P}\left(\left\{N_{v}(C_{v})\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\cap\left\{|C_{v}|\geq n-3n^{1-\alpha/2}\right\}\right)$
+$\leq\sum_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(\left\{N_{v}(C_{v})<k\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\mid C_{v}=A\right)\cdot\mathbb{P}\left(C_{v}=A\right)$
+$\leq\max_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(\left\{N_{v}(A)<k\right\}\cap\left\{\delta_{G}(v)\geq k+2/\alpha\right\}\right)$
+$\leq\max_{\begin{subarray}{c}A\subseteq V\\ |A|\geq n-3n^{\delta}\end{subarray}}\mathbb{P}\left(N_{v}(A^{c})\geq 1+2/\alpha\right)$ (21)
+For all sets $A$ of size larger than $n-3n^{\delta}$, we have that
+$\mathbb{P}\left(N_{v}(A^{c})\geq 1+2/\alpha\right)\!=\!\mathbb{P}\left(\sum_{i\in A^{c}}G_{vi}\geq 1+2/\alpha\right)\!\leq\mathbb{P}\left(\text{Bin}(3n^{\delta},n^{\alpha/2-1})\geq 1+2/\alpha\right),$ (22)
+where the last step uses that $\sum_{i\in A^{c}}G_{vi}\preceq\text{Bin}(|A^{c}|,p_{\max})\preceq\text{Bin}(3n^{\delta},n^{\alpha/2-1})$. Combining (21) and (22) yields
+$q_{2}$ $\leq\mathbb{P}\left(\text{Bin}(3n^{\delta},n^{\alpha/2-1})\geq 1+2/\alpha\right)\leq\left(\frac{3\alpha e}{2+\alpha}\cdot n^{-\alpha/2}\right)^{1+2/\alpha}\!=\text{const.}\times n^{-1-\alpha/2}=o(1/n).$
+[…]
+5 Proofs for Main Results
+This section presents proofs for Theorems 1 and 8.
+Section 5.1 presents the intuition behind the analysis of Algorithm 1, collecting useful lemmas along the way.
+These lemmas are proved in Section 5.3, after the proof of Theorem 1 in Section 5.2.
+Finally, the negative result (Theorem 8) is proved in Section 5.5.
+5.1 Analyzing Algorithm 1: Intuition behind the proof of Theorem 1
 Algorithm 1 succeeds if both step 1 and step 2 succeed, i.e.
   1. 1.
-    Each instance of pairwise matching using the $k$-core estimator is correct on its domain, i.e.
-    $\widehat{\mu}_{ij}(v)=\pi^{*}_{ij}(v)\ \forall v\in\mathsf{dom}(\widehat{\mu}_{ij}),\ \forall i,j.$
+    Each instance of pairwise matching using the $k$-core estimator is precise, i.e.
+it is correct on its domain with probability $1-o(1)$, i.e.
+$\widehat{\mu}_{ij}(v)=\pi^{*}_{ij}(v)$ for all $v\in\mathrm{dom}(\widehat{\mu}_{ij})$ for all $i$ and $j$.
   2. 2.
     For each node $v$ and any two graphs $G_{i}$ and $G_{j}$, there is a sequence of graphs such that $v$ can be transitively matched through those graphs between $G_{i}$ and $G_{j}$.
-On step 1
+5.1.1 On step 1
 This falls back to the regime of analyzing the performance of the $k$-core estimator in the setting of two graphs.
-Cullina and co-authors [CKMP19] showed that the $k$-core estimator is precise: For any two correlated graphs $G_{i}$ and $G_{j}$ with $p=C\log(n)/n$ and constant $s$, the $k$-core estimator correctly matches all nodes in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ with probability $1-o(1)$.
-In fact, this is true for any $C>0$ and for any $k\geq 13$ [RS23].
-Therefore, using the fact that the number of instances of pairwise matchings is constant whenever $m$ is constant, a union bound reveals
+Cullina and co-authors (CKMP, 19) showed that the $k$-core estimator is precise: For any two correlated Erdős-Rényi graphs $G_{i}$ and $G_{j}$ with $\mathbf{P}_{ij}=C\log(n)/n$ for all $1\leq i<j\leq n$, and constant subsampling probability $s$, the $k$-core estimator with sufficiently large $k$ correctly matches all nodes in $\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ with probability $1-o(1)$.
+Building on this analysis, Racz and Sridhar (RS, 23) showed that this is true more generally:
+Lemma 20 (Lemma III.4 from (RS, 23))
+Let $G_{1}$ and $G_{2}$ be correlated inhomogeneous graphs obtained by subsampling each edge independently with probability $s$ from a parent graph $G\sim\mathrm{RG}(n,\mathbf{P})$.
+Suppose that there exists some $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}\leq o(n^{\alpha/2-\varepsilon-1})$, and let $k>\frac{12}{1-\alpha+2\varepsilon}$.
+Then, with probability $1-o(1)$, it holds that the matching $\widehat{\mu}_{k}$ output by the $k$-core estimator satisfies
+$\mathrm{dom}(\widehat{\mu}_{k})=\mathrm{core}_{k}(G_{1}^{\prime}\wedge G_{2}^{\prime})\text{ and }\widehat{\mu}_{k}(v)=\pi^{*}(v)\ \forall\ v\in\mathrm{dom}(\widehat{\mu}_{k}).$
+This precision of the $k$-core estimator is critical to the success of Algorithm 1 since it ensures that there are no conflicts during the boosting step.
+Applying Lemma 20 to the setting of $m$ graphs yields that each instance of pairwise matching in step $1$ of Algorithm 1 is indeed precise.
+Corollary 21
+Let $\widehat{\mu}_{ij}$ denote the matching output by the $k$-core estimator on graphs $G_{i}$ and $G_{j}$. Under the conditions of Lemma 20,
+( $\mathbb{P}(\exists\ 1\leq i<j\leq m,\text{ and }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))=o(1).$
+Proof. Since the number of instances of pairwise matchings is constant whenever $m$ is constant, a union bound yields
+( $\mathbb{P}(\exists\ 1$ $\leq i<j\leq m,\text{ and }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))$
+$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\mathbb{P}\left(\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v)\text{ for some }v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right)=o(1).$
 […]
-$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\mathbb{P}\left(\widehat{\mu}_{i,j}(v)\neq\pi^{*}_{i,j}(v)\text{ for some }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right)$
-$=o(1).$
-We have proved the following.
-Proposition 5.
-Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model.
-Let $k\geq 13$ and let $\widehat{\mu}_{ij}$ denote the matching output by the $k$-core estimator on graphs $G_{i}$ and $G_{j}$.
-Then,
-( $\mathbb{P}(\exists\ 1\leq i<j\leq m,\text{ and }v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime}))\text{ such that }\widehat{\mu}_{ij}(v)\neq\pi^{*}_{ij}(v))=o(1).$
-On step 2
-The challenging part of the proof is to show that boosting through transitive closure matches all the nodes with probability $1-o(1)$ if $Cs(1-(1-s)^{m-1})>1$.
+5.1.2 On step 2
+The challenging part of the proof is to show that boosting through transitive closure matches all the nodes with probability $1-o(1)$ under the sufficient condition.
 It is instructive to visualize this using transitivity graphs.
-Definition 6 (Transitivity graph, $\mathcal{H}(v)$).
-For each node $v\in V$, let $\mathcal{H}(v)$ denote the graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that an edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$.
+Definition 22 (Transitivity graph, $\mathcal{H}(v)$)
+For each node $v\in[n]$, let $\mathcal{H}(v)$ denote the graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that an edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$.
 On the event that each instance of pairwise matching using the $k$-core is correct, the edge $\{g_{i},g_{j}\}$ is present in $\mathcal{H}(v)$ if and only if $v$ is correctly matched using the $k$-core estimator between $G_{i}$ and $G_{j}$, i.e.
@@
 Thus, in order for step 2 to succeed (i.e.
-to exactly match all vertices across all graphs), it suffices that the graph $\mathcal{H}(v)$ is connected for each node $v\in V$.
+to exactly match all vertices across all graphs), it suffices that the graph $\mathcal{H}(v)$ be connected for each node $v\in[n]$.
 However, studying the connectivity of the transitivity graphs is challenging because in any graph $\mathcal{H}(v)$, no two edges are independent.
@@
 To overcome this, we introduce another graph ~$\widetilde{\mathcal{H}}(v)$ that relates to $\mathcal{H}(v)$ and is amenable to analysis.
-Definition 7.
-For each node $v\in V$, let ~$\widetilde{\mathcal{H}}(v)$ denote a complete weighted graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that the weight on any edge $\{g_{i},g_{j}\}$ is $\widetilde{c}_{v}\left(i,j\right):=\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v).$
-The relationship between the graphs $\mathcal{H}(v)$ and ~$\widetilde{\mathcal{H}}(v)$ stems from a useful relationship between the degree of node $v$ in $G_{i}^{\prime}\wedge G_{j}^{\prime}$ and the inclusion of $v$ in $\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ for each $i$ and $j$.
-Since this result is of independent interest in the study of random graphs, we state it below for general Erdős-Rényi graphs.
-Lemma 8.
-Let $n$ and $k$ be positive integers and let $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$ for some $\alpha>0$. Let $v$ be a node of $G$ and let $\delta_{G}(v)$ denote the degree of $v$ in $G$. Then,
-($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)=o\left(1/n\right).$ (1)
-For any $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}(n,Cs^{2}\log(n)/n)$.
-Thus, Lemma 8 implies that with probability $1-o(1/n)$, if a pair $\{g_{i},g_{j}\}$ has edge weight $\widetilde{c}_{ij}\geq k+1/\alpha$ in ~$\widetilde{\mathcal{H}}(v)$, then the corresponding edge $\{g_{i},g_{j}\}$ is present in the transitivity graph $\mathcal{H}(v)$.
-Equivalently, $v$ is correctly matched between $G_{i}$ and $G_{j}$ in the instance of pairwise $k$-core matching between them.
+Let $\delta_{G}(v)$ denote the degree of a node $v$ in graph $G$.
+Definition 23
+For each node $v\in[n]$, let ~$\widetilde{\mathcal{H}}(v)$ denote a complete weighted graph on the vertex set $\{g_{1},\cdots,g_{m}\}$ such that the weight on any edge $\{g_{i},g_{j}\}$ is $\widetilde{c}_{v}\left(i,j\right):=\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v).$
+The relationship between the graphs $\mathcal{H}(v)$ and ~$\widetilde{\mathcal{H}}(v)$ stems from Lemma 13 which relates the degree of node $v$ in $G_{i}^{\prime}\wedge G_{j}^{\prime}$ to the inclusion of $v$ in $\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})$ for each $i$ and $j$.
+Since the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathrm{RG}(n,\mathbf{P}s^{2})$Lemma 13 implies that with probability $1-o(1/n)$, if a pair $\{g_{i},g_{j}\}$ has edge weight $\widetilde{c}_{ij}\geq k+2/\alpha$ in ~$\widetilde{\mathcal{H}}(v)$, then the corresponding edge $\{g_{i},g_{j}\}$ is present in the transitivity graph $\mathcal{H}(v)$, i.e.
+$v$ is correctly matched between $G_{i}$ and $G_{j}$ in the instance of pairwise $k$-core matching between them.
 The graph $\mathcal{H}(v)$ is not connected only if it contains a (non-empty) vertex cut $U\subset\{1,\cdots,m\}$ with no edge crossing between $U$ and $U^{c}$.
-Let $c_{v}(U)$ denote the number of such crossing edges in $\mathcal{H}(v)$.
+For a cut $U$, let $c_{v}(U)$ denote the number of such crossing edges in $\mathcal{H}(v)$.
 Furthermore, define the cost of the cut $U$ in ~$\widetilde{\mathcal{H}}(v)$ as
 $\widetilde{c}_{v}(U):=\sum_{i\in U}\sum_{j\in U^{c}}\widetilde{c}_{v}\left(i,j\right).$
-Lemma 8 is a statement about a single graph, but we show it can be invoked to prove the following.
-Theorem 9.
-Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model with parameters $C$ and $s$.
-Let $v\in V$ and let $U$ be a vertex cut of $\{1,\cdots,m\}$ such that $|U|\leq\lfloor m/2\rfloor$.
+Lemma 13 is a statement about a single graph, but it can be invoked to prove the following.
+Lemma 24
+Let $G_{1},\cdots,G_{m}$ be correlated inhomogeneous ranodm graphs from the subsampling model with parameters $\mathbf{P}$ and $s$.
+Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$.
+Let $v\in[n]$ and let $U$ be a vertex cut of $\{1,\cdots,m\}$ such that $|U|\leq\lfloor m/2\rfloor$.
 Then,
-($\mathbb{P}\left(\left\{c_{v}(U)=0\right\}\cap\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}\right)=o(1/n).$ (2)
-It suffices therefore to analyze the probability that the graph ~$\widetilde{\mathcal{H}}(v)$ has a cut $U$ such that its cost $\widetilde{c}_{v}(U)$ is too small.
+($\mathbb{P}\left(\left\{c_{v}(U)=0\right\}\cap\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)\right\}\right)=o(1/n).$ (23)
+The proof of Lemma 24 is deferred to Section 5.3.
+It suffices to analyze the probability that the graph ~$\widetilde{\mathcal{H}}(v)$ has a cut $U$ such that its cost $\widetilde{c}_{v}(U)$ is too small.
 To that end, we show that the bottleneck arises from vertex cuts of small size.
 Formally,
-Theorem 10.
-Let $G_{1},\cdots,G_{m}$ be correlated graphs from the subsampling model.
-Let $v\in V$ and let $U_{\ell}$ denote the set $\{1,\cdots,\ell\}$ for $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$.
+Lemma 25
+Let $G_{1},\cdots,G_{m}$ be correlated inhomogeneous graphs obtained from the subsampling model.
+Let $v\in[n]$ and let $U_{\ell}$ denote the set $\{1,\cdots,\ell\}$ for $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$.
 For any vertex cut $U$ of $\{1,\cdots,m\}$, let $\widetilde{c}_{v}(U)$ denote its cost in the graph ~$\widetilde{\mathcal{H}}(v)$.
@@
 $\widetilde{c}_{v}(U_{1})\preceq\widetilde{c}_{v}(U_{2})\preceq\cdots\preceq\widetilde{c}_{v}(U_{\lfloor m/2\rfloor}).$
-Theorems 9 and 10 imply that the tightest bottleneck to the connectivity of $\mathcal{H}(v)$ is the event that $\widetilde{c}_{v}(U_{1})$ is below the threshold $r:=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$, i.e.
+The proof of Lemma 25 is deferred to Appendix C due to space constraints.
+Lemmas 24 and 25 imply that the tightest bottleneck to the connectivity of $\mathcal{H}(v)$ is the event that $\widetilde{c}_{v}(U_{1})$ is below the threshold $r:=\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)$, i.e.
 the sum of degrees of $v$ over the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$ is less than $r$.
-This event occurs only if the degree of $v$ is less than $r$ in each of the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$.
-However, under the condition $Cs(1-(1-s)^{m-1})>1$, it turns out that this event occurs with probability $o(1/n)$.
-Theorem 11.
-Let $G_{1},\cdots,G_{m}$ be obtained from the subsampling model with parameters $C$ and $s$.
-Let $r=\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)$.
-Let $v\in[n]$ and suppose that $Cs(1-(1-s)^{m-1})>1$.
-Then,
-$\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq r\right)\leq\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\right\}\right)=o\left(1/n\right).$
-4.3 Piecing it all together: Proof of Theorem 3
+This event occurs only if the degree of $v$ is simultaneously less than $r$ in each of the intersection graphs $(G_{1}\wedge G_{j}^{\prime}:j=2,\cdots,m)$.
+The last ingredient to the analysis is that under the condition $\sum_{i=1}^{n}e^{-d_{i}s\left(1-(1-s)^{m-1}\right)}=o(n^{-\alpha})$, this event occurs with probability $o(1/n)$.
+Lemma 26
+Let $G_{1},\cdots,G_{m}$ be obtained from the subsampling model with parameters $\mathbf{P}$ and $s$.
+Let $d_{i}=\sum_{j\neq i}p_{ij}$ and $v^{*}\in\operatorname*{arg\,min}_{v\in V}d_{v}$.
+Let $k\geq\frac{12}{1-\alpha+2\varepsilon}$ be constant and let $r\triangleq\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)$.
+Suppose there exist $\alpha\in(0,1]$ and $\varepsilon>0$ such that $p_{\max}=o(n^{\alpha/2-\varepsilon-1})$ and
+$\sum_{v\in V\setminus v^{*}}e^{-d_{v}s\left(1-(1-s)^{m-1}\right)}=o(n^{-\alpha}).$ (24)
+Then
+$(\star)\triangleq\mathbb{P}\left(\exists\ v\in V\setminus v^{*}:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\big{\}}\right)=o(1).$
+Consequently, under the sufficient condition (1), the transitivity graph $\mathcal{H}(v)$ is connected for each node $v$. The above ideas are used to prove Theorem 1.
+5.2 Piecing it together: Proof of Theorem 1
 Proof of Theorem 1.
-Let $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ denote the output of Algorithm 2 with $k\geq 13$.
-Let $E_{1}$ (resp.
-$E_{2}$) denote the event that Algorithm 1 (resp.
-Algorithm 2) fails to match all $m$ graphs exactly, i.e.
-$E_{1}=\left\{\widehat{\pi}_{12}^{\text{ML}}\neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}^{\text{ML}}\neq\pi^{*}_{1m}\right\},\quad\quad E_{2}=\left\{\widehat{\pi}_{12}\ \neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}\neq\pi^{*}_{1m}\right\}.$
-First, we show that the output of Algorithm 2 is correct with probability $1-o(1)$ whenever $Cs(1-(1-s)^{m-1})>1$.
-If the event $E_{2}$ occurs, then either step 1 failed, i.e.
+Let $\widehat{\pi}_{12},\cdots,\widehat{\pi}_{1m}$ denote the output of Algorithm 1 with $k\geq\frac{12}{1-\alpha+2\varepsilon}$.
+Let $E_{1}$ denote the event that Algorithm 1 fails to match all $m$ graphs exactly, i.e.
+$E_{1}=\left\{\widehat{\pi}_{12}\ \neq\pi^{*}_{12}\right\}\cup\cdots\cup\left\{\widehat{\pi}_{1m}\neq\pi^{*}_{1m}\right\}.$
+First, we show that the output of Algorithm 1 is correct with probability $1-o(1)$ whenever the sufficient condition holds.
+If the event $E_{1}$ occurs, then either step 1 failed, i.e.
 there is a $k$-core matching $\widehat{\mu}_{ij}$ that is incorrect, or step 2 failed, i.e.
@@
 Therefore,
-$\mathbb{P}\left(E_{2}\right)$ $\leq\mathbb{P}\left(\bigcup_{i,j}\ \bigcup_{v\in\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})}\!\!\!\!\left\{\widehat{\mu}_{ij}\neq\pi^{*}_{ij}\right\}\!\right)\!+\mathbb{P}\left(\bigcup_{v\in V}\left\{\mathcal{H}(v)\text{ is not connected}\right\}\!\right)\!\leq o(1)+\!\sum_{v\in V}q_{v},$
-where the last step uses Proposition 5, and $q_{v}$ denotes the probability that the transitivity graph $\mathcal{H}(v)$ is not connected.
+$\mathbb{P}\left(E_{1}\right)$ $\leq\mathbb{P}\left(\bigcup_{i,j}\ \bigcup_{v\in\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})}\!\!\!\!\left\{\widehat{\mu}_{ij}\neq\pi^{*}_{ij}\right\}\!\right)\!+\mathbb{P}\left(\bigcup_{v\in[n]}\left\{\mathcal{H}(v)\text{ is not connected}\right\}\!\right)\!\leq o(1)+q,$
+where the last step uses Corollary 21, and $q$ denotes the probability that a transitivity graph $\mathcal{H}(v)$ is not connected for some $v$ in $[n]$.
 For each $\ell$ in the set $\{1,\cdots,\lfloor m/2\rfloor\}$, let $U_{\ell}$ denote the set $\left\{1,\cdots,\ell\right\}$.
 Then,
-$q_{v}\!$ $=\mathbb{P}\left(\bigcup_{\ell=1}^{\lfloor m/2\rfloor}\left\{\exists\ U\subset\{1,\cdots,m\}:|U|=\ell\text{ and }c_{v}(U)=0\right\}\right)$
-$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\cdot\mathbb{P}\left(c_{v}(U_{\ell})=0\right)$
-$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!\mathbb{P}\left(\left\{c_{v}(U_{\ell})=0\right\}\cap\bigg{\{}\widetilde{c}_{v}(U_{\ell})>\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}\right)\right]$
-$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$
-$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\left[\mathbb{P}\left(\widetilde{c}_{v}(U_{1})\leq\frac{m^{2}}{4}\bigg{(}k\!+\!\frac{1}{Cs^{2}}\bigg{)}\!\right)\!+\!o\left(\frac{1}{n}\right)\right]$
-$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\left[o\left(\frac{1}{n}\right)+o\left(\frac{1}{n}\right)\right]=o\left(\frac{1}{n}\right).$
-Here, (a) uses Theorem 9, and (b) uses the fact that for any $\ell\geq 2$, the random variable $\widetilde{c}_{v}(U_{\ell})$ stochastically dominates $\widetilde{c}_{v}(U_{1})$ (Theorem 10).
-Finally, (c) uses Theorem 11 and the fact that $Cs(1-(1-s)^{m-1})>1$.
-Therefore, a union bound over all the nodes yields
-$\mathbb{P}\left(E_{2}\right)\leq o(1)+\sum_{v\in V}q_{v}\leq o(1)+n\times o(1/n)=o(1).$
-Finally, by optimality of the MLE, it follows that
-$\mathbb{P}\left(E_{1}\right)\leq\mathbb{P}\left(E_{2}\right)=o(1),$
-whenever $Cs(1-(1-s)^{m-1})>1$. This concludes the proof. ∎
-5 Discussion and Future Work
-In this work, we introduced and analyzed matching through transitive closure - an approach that combines information from multiple graphs to recover the underlying correspondence between them.
-Despite its simplicity, it turns out that matching through transitive closure is an optimal way to combine information in the setting where the graphs are pairwise matched using the $k$-core estimator.
-A limitation of our algorithms is the runtime: Algorithm 2 does not run in polynomial time because it uses the $k$-core estimator for pairwise matching, which involves searching over the space of permutations.
+$q\!$ $=\mathbb{P}\left(\bigcup_{v\in[n]}\bigcup_{\ell=1}^{\lfloor m/2\rfloor}\left\{\exists\ U\subset\{1,\cdots,m\}:|U|=\ell\text{ and }c_{v}(U)=0\right\}\right)$
+$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!\binom{m}{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:c_{v}(U_{\ell})=0\right)$
+$\leq\!\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\Bigg{[}\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\big{(}k+2/\alpha\big{)}\right)$
+$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\mathbb{P}\left(\exists\ v\in[n]:\left\{c_{v}(U_{\ell})=0\right\}\cap\big{\{}\widetilde{c}_{v}(U_{\ell})>\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\big{\}}\right)\Bigg{]}$
+$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\!\sum_{\ell=1}^{\lfloor m/2\rfloor}\!m^{\ell}\left[\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{\ell})\leq\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\!\right)\!+\!o\left(1\right)\right]$
+$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\!o(1)+\sum_{\ell=1}^{\lfloor m/2\rfloor}\!m^{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:\widetilde{c}_{v}(U_{1})\leq\frac{m^{2}}{4}\big{(}k\!+\!2/\alpha\big{)}\!\right)$
+$\stackrel{{\scriptstyle\text{(c)}}}{{\leq}}\!o(1)+\sum_{\ell=1}^{\lfloor m/2\rfloor}m^{\ell}\cdot\mathbb{P}\left(\exists\ v\in[n]:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)\leq r\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)\leq r\big{\}}\right)\stackrel{{\scriptstyle\text{(d)}}}{{\leq}}o(1)$
+Here, (a) uses Lemma 24 and a union bound, and (b) uses the fact that for any $\ell\geq 2$, the random variable $\widetilde{c}_{v}(U_{\ell})$ stochastically dominates $\widetilde{c}_{v}(U_{1})$ (Lemma 25).
+Lastly, (c) uses the definition of $\widetilde{c}_{v}(U_{1})$ and (d) uses Lemma 26.
+Therefore, transitive closure successfully matches all nodes in $[n]$ with probability $1-o(1)$.
+It follows that $\mathbb{P}\left(E_{2}\right)=o(1)$.
+This concludes the proof.
+$\Box$
+5.3 Proofs of Lemmas 24 and 26
+5.3.1 Proof of Lemma 24.
+For any vertex cut $U$,
+$\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{2}{\alpha}\right)\right\}$ $\stackrel{{\scriptstyle\text{(a)}}}{{\subseteq}}\left\{\widetilde{c}_{v}(U)>|U|\left(m-|U|\right)\left(k+\frac{2}{\alpha}\right)\right\}$
+=$=\bigg{\{}\sum_{i\in U}\sum_{j\in U^{c}}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>|U|(m-|U|)\bigg{(}k+\frac{2}{\alpha}\bigg{)}\bigg{\}}$
+$\subseteq\bigcup_{i\in U}\bigcup_{j\in U^{c}}\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{2}{\alpha}\right\},$
+where (a) uses the fact that the maximum of a set of numbers is greater than or equal to the average. On the other hand,
+$\left\{c_{v}(U)=0\right\}=\bigcap_{i\in U}\bigcap_{j\in U^{c}}\left\{v\notin\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}.$
+Let $p_{1}$ denote the probability in the LHS of (23). It follows from the union bound that
+$p_{1}\leq\sum_{i\in U}\sum_{j\in U^{c}}\mathbb{P}\left(\left\{v\notin\mathrm{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}\cap\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{2}{\alpha}\right\}\right)=o(1/n),$
+[…]
+5.3.2 Proof of Lemma 26
+If $s=1$, then the graphs $G_{1},\cdots,G_{m}$ are all isomorphic, and
+$(\star)=\mathbb{P}\left(\exists v\in V\setminus v^{*}:\delta_{G_{1}}(v)\leq r\right)\leq\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(\delta_{G_{1}}(v)=0\right)\leq\sum_{i=1}^{n}e^{-d_{i}}=o(1),$
+and so the result follows. We therefore assume without loss of generality that $s<1$. Observe that
+(($\star$)( ⋆ ) $\leq\mathbb{P}\left(\bigcup_{r_{2}=0}^{r}\cdots\bigcup_{r_{m}=0}^{r}\left(\exists\ v\in V\setminus v^{*}:\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\big{\}}\right)\right)$
+$\leq\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\ \ \sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\mathbb{P}\left(\big{\{}\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\big{\}}\cap\cdots\cap\big{\{}\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\big{\}}\mid\delta_{G_{1}}(v)=D_{v}\right)\right]$
+=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\ \ \sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}\binom{D_{v}}{r_{\ell}}s^{r_{\ell}}(1-s)^{D_{v}-r_{\ell}}\cdot\bm{1}_{\{r_{\ell}\leq D_{v}\}}\right]$
+$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\sum_{r_{2}=0}^{r}\cdots\sum_{r_{m}=0}^{r}\left[\prod_{\ell=2}^{m}\left(\frac{e}{r_{\ell}}\cdot\frac{s}{1-s}\right)^{r_{\ell}}\right]\cdot\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right],$ (25)
+where (a) uses that the degrees of node $v$ in the graphs $G_{1}\wedge G_{2}^{\prime},\cdots,G_{1}\wedge G_{m}^{\prime}$ are conditionally independent of the degree of $v$ in $G_{1}$, and that each edge of $G_{1}$ is independently present in $G_{1}\wedge G_{j}^{\prime}$ with probability $s$.
+Further, (b) uses that $\binom{n}{k}\leq\left(ne/k\right)^{k}$ with the implicit definition $(1/0)^{0}\triangleq 1$.
+Note that
+$\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right]$ $=\sum_{v\in V\setminus v^{*}}\left(\sum_{d=0}^{n-1}\mathbb{P}\left(D_{v}=d\right)\prod_{\ell=2}^{m}d^{r_{\ell}}(1-s)^{d}\right)$
+$=\sum_{d=0}^{n-1}\left(\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(D_{v}=d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\right).$ (26)
+Proceed by splitting the outer summation in (26) over $d$ at $d^{*}\triangleq n^{\frac{\alpha}{2(m-1)r}}$. We have
+$\sum_{d=0}^{d^{*}}\sum_{v\in V\setminus v^{*}}$ $\mathbb{P}\left(D_{v}=d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\leq\sum_{v\in V\setminus v^{*}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{d=0}^{d^{*}}\mathbb{P}\left(D_{v}=d\right)\cdot(1-s)^{(m-1)d}$
+$\leq(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[(1-s)^{(m-1)D_{v}}\right]$
+=$\stackrel{{\scriptstyle\text{(c)}}}{{=}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\prod_{u\neq v}\left(1-p_{uv}s+p_{uv}s(1-s)^{m-1}\right)$
+$\stackrel{{\scriptstyle\text{(d)}}}{{\leq}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}\sum_{v\in V\setminus v^{*}}\exp\bigg{(}-\sum_{u\neq v}p_{uv}s\left(1-(1-s)^{m-1}\right)\bigg{)}$
+$\stackrel{{\scriptstyle\text{(e)}}}{{\leq}}n^{\alpha/2}\cdot\sum_{v\in V\setminus v^{*}}e^{-d_{v}s\left(1-(1-s)^{m-1}\right)}=o(1).$ (27)
+Here, (c) uses the moment generating function of $D_{v}$, which is distributed as a sum of independent Bernoulli random variables with means $p_{uv}$ respectively.
+Further, (d) uses that $1-x\leq e^{-x}$ for any $x$, and finally (e) uses the definition of $d^{*}$ and the condition (24).
+The other part of the split sum can be bounded as
+$\sum_{d=d^{*}}^{n-1}\sum_{v\in V\setminus v^{*}}$ $\mathbb{P}\left(D_{v}\!=\!d\right)\cdot d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\leq\sum_{v\in V\setminus v^{*}}\left[\max_{d\in[d^{*},n]}d^{\sum_{\ell=2}^{m}r_{\ell}}\cdot(1-s)^{(m-1)d}\right]\mathbb{P}\left(D_{v}\geq d^{*}\right)$
+$\stackrel{{\scriptstyle\text{(f)}}}{{\leq}}(d^{*})^{\sum_{\ell=2}^{m}r_{\ell}}(1-s)^{(m-1)d^{*}}\cdot\sum_{v\in V\setminus v^{*}}\mathbb{P}\left(D_{v}\geq d^{*}\right)$
+$\stackrel{{\scriptstyle\text{(g)}}}{{\leq}}n^{1+\alpha/2}\cdot\left((1-s)^{m-1}\right)^{n^{\frac{\alpha}{2(m-1)r}}}=o(1),$ (28)
+whenever $s>0$.
+Here, (f) uses that the function $d\mapsto d^{\sum r_{i}}(a-s)^{md}$ is decreasing on the interval $[d^{*},n]$ for all sufficiently large $n$.
+Finally, (g) uses the definition of $d^{*}$ and that $\mathbb{P}\left(D_{v}\geq d^{*}\right)\leq 1$ for all $v$.
+Plugging (27) and (28) in (26) yields that $\sum_{v\in V\setminus v^{*}}\mathbb{E}_{D_{v}}\left[\prod_{\ell=2}^{m}D_{v}^{r_{\ell}}(1-s)^{D_{v}}\right]=o(1)$.
+Therefore, continuing from (25),
+[…]
+$\leq(r+1)(m-1)\cdot\max_{1\leq r^{\prime}\leq r}\left(\frac{es}{r^{\prime}(1-s)}\right)^{r^{\prime}(m-1)}\cdot o(1)=o(1),$
+since $r$ and $m$ are both constants independent of $n$. This completes the proof. $\Box$
+5.4 Proof of Theorem 5
+In this section, we prove Theorem 5, which implies that our sufficient condition (2) is not necessary for matching inhomogeneous graphs in general.
+Proof of Theorem 5.
+The proof is nonconstructive; the idea is to show that recovery is possible if the parameter matrix $P$ is random with a suitable distribution, and then show that a set of deterministic choices for $P$ satisfying the requirements of the proposition has positive probability (actually probability converging to one) under the distribution of $P.$ Let $G_{-1}$ denote an Erdős-Rényi graph with parameters $n$ and $q$, where $q$ is to be determined later.
+Let $P$ denote the random matrix equal to $p_{\max}$ times the incidence matrix of $G_{-1}.$
+Given $P$, let $G\sim\mathrm{RG}(n,P)$, let $G^{\prime}_{1},\cdots G^{\prime}_{m}$ denote the graphs obtained from $G$ by subsampling each edge with probability $s$, and let $G_{k}$ be obtained from $G^{\prime}_{k}$ by randomly uniformly permuting the labels, for $k\in[m]$.
+We will consider the algorithm that separately matches $G_{-1}$ to $G_{k}$ for each $k$, and then matches $G_{1},\cdots,G_{m}$ to each other via transitive closure through $G_{-1}$.
+For $k\in[m]$, $G^{\prime}_{k}$ can be viewed as being obtained from $G_{-1}$ by independently subsampling the edges of $G_{-1}$ with probability $p_{\max}s.$ Therefore, if $r_{ij}\triangleq\mathbb{P}\left(\bm{1}_{\{e\in E(G_{-1})\}}=i,\bm{1}_{\{e\in E(G_{k})\}}=j\right)$, then
+$\left(\begin{array}[]{cc}r_{00}{}{}&r_{01}\\ r_{10}{}{}&r_{11}\end{array}\right)=\left(\begin{array}[]{cc}1-q&~{}~{}0\\ q(1-p_{\max}s)&~{}~{}qp_{\max}s\end{array}\right)$
+It is known (CK, 17; WXY, 22) for correlated Erdős-Rényi graphs such as this that exact recovery is possible if there is some $\varepsilon^{\prime}>0$ such that
+$n\left(\sqrt{r_{00}r_{11}}-\sqrt{r_{01}r_{10}}\right)^{2}\geq(1+\varepsilon^{\prime})\log n.$
+Assume without loss of generality that $\epsilon<1.$ Let $\epsilon^{\prime}=\epsilon/4$ and select $q$ so that $qs=\frac{1+2\epsilon^{\prime}}{\log n}.$ With that choice we conclude that the algorithm described exactly matches $G_{k}$ to $G_{-1}$ with probability $1-o(1)$ for each $k$, and hence matches $G_{1},\cdots,G_{m}$ to each other with probability $1-o(1).$
+Let $\delta_{G_{-1}}(i)$ denote the degree of node $i$ in graph $G_{-1}$.
+Note that $\delta_{G_{-1}}(i)\sim\text{Bin}(n-1,q)\preceq\text{Bin}(n,q)$ for all nodes $i$.
+Therefore by the union bound and Chernoff inequality [Theorem 4.4(MU, 17)],
+$\mathbb{P}\left(\bigcup_{i=1}^{n}\left\{\delta_{G_{-1}}(i)>(1+\varepsilon^{\prime})nq\right\}\right)\leq n\cdot\mathbb{P}\left(\text{Bin}\left(n,q\right){>}(1+\varepsilon^{\prime})nq\right)\leq ne^{-nq(\varepsilon^{\prime})^{2}/3}=o(1).$ (29)
+Let $D_{i}=p_{\max}\cdot\delta_{G_{-1}}(i)$ for $i\in[n],$ so that $D_{i}$ is the random version of $d_{i}.$ Then (3), (29), and the fact $(1+\epsilon^{\prime})(1+2\epsilon^{\prime})<1+\epsilon$ imply that $\mathbb{P}\left(D_{i}\leq d~{}~{}\mbox{for}~{}i\in[n]\right)=1-o(1).$ In summary, the intersection of the following three events has probability $1-o(1):$
+  1. 1.
+    $\max_{i,j}P_{ij}=p_{\max}=(\log n)^{2}/n$.
+  2. 2.
+    $G_{1},\cdots,G_{m}$ are exactly matched by the transitive closure algorithm via $G_{-1}$.
+  3. 3.
+    $D_{i}\leq d~{}~{}\mbox{for}~{}i\in[n]$.
+[…]
+5.5 On Necessary Conditions: Proofs of Theorem 8
+In this section, we prove Theorem 8, i.e.
+our sufficient condition is also necessary for the homogeneous case.
+Theorem 8 has a simple proof following a genie-aided converse argument: reduce the problem to that of matching two graphs by providing extra information to the estimator.
+[…]
+6 Simulation Results and Discussion
+A result of Rácz and Sridhar (RS, 23) establishes that with high probability, the $k$-core estimator for two inhomogeneous graphs matches only the nodes in the $k$-core of their true intersection graph.
+This is true even if the matrix $\mathbf{P}$ has constant entries that do not depend on $n$.
+Although the $k$-core estimator itself is not efficient to implement, its output can be efficiently simulated by computing the $k$-core of the true intersection graph.
+This is possible only because the simulator has access to the ground truth.
+Under the assumption that the $k$-core estimator outputs the $k$-core of the true intersection graph, we can simulate the performance of Algorithm 1.
+Figure 3 shows the mean fraction of matched nodes as a function of $m$ before and after transitive closure, for a variety of graph models.
+For each value of $m$, the $\binom{m}{2}$ pairs of graphs are pairwise matched by simulating the $k$-core estimator, and the average fraction of matched nodes is recorded.
+The transitive closure subroutine is then applied to the partial matchings and the improvement in the fraction of matched nodes is recorded.
+In all figures, the shaded region represents the standard deviation for the fraction of matched nodes.
+Three settings are studied:
+(a) Erdős-Rényi graphs, $(n,p,s,k)=(10^{4},0.003,0.8,13)$
+(b) SBM: $(n,p,q,s,k)=(10^{4},0.04,0.01,0.25,14)$
+(c) RGG, $(n,p,r,s,k)=(10^{4},0.15,0.2,0.4,14)$
+Figure 3: Mean fraction of correctly matched nodes for various graphs obtained by simulating Algorithm 1
+  1. 1.
+    Erdős-Rényi graphs: Figure 3(a) studies ER graphs on $10^{4}$ nodes.
+An edge is present in the parent graph with probability $p=0.003$ and subsampled independently to the graphs $G_{1}$ and $G_{2}$ with probability $0.8$.
+  2. 2.
+    Stochastic Block Model: Figure 3(b) similarly presents the fraction of matched nodes before and after transitive closure for a balanced $5$-community stochastic block model on $10^{4}$ nodes, with intra-community edge probability $p=0.04$ and inter-community edge probability $q=0.01$ respectively.
+The subsampling probability $s=0.25$.
+  3. 3.
+    Random Geometric Graph: Figure 3(c) considers random geometric graphs generated as follows: First, $n=10^{4}$ points are sampled independently and uniformly on the unit square $[0,1]\times[0,1]$.
+The parent graph $G$ is generated by connecting two points within Euclidean distance $r=0.2$ with probability $p=0.15$, and the graphs $G_{1}$ and $G_{2}$ are obtained by subsampling each edge independently with probability $s=0.4$.
+All data points are obtained by averaging across $50$ independent runs of the algorithm. As Figure 3 shows, the transitive closure step boosts the fraction of matched nodes across any two graphs.
+A limitation of Algorithm 1 is the runtime – it does not run in polynomial time because it uses the $k$-core estimator for pairwise matching, which involves searching over the space of permutations.
 Even so, it is useful to establish the fundamental limits of exact recovery, and serve as a benchmark to compare the performance of any other algorithm.
-The transitive closure subroutine (Step 2) itself is efficient because it runs in polynomial time $O(mn)$.
+The transitive closure subroutine (Step 2) itself is efficient because it runs in polynomial time.
 Therefore, a natural next step is to modify Step 1 in our algorithm so that the pairwise matchings are done by an efficient algorithm.
@@
 For example, there is a possibility that the pairwise matchings resulting from the efficient algorithm are heavily correlated, and transitive closure is unable to boost them.
-In Figure 3, we show experimentally that this is not the case for two algorithms of interest: GRAMPA [FMWX22] and Degree Profiles [DMWX21].
-  1. 1.
-    GRAMPA is a spectral algorithm that uses the entire spectrum of the adjacency matrices to match the two graphs. The code is available in [FMWX20].
-  2. 2.
-    Degree Profiles associates with each node a signature derived from the degrees of its neighbors, and matches nodes by signature proximity. The code is available in [DMWX20].
-Evidently, both algorithms benefit substantially from using transitive closure to boost the number of matched nodes.
-This suggests that transitive closure can be a practical algorithm to boost matchings between networks by using other networks as side-information.
-Unfortunately, both GRAMPA and Degree Profiles require the graphs to be close to isomorphic in order to perform well, and so they do not perform well when the model parameters are close to the information theoretic threshold for exact recovery.
-Subsequently, they cannot be used to answer the question in 1.
+7 Conclusion
+In this work, we introduced and analyzed matching through transitive closure - a black-box algorithm to combine information from multiple graphs to recover the underlying correspondence between them.
+Despite its simplicity, it turns out that matching through transitive closure is an optimal way to combine information in the homogeneous setting where the graphs are pairwise matched using the $k$-core estimator.
+The analysis of this algorithm yielded a sufficient condition to exactly match $m$ inhomogeneous random graphs.
 Our work presents several directions for future research.
-Figure 3: Matching through transitive closure
+  • Impossibility results for inhomogeneous graphs.
+The derivation and analysis of the maximum likelihood estimator for exact recovery can help quantify the sharpness of the sufficient condition (2).
+Can the heterogeneity in the model be exploited to match two graphs even when their true intersection graph is disconnected?
   • Polynomial-time algorithms.
 Using a polynomial-time estimator in place of the $k$-core estimator in Step 1 of Algorithm 1 yields a polynomial-time algorithm to match $m$ graphs.
-It is critical that the estimator in question is able to identify for itself the nodes that it has matched correctly - this precision is present in the $k$-core estimator and enables the transitive closure subroutine to work correctly.
+It is critical that this estimator is precise.
 Can the performance guarantees of the $k$-core estimator be realized through polynomial time algorithms that meet this constraint?
-  • Beyond Erdős-Rényi graphs.
-The study of matching two ER graphs provided tools and techniques that extended to the analysis of more realistic models.
-For instance, the $k$-core estimator itself played a crucial role in establishing limits to matching two correlated stochastic block models [GRS22] and two inhomogeneous random graphs [RS23].
-Can the techniques developed in the present work be used to identify the information theoretic limits to exact recovery in these models in the general setting of $m$ graphs?
   • Boosting for partial recovery.
@@
 Is it possible to quantify the extent to which an adversary may perturb edges in some of the graphs without losing the performance guarantees of the matching algorithm?
-Algorithms that perform well on models such as ER graphs and are further generally robust are expected to also work well with real-world networks.
-Acknowledgments and Disclosure of Funding
-This work was supported by NSF under Grant CCF 19-00636.
-Appendix A Maximum Likelihood Estimator
-Recall the form of the maximum likelihood estimator as claimed in Theorem 4. See 4
-Proof.
-Notice that
+Algorithms that perform well on models such as inhomogeneous graphs and are further generally robust are expected to also work well with real-world networks.
+Appendix A Maximum Likelihood Estimator for Exactly Matching Multiple Homogeneous Graphs
+We show that in the homogeneous setting, the maximum likelihood estimator is given by the permutation profile that minimizes the number of edges in the corresponding union graph, i.e.
+$(\widehat{\pi}^{\text{MLE}}_{12},\cdots,\widehat{\pi}^{\text{MLE}}_{1m})\in\operatorname*{arg\,min}_{\pi_{12},\cdots,\pi_{1m}}|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|.$
+Lemma 27
+Let $G_{1},\cdots,G_{m}$ be correlated Erdős-Rényi graphs obtained from the subsampling model with parameters $p$ and $s$.
+Further, let $\pi_{12},\cdots,\pi_{1m}$ denote a collection of permutations on $[n]$.
+Then
+$\log\mathbb{P}\left(G_{1},\cdots,G_{m}\mid\pi^{*}_{12}=\pi_{12},\cdots,\pi^{*}_{1m}=\pi_{1m}\right)\propto\mathrm{const}.-|E(G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}})|,$
+where $\mathrm{const}.$ depends only on $p$, $s$ and $G_{1},\cdots,G_{m}$.
+Proof. Consider the setting where $p_{ij}=p$ for all $i\neq j$. Notice that
 $\mathbb{P}\left(G_{1},\cdots,G_{m}|\pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)$ $=\prod_{e\in\binom{[n]}{2}}\mathbb{P}\left(G_{1}(e),G_{2}(\pi^{*}_{12}(e))\cdots,G_{m}(\pi^{*}_{1m}(e))\mid\pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)$
@@
 $\log\left(\mathbb{P}\left(G_{1},\cdots,G_{m}\ |\ \pi^{*}_{12},\cdots,\pi^{*}_{1m}\right)\right)\propto\text{const.}+\alpha_{0},$
-i.e.
-maximizing the likelihood corresponds to selecting $\pi_{12},\cdots,\pi_{1m}$ to maximize $\alpha_{0}$ the number of node pairs $e$ for which $G_{1}(e)=G_{2}(\pi_{12}(e))=\cdots=G_{m}(\pi_{1m}(e))=0$.
-This is equivalent to minimizing the number of edges in the union graph $G_{1}\vee G_{2}^{\pi_{12}}\vee\cdots\vee G_{m}^{\pi_{1m}}$, as desired.
-∎
-Remark 12.
+[…]
+Remark 28
 In the case of two graphs, minimizing the number of edges in the union graph $G_{1}\vee_{\pi}G_{2}$ is equivalent to maximizing the number of edges in the intersection graph $G_{1}\wedge_{\pi}G_{2}$.
-This is consistent with existing literature on two graphs [CK16, WXY22].
-Appendix B Concentration Inequalities for Binomial Random Variables
-The following bounds for the binomial distribution are used frequently in the analysis.
-Lemma 13.
-Let $X\sim\mathsf{Bin}(n,p)$. Then,
-  1. 1.
-    For any $\delta>0$,
-    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{np}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)np}.$ (5)
-  2. 2.
-    For any $\delta>5$,
-    $\mathbb{P}\left(X\geq(1+\delta)np\right)\leq 2^{-(1+\delta)np}.$ (6)
-  3. 3.
-    For any $\delta\in(0,1)$,
-    $\mathbb{P}\left(X\leq(1-\delta)np\right)\leq\left(\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right)^{np}.$ (7)
+This is consistent with existing literature on two graphs [5, 45].
+Appendix B On negative results for exact recovery with two graphs
+We build up to the proof of Lemma 9 by presenting first a useful lemma about inhomogeneous random graphs.
+Lemma 29
+Let $G\sim\mathrm{RG}(n,\mathbf{Q})$ be an inhomogeneous random graph, and let $N\triangleq\binom{K}{2}$, where $K$ is the number of isolated nodes in $G$. Then,
+$\mathbb{P}\left(N\geq 1\right)\geq\exp(-(2n-4)p_{\max}^{2}-6p_{\max})\cdot\frac{\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}}{4(\mu^{2}+\mu+1)},$ (32)
+where $\mu\triangleq\sum_{i=1}^{n}e^{-d_{i}}$.
 Proof.
-All proofs follow from the Chernoff bound and can be found, or easily derived, from Theorems 4.4 and 4.5 of [MU17]. ∎
-Appendix C Proof of Lemma 8
-We restate Lemma 8 for convenience. See 8
-Before presenting the proof, we present the intuition behind it.
-The events $\{v\notin\mathsf{core}_{k}(G)\}$ and $\left\{\delta_{G}(v)\geq k+1/\alpha\right\}$ are highly negatively correlated.
-However, consider the subgraph $(G-v)$ of $G$ induced on the vertex set $V-\{v\}$, and note that the $k$-core of this subgraph does not depend on the degree of $v$.
-Furthermore, if $v\notin\mathsf{core}_{k}(G)$, then it must be that $v$ has fewer than $k$ neighbors in $\mathsf{core}_{k}(G-v)$.
-Intuitively, this event has low probability if $\mathsf{core}_{k}(G-v)$ is sufficiently large.
-Notice that $(G-v)\sim\mathsf{ER}(n-1,\alpha\log(n)/n)$, and so standard results about the size of the $k$-core of Erdős-Rényi graphs apply.
-However, we require the error probability that the $k$-core of $G-v$ is too small to be $o(1/n)$ this is crucial since we will later use a union bound over all the nodes $v$.
-Unfortunately, standard results such as [Łuc91] can only be invoked directly to show that the corresponding probability is $o(1)$, which is insufficient for our purpose.
-Later in this section, we refine the analysis in [Łuc91] to obtain the desired convergence rate.
-The refinement culminates in the following.
-Lemma 14.
-Let $\alpha>0$ and $G\sim\mathsf{ER}(n,\alpha\log(n)/n)$. Let $v$ be a node of $G$. The size of the $k$-core of $G-v$ satisfies
-$\mathbb{P}\left(\left|\mathsf{core}_{k}(G-v)\right|<n-3n^{1-\alpha}\right)=o(1/n).$
-The proof of Lemma 14 is deferred to Section C.1.
-It remains to study the error event that $v$ has too few neighbors in $\mathsf{core}_{k}(G-v)$.
-To count the number of neighbors of $v$ in $\mathsf{core}_{k}(G-v)$, we exploit the independence of $\mathsf{core}_{k}(G-v)$ and $v$ as follows: each neighbor of $v$ is considered a success if it belongs to $\mathsf{core}_{k}(G-v)$ and a failure otherwise.
-Counting the number of successes is equivalent to sampling with replacement $\delta_{G}(v)$ elements, each of which is independently a success with probability $|\mathsf{core}_{k}(G-v)|/\left(n-1\right)$.
-The number of successes then follows precisely a hypergeometric distribution.
-This intuition is made rigorous in the proof below.
-Let us recall some facts about the hypergeometric distribution because it plays an important role in the proof.
-Denote by $\mathsf{HypGeom}(N,K,n)$ a random variable that counts the number of successes in a sample of $n$ elements drawn without replacement from a population of $N$ individuals, of which $K$ elements are considered successes.
-Note that if this sampling were done with replacement, then the number of successes would follow a $\mathsf{Bin}\left(n,K/N\right)$ distribution.
-A result of Hoeffding [Hoe94] establishes that the $\mathsf{HypGeom}(N,K,n)$ distribution is convex-order dominated by the $\mathsf{Bin}\left(n,K/N\right)$ distribution, i.e.
-$\mathbb{E}\left[f(\mathsf{HypGeom}(N,K,n))\right]\leq\mathbb{E}\left[f(\mathsf{Bin}(n,K/N))\right]\ \text{ for all convex functions $f$}.$
-In particular, the function $f(x)=e^{tx}$ is convex for any value of $t$, and so Chernoff bounds that hold for the binomial distribution also hold for the corresponding hypergeometric distribution.
-This yields the following proposition.
-Proposition 15.
-Let $X\sim\mathsf{HypGeom}(N,K,n)$. It follows for any $\delta>0$ that
-$\mathbb{P}\left(X>(1+\delta)\times\frac{nK}{N}\right)\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{nK/N}\leq\left(\frac{e}{1+\delta}\right)^{(1+\delta)nK/N}.$
-Our final remark about the hypergeometric distribution is a symmetry property. By interchanging the success and failure states, it follows that
-$\mathbb{P}\left(\mathsf{HypGeom}(N,K,n)=k\right)=\mathbb{P}\left(\mathsf{HypGeom}(N,N-K,n)=n-k\right).$
-The above intuition for the proof of Lemma 8 is formalized below.
-Proof of Lemma 8.
-Let $V$ denote the vertex set of $G$, and let $G-v$ denote the induced subgraph of $G$ on the vertex set $V-\{v\}$.
-For any set $A\subseteq V$, let $N_{v}(A)$ denote the set of neighbors of $v$ in the set $A$, i.e.
-$N_{v}(A):=\left\{u\in A:\{u,v\}\in E(G)\right\}.$
-Since $\mathsf{core}_{k}(G-v)\subseteq\mathsf{core}_{k}(G)$, it is true that
-$\left\{v\notin\mathsf{core}_{k}(G)\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G))|\leq k-1\right\}\subseteq\left\{|N_{v}(\mathsf{core}_{k}(G-v))|\leq k-1\right\}.$
-It follows that
-($\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G)\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\right)\leq p_{1}+p_{2},$
-where
-$p_{1}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right\}\right)$
-$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
-It suffices to show that both $p_{1}$ and $p_{2}$ are $o(1/n)$. The term $p_{1}$ deals with the probability that the $k$-core of $G-v$ is too small. In fact, by Lemma 14, it follows directly that
-$p_{1}$ $\leq\mathbb{P}\left(|\mathsf{core}_{k}(G-v)|<n-3n^{1-\alpha}\right)=o(1/n),$
-Next, the probability $p_{2}$ is analyzed. Enumerate arbitrarily but independently the elements of sets $N_{v}(V)$ and $\mathsf{core}_{k}(G-v)$, so that
-$N_{v}(V)=\left\{v_{1},\cdots,v_{\delta_{G}(v)}\right\},\quad\mathsf{core}_{k}(G-v)=\left\{a_{1},\cdots,a_{|\mathsf{core}_{k}(G-v)|}\right\}.$
-Given that $N_{v}(V)$ has more than $k+1/\alpha$ nodes and $\mathsf{core}_{k}(G-v)$ has more than $n-3n^{1-\alpha}$ nodes, it is true that
-$N_{v}(\mathsf{core}_{k}(G-v))$ $=N_{v}(V)\cap\mathsf{core}_{k}(G-v)$
-$\supseteq\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}\cap\left\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\right\}=:\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v)).$
-In words, $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ counts among the first $\lceil k+1/\alpha\rceil$ neighbors of $v$ those nodes that are also in the first $\lceil n-3n^{1-\alpha}\rceil$ nodes of $\mathsf{core}_{k}(G-v)$.
+Let $Z_{i}$ denote the indicator random variable that node $i$ is isolated in $G$.
+Further, let $\widetilde{Z}_{i}\sim\mathrm{Bern}(e^{-d_{i}})$ denote a collection of independent random variables and define $\widetilde{N}\triangleq\sum_{1\leq i<j\leq n}\widetilde{Z}_{i}\cdot\widetilde{Z}_{j}$.
+We use the second moment method by relating the moments of $N$ and $\widetilde{N}$.
+Bounding the second moment. Notice that
+$\mathbb{E}[N^{2}]$ $=\sum_{1\leq i<j\leq n}\ \sum_{1\leq k<\ell\leq n}\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]\,,$ (33)
+$\mathbb{E}[\widetilde{N}^{2}]$ $=\sum_{1\leq i<j\leq n}\ \sum_{1\leq k<\ell\leq n}\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}].$ (34)
+Consider the following three exhaustive cases.
+  • Case $1$: $i=k$ and $j=\ell$. Notice that
+    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}^{2}]\cdot\mathbb{E}[\widetilde{Z}_{j}^{2}]=e^{-d_{i}-d_{j}}.$
     Therefore,
-$p_{2}$ =($=\mathbb{P}\left(\left\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\right\}\cap\left\{\delta_{G}(v)\geq\ k+1/\alpha\right\}\cap\left\{|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right\}\right)$
-($\leq\mathbb{P}\left(\{N_{v}(\mathsf{core}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$
-($\leq\mathbb{P}\left(\{\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))\leq k-1\}\ \big{|}\ \delta_{G}(v)\geq\ k+1/\alpha,\ |\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}\right)$ (8)
-Note that $\mathsf{core}_{k}(G-v)$ is entirely determined by the graph $G-v$, i.e.
-it is independent of the neighbors of $v$.
-Consequently, the two sets $\left\{v_{1},\cdots,v_{\lceil k+1/\alpha\rceil}\right\}$ and $\{a_{1},\cdots,a_{\lceil n-3n^{1-\alpha}\rceil}\}$ are selected independent of each other.
-Equivalently, given that $|\mathsf{core}_{k}(G-v)|\geq n-3n^{1-\alpha}$ and $\delta_{G}(v)\geq k+1/\alpha$, the size of the intersection set $\widetilde{N_{v}}(\widetilde{\mathsf{core}}_{k}(G-v))$ follows a hypergeometric distribution with parameters $(n-1,\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)$.
+    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]\!=\!(1-p_{ij})\prod_{\begin{subarray}{c}{(i,u)}\\ u\neq j\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}{(j,v)}\\ v\neq i\end{subarray}}(1-p_{jv})\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}e^{-d_{i}-d_{j}+p_{ij}}\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}],$ (35)
+    where (a) uses that $1-x\leq e^{-x}$ for all $x$, and (b) uses that $p_{ij}\leq p_{\max}\leq 6p_{\max}$.
+  • Case $2$: $j=k$. Notice that
+    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}]\cdot\mathbb{E}[\widetilde{Z}_{j}^{2}]\cdot\mathbb{E}[\widetilde{Z}_{\ell}]=e^{-d_{i}-d_{j}-d_{\ell}}.$
+    Therefore, similar to case $1$, we have
+    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]$ $=\prod_{f\in\binom{\{i,j,\ell\}}{2}}(1-p_{f})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j,\ell\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i,\ell\end{subarray}}(1-p_{jv})\prod_{\begin{subarray}{c}(\ell,w)\\ w\neq i,j\end{subarray}}(1-p_{\ell w})$
+    $\leq e^{-d_{i}-d_{j}-d_{\ell}+p_{ij}+p_{i\ell}+p_{j\ell}}\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]$ (36)
+  • Case $3$: $i\neq j\neq k\neq\ell$. Notice that
+    $\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]=\mathbb{E}[\widetilde{Z}_{i}]\cdot\mathbb{E}[\widetilde{Z}_{j}]\cdot\mathbb{E}[\widetilde{Z}_{k}]\cdot\mathbb{E}[\widetilde{Z}_{\ell}]=e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}.$
+    Therefore, similar to case $1$, we have
+    $\mathbb{E}[Z_{i}Z_{j}Z_{k}Z_{\ell}]$ $=\prod_{f\in\binom{\{i,j,k,\ell\}}{2}}\!\!\!(1-p_{f})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j,k,\ell\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i,k,\ell\end{subarray}}(1-p_{jv})\prod_{\begin{subarray}{c}(k,w)\\ w\neq i,j,\ell\end{subarray}}(1-p_{kw})\prod_{\begin{subarray}{c}(\ell,x)\\ x\neq i,j,k\end{subarray}}(1-p_{\ell x})$
+    $\leq e^{-d_{i}-d_{j}-d_{k}-d_{\ell}+p_{ij}+p_{ik}+p_{i\ell}+p_{jk}+p_{j\ell}+p_{k\ell}}\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}\widetilde{Z}_{k}\widetilde{Z}_{\ell}]$ (37)
+Consequently,
+$\mathbb{E}[\widetilde{N}^{2}]$ $=\sum_{1\leq i<j\leq n}\sum_{1\leq k<\ell\leq n}\Big{[}e^{-d_{i}-d_{j}}\bm{1}_{\{i=k,j=\ell\}}+e^{-d_{i}-d_{j}-d_{\ell}}\bm{1}_{\{j=k\}}+e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}\bm{1}_{\{i\neq j\neq k\neq\ell\}}\Big{]}$
+$\leq\sum_{1\leq i<j\leq n}\sum_{1\leq k<\ell\leq n}\Big{[}e^{-d_{i}-d_{j}}+e^{-d_{i}-d_{j}-d_{\ell}}+e^{-d_{i}-d_{j}-d_{k}-d_{\ell}}\Big{]}\leq\mu^{2}+\mu^{3}+\mu^{4},$ (38)
+where $\mu\triangleq\sum_{1\leq i\leq n}e^{-d_{i}}$. Combining equations (33) and (34) with (35), (36), (37) and (38) yields that
+$\mathbb{E}[N^{2}]\leq e^{6p_{\max}}\cdot\mathbb{E}[\widetilde{N}^{2}]\leq e^{6p_{\max}}\mu^{2}(\mu^{2}+\mu+1).$ (39)
+Bounding the first moment. Notice that
+$\mathbb{E}[\widetilde{N}]=\sum_{1\leq i<j\leq n}\mathbb{E}[\widetilde{Z}_{i}\widetilde{Z}_{j}]=\sum_{1\leq i<j\leq n}e^{-d_{i}-d_{j}}=\frac{1}{2}\left(\mu^{2}-\sum_{i=1}^{n}e^{-2d_{i}}\right)\geq\frac{\mu}{2}\left(\mu-\max_{1\leq i\leq n}e^{-d_{i}}\right),$ (40)
+where the last step uses that $e^{-2d_{k}}\leq e^{-d_{k}}\cdot\max_{1\leq i\leq n}e^{-d_{i}}$ for each $k$. On the other hand,
+$\mathbb{E}[N]$ $=\sum_{1\leq i<j\leq n}\mathbb{E}\left[Z_{i}Z_{j}\right]=\sum_{1\leq i<j\leq n}\Big{[}(1-p_{ij})\prod_{\begin{subarray}{c}(i,u)\\ u\neq j\end{subarray}}(1-p_{iu})\prod_{\begin{subarray}{c}(j,v)\\ v\neq i\end{subarray}}(1-p_{jv})\Big{]}$
+$\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\sum_{1\leq i<j\leq n}\Big{[}e^{-p_{ij}-p_{ij}^{2}}\prod_{\begin{subarray}{c}(i,u)\\ u\neq j\end{subarray}}e^{-p_{iu}-p_{iu}^{2}}\prod_{\begin{subarray}{c}(j,v)\\ v\neq i\end{subarray}}e^{-p_{jv}-p_{jv}^{2}}\Big{]}$
+$\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}e^{-(2n-1)p_{\max}^{2}}\cdot\sum_{1\leq i<j\leq n}e^{-d_{i}-d_{j}+p_{ij}}\stackrel{{\scriptstyle\text{(c)}}}{{\geq}}e^{-(2n-1)p_{\max}^{2}}\cdot\mathbb{E}[\widetilde{N}],$ (41)
+where (a) uses that $1-x\geq e^{-x-x^{2}}$ for all $x\in[0,1/2]$, (b) uses that $e^{-p_{ij}^{2}}\geq e^{-p_{\max}^{2}}$, and (c) uses that $e^{p_{ij}}\geq 1$ for all $i$ and $j$.
+Finally, combining equations (39), (40) and (41), we have that
+$\frac{\mathbb{E}[N]^{2}}{\mathbb{E}[N^{2}]}\geq e^{-(4n-2)p_{\max}^{2}-6p_{\max}}\frac{\mathbb{E}[\widetilde{N}]^{2}}{\mathbb{E}[\widetilde{N}^{2}]}\geq\exp\left(-(4n-2)p_{\max}^{2}-6p_{\max}\right)\frac{\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}}{4(\mu^{2}+\mu+1)},$
+where the last step uses that
+$\mathbb{E}[\widetilde{N}]^{2}\geq\frac{\mu^{2}}{4}\left(\mu^{2}-2\mu\cdot\max_{1\leq i\leq n}e^{-d_{i}}+\max_{1\leq i\leq n}e^{-2d_{i}}\right)\geq\frac{\mu^{2}}{4}\left(\mu^{2}-2\mu\max_{1\leq i\leq n}e^{-d_{i}}\right).$
+[…]
+B.1 Proof of Lemma 9
+Proof of Lemma 9.    Let $N$ be the number of pairs of isolated nodes in $H$. It follows from Lemma 29 that
+$\mathbb{P}\left(N\geq 1\right)$ $\geq\exp\left(-(2n-4)q_{\max}^{2}-6q_{\max}\right)\cdot\frac{\lambda^{2}-2\lambda\max_{1\leq i\leq n}e^{-r_{i}}}{4(\lambda^{2}+\lambda+1)}$
+$\geq(1-o(1))\cdot\frac{\lambda^{2}-2\lambda\max_{1\leq i\leq n}e^{-r_{i}}}{4(\lambda^{2}+\lambda+1)},$ (42)
+where $\lambda\triangleq\sum_{i=1}^{n}e^{-r_{i}}$, and the last step in (42) uses that $q_{\max}=o(1/\sqrt{n})$.
+Let $r_{[1]}\leq r_{[2]}\leq\cdots\leq r_{[n]}$ denote a sorted copy of the sequence $r_{1},\cdots,r_{n}$.
+Consider the following three cases.
+  • Case $1$. $r_{[1]}=\omega(1)$. In this setting, it follows that
+    $\lambda\max_{1\leq i\leq n}e^{-r_{i}s(1-(1-s)^{m-1})}=\lambda e^{-r_{[1]}s(1-(1-s)^{m-1})}=o(1),$
+    since $\lambda=\Omega(1)$. Thus,
+    $\mathbb{P}\left(N\geq 1\right)\geq(1-o(1))\cdot\frac{\lambda^{2}-o(1)}{4(\lambda^{2}+\lambda+1)}=\Omega(1).$
+  • Case $2$.
+$r_{[2]}=O(1)$.
+In this setting, there exists a constant $C>0$ such that $C\leq e^{-r_{[2]}}\leq e^{-r_{[1]}}$.
+Let $A$ denote the adjacency matrix of $G$ and let $R_{[1]}$ and $R_{[2]}$ denote the degree of two nodes with mean degree $r_{[1]}$ and $r_{[2]}$ respectively.
+Then,
+    $\mathbb{P}\left(R_{[1]}=0,R_{[2]}=0\right)$ $=\mathbb{P}\left(R_{[1]}=0\right)\cdot\mathbb{P}\left(R_{[2]}=0\mid R_{[1]}=0\right)\geq\mathbb{P}\left(R_{[1]}=0\right)\mathbb{P}\left(R_{[2]}=0\right)$
+    $=\prod_{i\neq[1]}(1-q_{[1]i})\prod_{j\neq[2]}(1-q_{[2]j})\stackrel{{\scriptstyle\text{(a)}}}{{\geq}}\prod_{i\neq[1]}e^{-q_{[1]i}-q_{[1]i}^{2}}\prod_{j\neq[2]}e^{-q{[2]j}-q_{[2]j}^{2}}$
+    $=e^{-r_{[1]}}e^{-r_{[2]}}\cdot\exp\left(-\sum_{i\neq[1]}q_{[1]i}^{2}-\sum_{j\neq[2]}q_{[2]j}^{2}\right)$
+    $\stackrel{{\scriptstyle\text{(b)}}}{{\geq}}\exp\left(-2(n-1)q_{\max}^{2}\right)\cdot C^{2}\stackrel{{\scriptstyle\text{(c)}}}{{=}}\Omega(1).$
+    Here, (a) uses the inequality $1-x\geq e^{-x-x^{2}}$ whenever $x\in[0,1/2]$, (b) uses that $e^{-q_{ij}^{2}}\geq e^{-q_{\max}^{2}}$ for all $i,j$ and also that $C\leq e^{-r_{[2]}}\leq e^{-r_{[1]}}$.
+Finally, (c) uses that $\exp\left(-2(n-1)q_{\max}^{2}\right)=\Omega(1)$ whenever $q_{\max}=o(1/\sqrt{n})$.
+  • Case $3$.
+$r_{[2]}=\omega(1)$.
+Consider the induced subgraph $G^{\prime}$ of $G$ on the vertex set $V=V-\{[1]\}$.
+The mean degrees of $G^{\prime}$ satisfy $r_{[i]}^{\prime}\leq r_{[i]}$, and so it follows that
+    $\lambda^{\prime}\triangleq\sum_{i=2}^{n}e^{-d^{\prime}_{[i]}}\leq\sum_{i=2}^{n}e^{-r_{[i]}}=\Omega(1),$
+    and the smallest mean degree in $G^{\prime}$ diverges, i.e.
+$r_{[2]}=\omega(1)$.
+Equivalently, the graph $G^{\prime}$ satisfies the conditions for Case $1$ above, and so there is a constant lower bound $C$ on the probability that the number of pairs of isolated nodes $N^{\prime}$ in $G^{\prime}$ is at least one.
+If node $[1]$ does not connect to an isolated pair of nodes in $G^{\prime}$, then the pair remains isolated in $G$ as well.
 Therefore,
+    $\mathbb{P}\left(N\geq 1\right)\geq(1-q_{\max})^{2}\cdot\mathbb{P}\left(N^{\prime}\geq 1\right)\geq(1-q_{\max})^{2}\cdot C=\Omega(1).$
 […]
-=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\mathbb{P}\left(\mathsf{HypGeom}(n-1,n-1-\lceil n-3n^{1-\alpha}\rceil,\lceil k+1/\alpha\rceil)\geq\lceil k+1/\alpha\rceil-(k-1)\right)$
-$=\mathbb{P}\left(\mathsf{HypGeom}(n-1,\lfloor 3n^{1-\alpha}\rfloor-1,\lceil k+1/\alpha\rceil)\geq 1+1/\alpha\right)$ (9)
-where (a) uses the symmetry of the hypergeometric distribution. Using Proposition 15 and the fact that $n-1\geq n/2$ for any $n\geq 1$ yields
-[…]
-$\leq\left(\frac{6e\lceil k+1/\alpha\rceil}{1+1/\alpha}\right)^{1+1/\alpha}\times n^{-1-\alpha}$
-$=o(1/n),$
-whenever $\alpha>0$. ∎
-C.1 Proof of Lemma 14
-A key ingredient towards proving Lemma 14 is a useful result about the number of low-degree vertices in an Erdős-Rényi graph, presented next.
-Proposition 16.
-Let $\alpha>0$ and $G\sim\mathsf{ER}\left(n-1,\alpha\log(n)/n\right)$. Let $r$ be a positive integer and let $Z_{r}$ denote the set of vertices in $G$ with degree no more than $r$, i.e.
-$Z_{r}=\left\{v\in V(G):\delta_{G}(v)\leq r\right\}.$
-For any $\delta$ such that $\delta>1-\alpha$, it is true that
-$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)=o(1/n).$
-Proof.
-Notice that
-$\mathbb{P}\left(|Z_{r}|\geq n^{\delta}\right)$ $=\mathbb{P}\left(\exists S^{\prime}\subseteq V:\left\{|S^{\prime}|\geq n^{\delta}\right\}\cap\left\{\max_{i\in S^{\prime}}\delta_{G}(i)\leq r\right\}\right)$
-$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\max_{i\in S}\delta_{G}(i)\leq r\right\}\right)$
-$\leq\mathbb{P}\left(\exists\ S\subseteq V:\ \left\{|S|=n^{\delta}\right\}\cap\left\{\sum_{i\in S}\delta_{G}(i)\leq r|S|\right\}\right).$ (10)
-If $|S|=n^{\delta}$, then the sum of degrees of vertices in $S$ is the total number of edges with exactly end point in $S$, plus twice the number of edges with both end points in $S$.
-There are exactly $\binom{|S|}{2}+|S|\left(n-1-|S|\right)\leq n^{1+\delta}$ such vertex pairs, and each of them independently has an edge with probability $\alpha\log(n)/n$.
-Therefore, a union bound over all possible choices of $S$ yields
-[…]
-$\leq\binom{n-1}{n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(n^{1+\delta},\alpha\log(n)/n\right)\leq rn^{\delta}\right)$
-$\stackrel{{\scriptstyle\text{(a)}}}{{\leq}}\left(\frac{ne}{n^{\delta}}\right)^{n^{\delta}}\times\left(\frac{\exp\left(r/(\alpha\log n)-1\right)}{(r/(\alpha\log n))^{r/(\alpha\log n)}}\right)^{n^{\delta}\alpha\log(n)}$
-$=o(1/n),$
-whenever $\delta>1-\alpha$ as desired. Note that (a) uses the Binomial concentration inequality (7) and the fact that $\binom{n-1}{k}\leq\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$. ∎
-Our objective is to show that the $k$-core of $G-v$ is sufficiently large with probability $1-o(1/n)$.
-To that end, consider Algorithm 3 to identify a subset of the $k$-core, originally proposed by Łuczak [Łuc91].
-require :  Graph $G$, Set $U\subseteq V(G)$.
-1 $U_{0}\leftarrow U$
-2 for $i=0,1,2,3,\cdots$ do
-3       if  there exists $u\in V\setminus U_{i}$ such that $u$ has $3$ or more neighbors in $U_{i}$ then
-4             $U_{i+1}\leftarrow U_{i}\cup\{u\}$
-5
-6       else
-7             return $U_{i}$
-8
-9       end if
-10
-11 end for
-Algorithm 3 Łuczak expansion
-Note that the for loop eventually terminates - the set $V\setminus U_{i}$ is empty, for example when $i=n$ for any input set $U$.
-The key is to realize that the for loop terminates much faster when the input $U=Z_{k+1}$ , i.e the set of vertices of the input graph $G$ whose degree is $k+1$ or less.
-Furthermore, the complement of the set output by the algorithm is contained in the $k$-core.
-Formally,
-Lemma 17.
-Let $U_{f}$ be the output of Algorithm 3 with input graph $G-v$ and set $U=Z_{k+1}$. Then,
-  1. (a)
-    $U_{f}^{c}\subseteq\mathsf{core}_{k}(G-v)$.
-  2. (b)
-    For any $\delta>1-\alpha$,
-    $\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)=o(1/n).$
-Proof.
-(a) The proof is by construction: Since $U_{f}$ is obtained by adding exactly $f$ nodes to $U_{0}$, it follows that $U_{f}^{c}\subseteq U_{0}^{c}=Z_{k+1}^{c}$, so each node in $U_{f}^{c}$ has degree $k+2$ or more in $G-v$.
-Further, each node in $U_{f}^{c}$ has at most $2$ neighbors in $U_{f}$, else the for loop would not have terminated.
-Thus, the subgraph of $G-v$ induced on the set $U_{f}^{c}$ has minimum degree at least $k$, and the result follows.
-(b) If $|U_{f}|>3n^{\delta}$, then either $|U_{0}|>3n^{\delta}$ or there is some $M$ in $\{0,1,\cdots,f\}$ for which $|U_{M}|=3n^{\delta}$. Therefore,
-$\mathbb{P}\left(|U_{f}|>3n^{\delta}\right)$ $\leq\mathbb{P}\left(|U_{0}|>3n^{\delta}\right)+\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)$
-$=o(1/n)+\underbrace{\mathbb{P}\left(\exists\ M\in\{0,1,\cdots,f\}\text{ s.t. }|U_{M}|=3n^{\delta}\right)}_{(\star)},$
-by Proposition 16.
-Note that each iteration $i=0,1,\cdots,M-1$ of the for loop adds exactly $1$ vertex and at least $3$ edges to the subgraph of $G-v$ induced on $U_{M}$.
-Therefore, the induced subgraph $G|_{U_{M}}$ has $3n^{\delta}$ vertices and at least $3\left(|U_{M}|-|U_{0}|\right)$ edges.
-Thus,
-(($\star$)( ⋆ ) ( $\leq\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{ and }|F|\geq 3\left(3n^{\delta}-|U_{0}|\right)\right)$
-$\leq\mathbb{P}\left(|U_{0}|>n^{\delta}\right)+\mathbb{P}\left(\exists\text{ subgraph $H\!=\!(W,F)$ of }G-v\text{ s.t. }|W|=3n^{\delta}\text{and }|F|\geq 6n^{\delta}\right)$
-$\leq o(1/n)+\underbrace{\binom{n}{3n^{\delta}}\cdot\mathbb{P}\left(\mathsf{Bin}\left(\binom{3n^{\delta}}{2},\frac{\alpha\log(n)}{n}\right)>6n^{\delta}\right)}_{(\star\star)},$
-where the last step uses Proposition 16 and a union bound over all possible choices of $W$.
-Finally, using the relation $\binom{n}{k}\leq\left(\frac{ne}{k}\right)^{k}$ and the concentration inequality (5) from Lemma 13 yields
-[…]
-$\leq(n^{1-\delta})^{3n^{\delta}}\times\left(\frac{3\alpha e\log n}{4n^{1-\delta}}\right)^{6n^{\delta}}$
-$=\left(\frac{3\alpha e\log n}{4n^{(1-\delta)/2}}\right)^{6n^{\delta}}$
-$=o(1/n),$
-whenever $10<\delta<$ 1. The result follows. ∎
-Finally, notice that Lemma 17 directly implies Lemma 14.
-Appendix D Proof of Theorem 9
-See 9
-Proof.
-For any vertex cut $U$,
-$\left\{\widetilde{c}_{v}(U)>\frac{m^{2}}{4}\left(k+\frac{1}{Cs^{2}}\right)\right\}$ $\stackrel{{\scriptstyle\text{(a)}}}{{\subseteq}}\left\{\widetilde{c}_{v}(U)>|U|\left(m-|U|\right)\left(k+\frac{1}{Cs^{2}}\right)\right\}$
-=$=\bigg{\{}\sum_{i\in U}\sum_{j\in U^{c}}\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>|U|(m-|U|)\bigg{(}k+\frac{1}{Cs^{2}}\bigg{)}\bigg{\}}$
-$\subseteq\bigcup_{i\in U}\bigcup_{j\in U^{c}}\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\},$
-where (a) uses the fact that the maximum of a set of a numbers is greater than or equal to the average. On the other hand
-$\left\{c_{v}(U)=0\right\}=\bigcap_{i\in U}\bigcap_{j\in U^{c}}\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}.$
-Let $p_{1}$ denote the probability in the LHS of (2). It follows from the union bound that
-$p_{1}\leq\sum_{i\in U}\sum_{j\in U^{c}}\mathbb{P}\left(\left\{v\notin\mathsf{core}_{k}(G_{i}^{\prime}\wedge G_{j}^{\prime})\right\}\cap\left\{\delta_{G_{i}^{\prime}\wedge G_{j}^{\prime}}(v)>k+\frac{1}{Cs^{2}}\right\}\right)=o(1/n),$
-since for any choice of $i$ and $j$, the graph $G_{i}^{\prime}\wedge G_{j}^{\prime}\sim\mathsf{ER}\left(n,Cs^{2}\log(n)/n\right)$. ∎
-Appendix E On Stochastic Dominance: Proof of Theorem 10
-The objective of this section is to build up to a proof of Theorem 10. We start by making a simple observation about products of Binomial random variables.
-Lemma 18.
-Let $X_{1},\cdots,X_{m}\sim\mathsf{Bern}(s)$ be i.i.d. random variables, and let $B=X_{1}+\cdots+X_{m}$ denote their sum. For each $\ell$ in $\{1,2,\cdots,\lfloor m/2\rfloor\}$, define
+Appendix C Stochastic domination of crossing edges in ~$\widetilde{\mathcal{H}}(v)$: Proof of Lemma 25
+The objective of this section is to build up to a proof of Lemma 25. We start by making a simple observation about products of Binomial random variables.
+Lemma 30
+Let $X_{1},\cdots,X_{m}\sim\mathrm{Bern}(s)$ be i.i.d. random variables, and let $B=X_{1}+\cdots+X_{m}$ denote their sum. For each $\ell$ in $\{1,2,\cdots,\lfloor m/2\rfloor\}$, define
 $T_{\ell}=\left(X_{1}+\cdots+X_{\ell}\right)\left(X_{\ell+1}+\cdots+X_{m}\right).$
@@
 This concludes the proof.
-∎
-Corollary 19.
+$\Box$
+Corollary 31
 Let $F$ be a collection of edges in the parent graph $G$.
-For any edge $e_{r}\in F$, let $X^{r}_{i}$ denote the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$.
+For any edge $e_{r}\in F$, let $X^{r}_{i}$ denote the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathrm{Bern}(ps)$.
 For each $\ell$ in $\{1,\cdots,\lfloor m/2\rfloor\}$, define
@@
 $\sum_{r=1}^{|F|}T^{r}_{\ell_{1}}\preceq\sum_{r=1}^{|F|}T^{r}_{\ell_{2}}.$
-Proof.
-It suffices to show that $T^{r}_{\ell_{1}}\preceq T^{r}_{\ell_{2}}$ for each $r$, since the edges are independent. Indeed, we have for any $t$ that
+Proof. It suffices to show that $T^{r}_{\ell_{1}}\preceq T^{r}_{\ell_{2}}$ for each $r$, since the edges are independent. For any $t$,
 $\mathbb{P}\left(T^{r}_{\ell_{1}}>t\right)=\sum_{b=0}^{m}\mathbb{P}\left(B=b\right)\mathbb{P}\left(T_{\ell_{1}}^{r}>t|B=b\right)\leq\sum_{b=0}^{m}\mathbb{P}\left(B=b\right)\mathbb{P}\left(T_{\ell_{2}}^{r}>t|B=b\right)=\mathbb{P}\left(T^{r}_{\ell_{2}}>t\right),$
-which concludes the proof. ∎
-With this, we are ready to prove Theorem 10. The theorem is restated for convenience.
-See 10
-Proof.
+[…]
+With this, we are ready to prove Lemma 25. The lemma is restated for convenience.
+See 25 Proof.
 Let $\ell_{1},\ell_{2}\in\{1,\cdots,\lfloor m/2\rfloor\}$ such that $\ell_{1}<\ell_{2}$.
@@
 Consider the parent graph $G$ and label the set of incident edges on $v$ as $\{e_{1},\cdots,e_{\delta_{G}(v)}\}$.
-Denote by $X_{i}^{r}$ the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathsf{Bern}(ps)$.
+Denote by $X_{i}^{r}$ the indicator random variable $G_{i}^{\prime}(e_{r})\sim\mathrm{Bern}(ps)$.
 It follows that
@@
 $=\mathbb{P}\left(\widetilde{c}_{v}(U_{\ell_{1}})>t\right),$
-as desired. Here, (a) uses Corollary 19. ∎
-Appendix F On Low Degree Nodes: Proof of Theorem 11
-See 11
-Proof.
-Consider fixed integers $r_{1},\cdots,r_{m}$ such that $0\leq r_{2},\cdots,r_{m}\leq r$. Since $r$ is constant, by a union bound argument it suffices to show
-()=:($(\star)=:\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{2}^{\prime}}(v)=r_{2}\right\}\cap\cdots\cap\left\{\delta_{G_{1}\wedge G_{m}^{\prime}}(v)=r_{m}\right\}\right)=o\left(1/n\right).$
-Proceed by conditioning on the degree of $v$ in $G_{1}$, which follows a $\mathsf{Bin}(n,ps)$ distribution.
-Since the degrees of $v$ in the intersection graphs $\{G_{1}\wedge G_{i}^{\prime}:i=2,\cdots,m\}$ are conditionally independent given the degree of $v$ in $G_{1}$, we have
-(($\star$)( ⋆ ) $=\mathbb{E}_{D}\left[\mathbb{P}\Big{(}\bigcap_{i=2}^{m}\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\Big{)}\right]$
-$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\mathbb{P}\left(\left\{\delta_{G_{1}\wedge G_{i}^{\prime}}(v)=r_{i}\right\}\Big{|}\delta_{G_{1}}(v)=D\right)\right]$
-$=\mathbb{E}_{D}\left[\prod_{i=2}^{m}\binom{D}{r_{i}}s^{r_{i}}(1-s)^{D-r_{i}}\right].$ (13)
-Using the fact that $\binom{D}{r_{i}}\leq\left(\frac{De}{r_{i}}\right)^{r_{i}}$, it follows that
 […]
-[…]
-Expanding out the expectation yields
-[…]
-Proceed by splitting the summation at $(\log n)^{2}$. The first part can be bounded as
-$\sum_{d=0}^{\left(\log n\right)^{2}}L_{d}$ $\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\sum_{d=0}^{(\log n)^{2}}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
-$\leq\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\sum_{d=0}^{n}\left(1-s\right)^{(m-1)d}\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)=d\right)$
-$=\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\mathbb{E}_{D}\left[(1-s)^{(m-1)D}\right]$
-=$\stackrel{{\scriptstyle\text{(a)}}}{{=}}\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}\cdot\left(1-\frac{Cs\left(1-(1-s)^{m-1}\right)\log n}{n}\right)^{n}$
-$=o(1/n),$
-whenever $Cs(1-(1-s))^{m-1}>1$. Here, (a) is obtained by evaluating the probability generating function of the $\mathsf{Bin}(n,ps)$ random variable at $(1-s)^{m-1}$ and setting $p=C\log(n)/n$.
-The other part of the sum can now be bounded as follows.
-$\sum_{d=\left(\log n\right)^{2}}^{n}L_{d}$ $\leq\left[\max_{d\colon(\log n)^{2}\leq d\leq n}d^{\sum_{i=2}^{m}r_{i}}(1-s)^{md}\right]\cdot\mathbb{P}\left(\mathsf{Bin}(n,ps)\geq(\log n)^{2}\right)$
-$\stackrel{{\scriptstyle\text{(b)}}}{{\leq}}\left[\left(\log n\right)^{2\sum_{i=2}^{m}r_{i}}(1-s)^{m\left(\log n\right)^{2}}\right]\times 2^{-\left(\log n\right)^{2}}$
-$=(\log n)^{2\sum_{i=2}^{m}r_{i}}\left(\frac{(1-s)^{m}}{2}\right)^{(\log n)^{2}}$
-$=o\left(1/n\right).$
-whenever $C>0$.
-Here, (b) is true because the function $d\mapsto d^{\sum r_{i}}(a-s)^{md}$ is decreasing on the interval $[(\log n)^{2},n]$ for all sufficiently large $n$.
-Finally, the concentration inequality for the Binomial distribution holds by (6) in Lemma 13.
-The inequality applies since $p=C\log(n)/n$ and since $(\log n)^{2}>6Cs\log(n)$ for all $n$ sufficiently large.
-This concludes the proof.
-∎